### v1

# Estimation of relative risk, odds ratio and their logarithms with guaranteed accuracy and controlled sample size ratio

Abstract
Given two populations from which independent binary observations are taken with parameters $p_{1}$ and $p_{2}$ respectively, estimators are proposed for the relative risk $p_{1}/p_{2}$, the odds ratio $p_{1}(1-p_{2})/(p_{2}(1-p_{1}))$ and their logarithms. The estimators guarantee that the relative mean-square error, or the mean-square error for the logarithmic versions, is less than a target value for any $p_{1},p_{2}\in(0,1)$, and the ratio of average sample sizes from the two populations is close to a prescribed value. The estimators can also be used with group sampling, whereby samples are taken in batches of fixed size from the two populations. The efficiency of the estimators with respect to the Cramér–Rao bound is good, and in particular it is close to $1$ for small values of the target error.
keywords:
Estimation, sequential sampling, group sampling, relative risk, odds ratio, log odds ratio, mean-square error, efficiency.
pacs:
MSC2010 Classification:]62F10, 62L12
1 Introduction
Let $p_{1},p_{2}\in(0,1)$ denote the probabilities of occurrence of a given dichotomous attribute in two different populations. The problem of estimating the relative risk (RR) or risk ratio,
$\theta=\frac{p_{1}}{p_{2}},$ (1)
from binary observations of the two populations arises frequently in medical and social sciences, as well as in other fields. For example, in a phase-III clinical trial of a vaccine [4, chapter 18] the relevant attribute is the presence of a disease, and the two populations are vaccinated and non-vaccinated people. The odds ratio (OR),
$\psi=\frac{p_{1}(1-p_{2})}{p_{2}(1-p_{1})},$ (2)
is sometimes used instead of RR, as are their logarithmic versions: log relative risk (LRR), $\Theta=\log\theta$, and log odds ratio (LOR), $\Psi=\log\psi$. The latter is especially important in connection with logistic regression [3], which is a prevalent tool in machine learning [7].
When estimating any of these parameters, it is crucial to have knowledge about the accuracy of the estimation. Ideally, the estimation should guarantee a target accuracy, regardless of the unknown $p_{1}$ and $p_{2}$. A common measure of accuracy is the mean-square error (MSE), or its square root, known as root-mean-square error (RMSE), which for unbiased estimators reduce to variance or standard deviation respectively. For non-logarithmic parameters such as RR and OR, these error measurements are meaningful in a relative sense [16], because the significance of a given estimation error can only be assessed by comparing it with the true value of the parameter. Thus, it is natural to require that the RMSE be proportional to the true value of the parameter. On the other hand, for LRR and LOR the estimation error is meaningful by itself, without comparing with the true value, as the logarithm transforms multiplication into addition.
A second desirable feature, along with guaranteed accuracy, is to have control on the proportion of sample sizes of the two populations. Consider first the case that the two populations are sampled separately, meaning that samples from one or the other population can be taken as needed. In many use cases, it may be required that the two sample sizes be similar, or that they approximately satisfy a given ratio. Another possible sampling procedure is group sampling, whereby samples are collected in groups or batches of $l_{1}$ and $l_{2}$ samples from the two populations respectively. This imposes a strict sample size ratio of $l_{1}/l_{2}$. Either with separate sampling or with group sampling, these conditions only refer to the sample size ratio; the actual sample sizes should be chosen to fulfill the target accuracy.
It will be assumed that each population is infinite, and observations are statistically independent. This implies that the observations can be modeled as Bernoulli trials. The requirement that the target accuracy be satisfied for all $p_{1},p_{2}\in(0,1)$ makes it necessary to use sequential sampling, because any fixed sample size will fail to satisfy that requirement for low enough $p_{1},p_{2}$. More specifically, this work will make extensive use of inverse binomial sampling (IBS) 12; 15, chapter 2. Denoting the presence or absence of the attribute of interest in a sample as “success” or “failure”, IBS consists in observing samples until a predefined number $r$ of successes is obtained. The number $r$ will be referred to as the parameter of the IBS procedure.
This paper presents unbiased estimators of the four parameters RR, LRR, OR and LOR, that guarantee a target accuracy and provide control on the sample size ratio, irrespective of $p_{1}$ and $p_{2}$. The estimators are based on the two-stage sampling procedure suggested in a previous work [16, section 4]. As argued previously, the target accuracy, $A$, is defined as relative MSE for RR and OR, or as MSE for LRR and LOR. The control on the sample size ratio means that, for separate sampling of each population, average sample sizes will approximately satisfy a specified proportion. The estimators can also be applied with group sequential sampling, using groups of $l_{1}$ and $l_{2}$ samples from each population. This incurs, as will be seen, a small increase in the average sample size compared to separate sampling, but ensures an exact ratio of the (random) sample sizes. In both cases the estimation efficiency, defined in terms of the Cramér–Rao bound, is considerably large, and is close to $1$ for small $A$.
Estimation of the RR, OR or their logarithmic versions is an important problem in statistics. For a review of results in this area, see Mendo [16, section 1]. Variants of the problem considering a desired ratio of sample sizes have been addressed in existing works (often assuming the ratio equal to $1$); see for example Siegmund [20], Agresti [2], Cho [9], Cho [10]. However, to the author’s knowledge, no previously proposed estimators for these parameters guarantee a target accuracy as defined above, i.e. relative error for RR and OR or absolute error for LRR and LOR, while offering control on the proportion of sample sizes.
The following notation and basic identities will be used. The function $\log x$ represents the natural logarithm of $x$. The $n$-th harmonic number is denoted as
$H_{n}=\sum_{k=1}^{n}\frac{1}{k}.$ (3)
Matrices are written in boldface letters, and $\mathbf{Q}^{\intercal}$ represents the transpose of a matrix $\mathbf{Q}$. The regularized incomplete beta function is defined as
$I(x;u,v)=\frac{1}{\mathrm{B}(u,v)}\int_{0}^{x}t^{u-1}(1-t)^{v-1}\,\mathrm{d}t,$ (4)
where $\mathrm{B}(u,v)$ is the beta function; and from Abramowitz and Stegun [1, equations (6.1.15), (6.2.2), (26.5.15)] it follows that
$(v-1)\mathrm{B}(u+1,v-1)$ $=u\,\mathrm{B}(u,v),$ (5)
$(u-1)\mathrm{B}(u-1,v+1)$ $=v\,\mathrm{B}(u,v),$ (6)
$I(x;u,v)-I(x;u+1,v-1)$ $=\frac{x^{u}(1-x)^{v-1}}{u\,\mathrm{B}(u,v)},$ (7)
$I(x;u,v)-I(x;u-1,v+1)$ $=-\frac{x^{u-1}(1-x)^{v}}{v\,\mathrm{B}(u,v)}.$ (8)
The probability density function of a beta prime distribution with parameters $u$, $v$ is denoted as $f(y;u,v)$:
$f(y;u,v)=\frac{y^{u-1}}{\mathrm{B}(u,v)(1+y)^{u+v}},\quad y>0.$ (9)
For a random variable $Y$ with this distribution [8, section 4.4],
$\Pr[Y\leq y]=\int_{0}^{y}f(t;u,v)\,\mathrm{d}t=I\left(\frac{y}{y+1};u,v\right).$ (10)
In IBS, the number $N$ of samples needed to obtain $r$ successes has a negative binomial distribution with parameters $r$ and $p$. Then, for $p\in(0,1)$ 18, equation (3.1); 19, section 4.8.2,
$\operatorname{E}[N]$ $=\frac{r}{p},$ (11)
$\operatorname{E}\left[\frac{1}{N-1}\right]$ $=\frac{p}{r-1}\quad\text{for }r\geq 2,$ (12)
$\operatorname{Var}[N]$ $=\frac{r(1-p)}{p^{2}},$ (13)
$\operatorname{Var}\left[\frac{1}{N-1}\right]$ $\leq\frac{p^{2}(1-p)}{(r-1)^{2}(r-2+2p)}<\frac{p^{2}(1-p)}{(r-1)^{2}(r-2)}\quad\text{for }r\geq 3.$ (14)
The rest of the paper is organized as follows. Section 2 describes the estimation procedure for RR, derives approximate expressions and bounds for the average sample sizes and estimation efficiency, and compares these with values obtained from Monte Carlo simulations. The estimation procedure and the results for LRR, OR and LOR are to some extent analogous, and are presented in Sections 3, 4 and 5. Concluding remarks are given in Section 6.
2 Estimation of relative risk
The general estimation procedure for RR is motivated and described in Subsection 2.1, considering separate sampling of the two populations. Subsections 2.22.4 complete the definition of the estimator, obtain theoretical bounds for the average sample sizes and estimation efficiency, and compare them with simulation results. Subsection 2.5 describes the estimation method with group sampling, and then Subsections 2.6 and 2.7 obtain approximations for the average number of groups and efficiency in this case, and compare with simulation results.
2.1 Separate sampling. Estimation procedure
The estimator to be presented is unbiased, and uses two sampling stages. Each stage is comprised of two independent IBS procedures, one for each population. Before explaining the purpose of each stage, it is necessary to define several parameters and variables. The IBS parameters of the first stage are denoted as $r_{1}$ and $r_{2}$ for the two populations respectively. The resulting numbers of samples are negative binomial random variables, $M_{1}$ and $M_{2}$. Similarly, the second-stage IBS procedures have parameters $s_{1}$ and $s_{2}$, and the numbers of samples are $N_{1}$ and $N_{2}$. The total number of samples used from each population $i=1,2$ is then $M_{i}+N_{i}$, and from (11) it stems that
$\operatorname{E}[M_{i}+N_{i}]=\frac{r_{i}+s_{i}}{p_{i}}.$ (15)
Consider a target relative MSE equal to $A$, and a desired ratio $\lambda$ of average sample sizes. Let $\hat{\theta}$ denote the estimation of $\theta$. Since MSE reduces to variance for an unbiased estimator, the conditions that $\hat{\theta}$ must satisfy are
$\frac{\operatorname{Var}[\hat{\theta}]}{\theta^{2}}$ $\leq A,$ (16)
$\frac{\operatorname{E}[M_{1}+N_{1}]}{\operatorname{E}[M_{2}+N_{2}]}$ $\approx\lambda$ (17)
for any $p_{1},p_{2}\in(0,1)$.
The purpose of the first sampling stage is to obtain two pilot sets of samples, one from each population, using predefined values for the IBS parameters $r_{1}$ and $r_{2}$; and from those acquire some knowledge about $\theta$. With this knowledge, suitable values for the second-stage IBS parameters $s_{1}$ and $s_{2}$ are computed such that (16) and (17) are satisfied. The results of the second stage, i.e. $N_{1}$ and $N_{2}$, are then used to produce the final estimate $\hat{\theta}$. The rationale is as follows. According to (15),
$\frac{\operatorname{E}[M_{1}+N_{1}]}{\operatorname{E}[M_{2}+N_{2}]}=\frac{r_{1}+s_{1}}{(r_{2}+s_{2})\theta}.$ (18)
Each of the IBS procedures in the second stage provides information about one of the two probabilities. Given a target accuracy for the second-stage estimate $\hat{\theta}$, there is a trade-off between $s_{1}$ and $s_{2}$: decreasing one of them causes the information on the corresponding probability to be less accurate, which can be compensated for by increasing the other parameter. In view of (18), this can be exploited for balancing $\operatorname{E}[M_{1}+N_{1}]$ and $\operatorname{E}[M_{2}+N_{2}]$; doing so requires knowledge about $\theta$, which is provided by the first stage.
An initial idea to obtain information about $\theta$ from the first-stage variables $M_{1}$ and $M_{2}$ is to compute an estimate of this parameter using a generalization of the method described in Mendo [16, section 2] for estimating $p/(1-p)$. Namely, from (11) and (12) it follows that $(r_{1}-1)/(M_{1}-1)$ is an unbiased estimator of $p_{1}$ and $M_{2}/r_{2}$ is an unbiased estimator of $1/p_{2}$. Therefore, since the observations used by those estimators are independent,
$\frac{(r_{1}-1)\,M_{2}}{r_{2}\,(M_{1}-1)}$
is an unbiased estimator of $\theta$. Replacing $\theta$ in (18) by this estimate suggests that the condition (17) will be roughly satisfied if $s_{1}$ and $s_{2}$ are chosen so that
$\frac{r_{1}+s_{1}}{r_{2}+s_{2}}\approx\lambda\,\frac{(r_{1}-1)\,M_{2}}{r_{2}\,(M_{1}-1)}.$ (19)
It is more convenient, however, to substitute the requirement (19) by a generalized version:
$\frac{s_{1}+\delta_{1}}{s_{2}+\delta_{2}}=\gamma\,X$ (20)
with $X$ defined as
$X=\frac{M_{2}-\epsilon_{2}}{M_{1}-\epsilon_{1}},$ (21)
where $\gamma>0$, $\delta_{i}\mathbb{\in}\mathbb{R}$, $\epsilon_{i}\in(0,1)$, $i=1,2$ are design parameters, whose values can be selected to facilitate meeting (17) with good approximation. (Observe that (20) indeed reduces to (19) for $\gamma=\lambda(r_{1}-1)/r_{2}$, $\delta_{i}=r_{i}$, $\epsilon_{1}=1$, $\epsilon_{2}=0$.) In addition, $s_{i}$ and $\delta_{i}$ must satisfy
$s_{i}+\delta_{i}>0,\quad i=1,2,$ (22)
to ensure that the left-hand side of (20) does not involve division by zero and remains positive. It should be noted that (20) may give non-integer values for $s_{1}$ and $s_{2}$, which thus have to be rounded. This will introduce a small additional error in (17).
For $r_{1}\geq 3$, a simple analysis based on (13) and (14) shows that the relative variance of $X$ is bounded uniformly on $p_{1}$, $p_{2}$. On the other hand, for $r_{1}=2$ it is easy to see that the relative variance takes arbitrarily large values as $p_{1}\rightarrow 0$. Thus, to ensure that the variability of $X$ is not too large, the following additional requirement is imposed:
$r_{i}\geq 3,\quad i=1,2.$ (23)
The second sampling stage uses the parameters $s_{1}$ and $s_{2}$, determined in the first stage, to obtain $N_{1}$ and $N_{2}$. By the same reasoning applied earlier,
$\hat{\theta}=\frac{(s_{1}-1)\,N_{2}}{s_{2}\,(N_{1}-1)}$ (24)
is a conditionally unbiased estimation of $\theta$ given $s_{1},s_{2}$; and thus it is also unconditionally unbiased. For $s_{1}\geq 3$ the conditional variance of $(s_{1}-1)/(N_{1}-1)$ can be bounded, using (14), as
$\operatorname{Var}\left[\frac{s_{1}-1}{N_{1}-1}\mathrel{}\middle|\mathrel{}s_{1}\right]\leq\frac{p_{1}^{2}(1-p_{1})}{s_{1}-2+2p_{1}};$ (25)
and for $s_{2}\geq 1$ the conditional variance of $N_{2}/s_{2}$ is, according to (13),
$\operatorname{Var}\left[\frac{N_{2}}{s_{2}}\mathrel{}\middle|\mathrel{}s_{1}\right]=\frac{1-p_{2}}{s_{2}p_{2}^{2}}.$ (26)
Therefore,
$\begin{split}\frac{\operatorname{E}\left[\hat{\theta}^{2}\mid s_{1},s_{2}\right]}{\theta^{2}}&=\frac{p_{2}^{2}}{p_{1}^{2}}\left(\operatorname{Var}\left[\frac{s_{1}-1}{N_{1}-1}\mathrel{}\middle|\mathrel{}s_{1}\right]+p_{1}^{2}\right)\left(\operatorname{Var}\left[\frac{N_{2}}{s_{2}}\mathrel{}\middle|\mathrel{}s_{2}\right]+\frac{1}{p_{2}^{2}}\right)\\ &\leq\left(\frac{1-p_{1}}{s_{1}-2+2p_{1}}+1\right)\left(\frac{1-p_{2}}{s_{2}}+1\right).\end{split}$ (27)
This implies that, for all $p_{1},p_{2}\in(0,1)$,
$\frac{\operatorname{E}\left[\hat{\theta}^{2}\mid s_{1},s_{2}\right]}{\theta^{2}}<\frac{1}{s_{1}-2}+\frac{1}{s_{2}}+\frac{1}{(s_{1}-2)s_{2}}+1.$ (28)
In view of (28), let the function $e(s_{1},s_{2})$ be defined as
$e(s_{1},s_{2})=\frac{1}{s_{1}-\mu_{1}}+\frac{1}{s_{2}-\mu_{2}}+\frac{\mu_{12}}{(s_{1}-\mu_{1})(s_{2}-\mu_{2})},\quad s_{1}>\mu_{1},\ s_{2}>\mu_{2},$ (29)
with
$\mu_{1}=2,\qquad\mu_{2}=0,\qquad\mu_{12}=1.$ (30)
This will be referred to as error function. The parameters $\mu_{1},\mu_{2},\mu_{12}$ are introduced for convenience; this way the expression of $e(s_{1},s_{2})$ for other estimators will be the same as (29), only with different values of these parameters. Then, requiring
$e(s_{1},s_{2})\leq A$ (31)
guarantees that condition (16) holds; in fact with strict inequality. Namely, from (28)–(31),
$\frac{\operatorname{Var}\left[\hat{\theta}\right]}{\theta^{2}}=\frac{\operatorname{E}\left[\hat{\theta}^{2}\right]}{\theta^{2}}-\frac{\left(\operatorname{E}\left[\hat{\theta}\right]\right)^{2}}{\theta^{2}}=\frac{\operatorname{E}\left[\operatorname{E}\left[\hat{\theta}^{2}\mid s_{1},s_{2}\right]\right]}{\theta^{2}}-1<\operatorname{E}\left[e(s_{1},s_{2})\right]\leq A.$ (32)
2.2 Estimator parameters. Approximate average sample sizes
According to (15), to reduce the average sample sizes, $r_{i}$ and $s_{i}$, $i=1,2$ should be as small as possible. In this subsection approximate expressions are first derived for the average sample sizes, from which the choice of $s_{1}$, $s_{2}$ is addressed and the values of the estimator parameters $\gamma$, $\delta_{1}$, $\delta_{2}$, $\epsilon_{1}$, $\epsilon_{2}$ are determined. (Subsection 2.3 will discuss how to select the values of $r_{1}$, $r_{2}$.)
Since $e(s_{1},s_{2})$ is a decreasing function of $s_{1}$ and $s_{2}$, to minimize these parameters (31) should be treated as an equality. Thus, $s_{1}$ and $s_{2}$ are determined by
$e(s_{1},s_{2})=A$ (33)
together with (20) (where the values of the design parameters $\gamma$, $\delta_{1}$ and $\delta_{2}$ are yet to be defined). Solving this quadratic equation system yields
$s_{1}$ $=\frac{\gamma X(A(\delta_{2}+\mu_{2})+1)-A(\delta_{1}-\mu_{1})+1+\sqrt{D}}{2A},$ (34)
$s_{2}$ $=\frac{s_{1}+\delta_{1}}{\gamma X}-\delta_{2},$ (35)
where the discriminant $D$ is
$\begin{split}D&=(\gamma X(A(\delta_{2}+\mu_{2})+1)-A(\delta_{1}-\mu_{1})+1)^{2}\\ &\quad-4A\left(\gamma X((A\mu_{1}+1)(\delta_{2}+\mu_{2})+\mu_{1}-\mu_{12})-(A\mu_{1}+1)\delta_{1}\right).\end{split}$ (36)
There would be another pair of solutions where the square root in (34) has a negative sign, but that pair is not valid because it does not satisfy $s_{1}>\mu_{1}$, $s_{2}>\mu_{2}$ as required by (29). This is illustrated in Figure 1, which makes it clear that there is only one solution pair in the valid range; and therefore it corresponds to the positive sign.
Figure 1: Solutions $s_{1},s_{2}$ to (20) and (33) (example with $A=0.05$, $\delta_{1}=1$, $\delta_{2}=3$, $X=2.5$, $\mu_{1}=2$, $\mu_{2}=0$, $\mu_{12}=1$)
As indicated in Subsection 2.1, the solutions (34) and (35) have to be rounded, because only integer numbers can be used as IBS parameters. Depending on their specific values it may be necessary to round both of them up, or it may be sufficient to round one up and the other down, if that satisfies (31). A simple criterion, which will be assumed in the rest of the paper, is as follows. First, randomly choose $(i,j)=(1,2)$ or $(2,1)$ with equal probability; then round $s_{i}$ up and $s_{j}$ down, and check if those values satisfy (31). If not, try rounding $s_{i}$ down and $s_{j}$ up. If not valid either, round both values up, which necessarily satisfies (31).
Expressions (34) and (35) give $s_{1}$ and $s_{2}$ as functions of $X$, up to the required rounding, but are difficult to deal with. A natural simplification, which will be helpful in fulfilling (17), is to replace them by the first-order approximations
$s_{1}$ $\approx a_{1}+b_{1}X,$ (37)
$s_{2}$ $\approx a_{2}+\frac{b_{2}}{X},$ (38)
where the coefficients $a_{1}$, $b_{1}$, $a_{2}$, $b_{2}$ are obtained from (34)–(36) as
$a_{1}$ $=\lim_{X\rightarrow 0}s_{1}$ $=1/A+\mu_{1},$ (39)
$b_{1}$ $=\lim_{X\rightarrow\infty}\frac{s_{1}}{X}$ $=\gamma(1/A+\delta_{2}+\mu_{2}),$ (40)
$a_{2}$ $=\lim_{X\rightarrow\infty}s_{2}$ $=1/A+\mu_{2},$ (41)
$b_{2}$ $=\lim_{X\rightarrow 0}s_{2}X$ $=\frac{1/A+\delta_{1}+\mu_{1}}{\gamma}.$ (42)
Figure 2 represents these approximations using the example values (same as in Figure 1) $A=0.05$, $\gamma=0.5$, $\delta_{1}=1$, $\delta_{2}=3$, and with $\mu_{1}$, $\mu_{2}$, $\mu_{12}$ given by (30), as specified for RR. The figure illustrates that the accuracy of the approximations depends on the curvature of $s_{1}$ and $s_{2}$ as functions of $X$ and $1/X$. In fact, for certain values of $A$, $\delta_{1}$ and $\delta_{2}$ the approximations (37) and (38) are exact. This will be analyzed in Subsection 2.3.
(a) $s_{1}$
(b) $s_{2}$
Figure 2: First-order approximations of $s_{1}$ and $s_{2}$ for RR (example with $A=0.05$, $\gamma=0.5$, $\delta_{1}=1$, $\delta_{2}=3$)
Substituting (37) and (38) into (15), and introducing an additional term $\xi$ to account for the effect of rounding $s_{1}$ and $s_{2}$, the average sample sizes are approximated as
$\operatorname{E}[M_{1}+N_{1}]$ $\approx\frac{a_{1}+r_{1}+\xi}{p_{1}}+\frac{b_{1}}{p_{1}}\operatorname{E}[X],$ (43)
$\operatorname{E}[M_{2}+N_{2}]$ $\approx\frac{a_{2}+r_{2}+\xi}{p_{2}}+\frac{b_{2}}{p_{2}}\operatorname{E}\left[\frac{1}{X}\right].$ (44)
The term $\xi$ models the average increase in $s_{1}$ and $s_{2}$ due to rounding. The impact of this on $\operatorname{E}[M_{i}+N_{1}]$, $i=1,2$ will usually be negligible, and thus $\xi$ could be taken as 00 with good approximation. However, choosing $\xi=1$ (together with appropriate values of $r_{1}$ and $r_{2}$) will be useful in Subsection 2.3 to obtain upper bounds on $\operatorname{E}[M_{i}+N_{i}]$.
The terms $\operatorname{E}[X]$ and $\operatorname{E}[1/X]$ in (43) and (44) can be obtained as follows. Since $M_{1}$ and $M_{2}$ in (21) are independent,
$\operatorname{E}[X]=\operatorname{E}[M_{2}-\epsilon_{2}]\operatorname{E}\left[\frac{1}{M_{1}-\epsilon_{1}}\right].$ (45)
From (11),
$\operatorname{E}[M_{2}-\epsilon_{2}]=\frac{r_{2}}{p_{2}}-\epsilon_{2}=\frac{r_{2}}{p_{2}}\left(1-\frac{\epsilon_{2}p_{2}}{r_{2}}\right).$ (46)
As regards $\operatorname{E}[1/(M_{1}-\epsilon_{1})]$, it is easy to see that for $\epsilon_{1}\in(0,1)$
$\frac{1}{M_{1}}<\frac{1}{M_{1}-\epsilon_{1}}<\frac{\epsilon_{1}}{M_{1}-1}+\frac{1-\epsilon_{1}}{M_{1}},$ (47)
where $\operatorname{E}[1/(M_{1}-1)]$ is given by (12) and $\operatorname{E}[1/M_{1}]$ is bounded as [17, section II]
$\frac{p_{1}}{r_{1}-1}\left(1-\frac{p_{1}}{r_{1}-2}\right)<\operatorname{E}\left[\frac{1}{M_{1}}\right]<\frac{p_{1}}{r_{1}-1}\left(1-\frac{p_{1}}{r_{1}-1+p_{1}}\right).$ (48)
Therefore,
$\frac{p_{1}}{r_{1}-1}\left(1-\frac{p_{1}}{r_{1}-2}\right)<\operatorname{E}\left[\frac{1}{M_{1}-\epsilon_{1}}\right]<\frac{p_{1}}{r_{1}-1}\left(1-\frac{(1-\epsilon_{1})p_{1}}{r_{1}-1+p_{1}}\right).$ (49)
It follows from (45), (46) and (49) that $\operatorname{E}[X]$ is bounded for $\epsilon_{1},\epsilon_{2}\in(0,1)$ as
$\left(1-\frac{\epsilon_{2}p_{2}}{r_{2}}\right)\left(1-\frac{p_{1}}{r_{1}-2}\right)<\frac{r_{1}-1}{r_{2}\theta}\operatorname{E}[X]<\left(1-\frac{\epsilon_{2}p_{2}}{r_{2}}\right)\left(1-\frac{(1-\epsilon_{1})p_{1}}{r_{1}-1+p_{1}}\right).$ (50)
By analogous arguments, $\operatorname{E}[1/X]$ satisfies the following bound for $\epsilon_{1},\epsilon_{2}\in(0,1)$:
$\left(1-\frac{\epsilon_{1}p_{1}}{r_{1}}\right)\left(1-\frac{p_{2}}{r_{2}-2}\right)<\frac{(r_{2}-1)\theta}{r_{1}}\operatorname{E}\left[\frac{1}{X}\right]<\left(1-\frac{\epsilon_{1}p_{1}}{r_{1}}\right)\left(1-\frac{(1-\epsilon_{2})p_{2}}{r_{2}-1+p_{2}}\right).$ (51)
A convenient choice for $\epsilon_{1}$ and $\epsilon_{2}$, which will be assumed in the sequel, is $\epsilon_{1}=\epsilon_{2}=1/2$. Then, in view of (50) and (51), $\operatorname{E}[X]$ and $\operatorname{E}[1/X]$ are well approximated, especially for small $p_{1}$, $p_{2}$, as
$\operatorname{E}[X]$ $\approx\frac{r_{2}\theta}{r_{1}-1},$ (52)
$\operatorname{E}\left[\frac{1}{X}\right]$ $\approx\frac{r_{1}}{(r_{2}-1)\theta}.$ (53)
Substituting (52) and (53) into (43) and (44) yields
$\operatorname{E}[M_{1}+N_{1}]$ $\approx\frac{a_{1}+r_{1}+\xi}{p_{1}}+\frac{b_{1}r_{2}}{(r_{1}-1)p_{2}},$ (54)
$\operatorname{E}[M_{2}+N_{2}]$ $\approx\frac{a_{2}+r_{2}+\xi}{p_{2}}+\frac{b_{2}r_{1}}{(r_{2}-1)p_{1}}.$ (55)
By means of (54) and (55), a desired ratio of average sample sizes can be approximately achieved. Specifically, (17) will hold for all $p_{1}$, $p_{2}$ if
$a_{1}+r_{1}+\xi$ $=\frac{\lambda b_{2}r_{1}}{r_{2}-1},$ (56)
$\lambda(a_{2}+r_{2}+\xi)$ $=\frac{b_{1}r_{2}}{r_{1}-1}.$ (57)
This equation system can be expressed, making use of (39)–(42), as
$\frac{1}{A}+r_{1}+\mu_{1}+\xi$ $=\frac{\lambda r_{1}}{\gamma(r_{2}-1)}\left(\frac{1}{A}+\delta_{1}+\mu_{1}\right),$ (58)
$\frac{\lambda}{\gamma}\left(\frac{1}{A}+r_{2}+\mu_{2}+\xi\right)$ $=\frac{r_{2}}{r_{1}-1}\left(\frac{1}{A}+\delta_{2}+\mu_{2}\right).$ (59)
Given $A$, $\lambda$, $\xi$, and with $\mu_{1}$, $\mu_{2}$, $\mu_{12}$ known, the design parameters $r_{1}$, $r_{2}$, $\delta_{1}$, $\delta_{2}$, $\gamma$ provide degrees of freedom that help meet (58) and (59). Namely, imposing the relationships
$r_{2}=r_{1}+\mu_{1}-\mu_{2},\qquad\delta_{2}=\delta_{1}+\mu_{1}-\mu_{2},$ (60)
it is straightforward to solve for $\gamma$ and $\delta_{1}$ in (58) and (59):
$\gamma$ $=\lambda\sqrt{\frac{r_{1}(r_{1}-1)}{r_{2}(r_{2}-1)}}=\lambda\sqrt{\frac{r_{1}(r_{1}-1)}{(r_{1}+\mu_{1}-\mu_{2})(r_{1}+\mu_{1}-\mu_{2}-1)}},$ (61)
$\begin{split}\delta_{1}&=\left(\frac{1}{A}+r_{1}+\mu_{1}+\xi\right)\frac{\lambda(r_{1}-1)}{\gamma\,(r_{1}+\mu_{1}-\mu_{2})}-\frac{1}{A}-\mu_{1}\\ &=\left(\frac{1}{A}+r_{1}+\mu_{1}+\xi\right)\sqrt{\frac{(r_{1}-1)(r_{1}+\mu_{1}-\mu_{2}-1)}{r_{1}(r_{1}+\mu_{1}-\mu_{2})}}-\frac{1}{A}-\mu_{1}.\end{split}$ (62)
The parameters $\mu_{1}$, $\mu_{2}$ $\mu_{12}$ for the RR estimator, given by (30), satisfy
$0\leq\mu_{1},\mu_{2}\leq 2,\quad 0\leq\mu_{12},$ (63)
and this will also be the case for the other estimators to be described later. Combining (63) with (23), (60) and (62), it follows that $\delta_{i}>-1/A-\mu_{i}$, $i=1,2$. On the other hand, (29) and (31) imply that $s_{i}-\mu_{i}>1/A$. From these two inequalities it stems that $\delta_{1}$ and $\delta_{2}$ as computed from (60) and (62) fulfill the requirement (22), and are thus valid. Making use of (60), the expressions (34)–(36) for $s_{1}$, $s_{2}$ are more conveniently written as
$s_{1}$ $=\frac{\gamma X(A(\delta_{1}+\mu_{1})+1)-A(\delta_{1}-\mu_{1})+1+\sqrt{D}}{2A},$ (64)
$s_{2}$ $=\frac{s_{1}+\delta_{1}}{\gamma X}-\delta_{1}-\mu_{1}+\mu_{2},$ (65)
$\begin{split}D&=(\gamma X(A(\delta_{1}+\mu_{1})+1)-A(\delta_{1}-\mu_{1})+1)^{2}\\ &\quad-4A\left(\gamma X((A\mu_{1}+1)(\delta_{1}+\mu_{1})+\mu_{1}-\mu_{12})-(A\mu_{1}+1)\delta_{1}\right).\end{split}$ (66)
Substituting (39)–(42) and (60)–(62) into (54) and (55) yields
$\operatorname{E}[M_{1}+N_{1}]$ $\approx\left(\frac{1}{A}+r_{1}+\mu_{1}+\xi\right)\left(\frac{1}{p_{1}}+\frac{\lambda}{p_{2}}\right),$ (67)
$\operatorname{E}[M_{2}+N_{2}]$ $\approx\left(\frac{1}{A}+r_{1}+\mu_{1}+\xi\right)\left(\frac{1}{\lambda p_{1}}+\frac{1}{p_{2}}\right).$ (68)
It is convenient to use a normalized version of the average sample sizes, $\operatorname{E}[M_{i}+N_{i}]\sqrt{p_{1}p_{2}}$, so that the normalized average sample sizes depend on the two probabilities only through their ratio $\theta$. Defining
$\phi=\sqrt{p_{1}p_{2}},$ (69)
the approximations (67) and (68) can be written as
$\operatorname{E}[M_{1}+N_{1}]\phi$ $\approx\left(\frac{1}{A}+r_{1}+\mu_{1}+\xi\right)\left(\frac{1}{\sqrt{\lambda\theta}}+\sqrt{\lambda\theta}\right)\sqrt{\lambda},$ (70)
$\operatorname{E}[M_{2}+N_{2}]\phi$ $\approx\left(\frac{1}{A}+r_{1}+\mu_{1}+\xi\right)\left(\frac{1}{\sqrt{\lambda\theta}}+\sqrt{\lambda\theta}\right)\frac{1}{\sqrt{\lambda}}.$ (71)
The average number of samples $\operatorname{E}[M_{i}+N_{i}]$, $i=1,2$ is the sum of two terms inversely proportional to $p_{1}$ and $p_{2}$, according to (67) and (68); and it is $\lambda$ times more sensitive to $p_{2}$ than to $p_{1}$. That is, the parameter of the population for which a smaller average sample size is desired (as specified by (17)) has a stronger influence on both average sample sizes. It is also noteworthy that, for $\lambda$ fixed, (70) and (71) are minimized when $\lambda\theta=1$, which according to (17) and (18) corresponds to $r_{1}+s_{1}\approx r_{2}+s_{2}$. Thus, disparity in the sum of IBS parameters between the two populations incurs an increase in average sample sizes.
2.3 Analysis of the approximation. Bounds on average sample sizes
Based on an analysis of the approximation error in (37) and (38), this subsection obtains bounds on the average sample sizes and selects appropriate values for $r_{1}$ and $r_{2}$.
The error in approximating (34) and (35) by (37) and (38) vanishes when $s_{1}$ is an affine function of $X$ and $s_{2}$ is an affine function of $1/X$ (see Figure 2). This happens when $A$, $r_{1}$ and $\xi$ satisfy a certain relationship, as discussed next. Expanding (66) and collecting terms, $D$ can be written as $d_{2}X^{2}+d_{1}X+d_{0}$ with
$d_{2}$ $=\gamma^{2}(A(\delta_{1}+\mu_{1})+1)^{2},$ (72)
$d_{1}$ $=2\gamma(-A^{2}(\delta_{1}+\mu_{1})^{2}-2A(\delta_{1}+\mu_{1}-\mu_{12})+1),$ (73)
$d_{0}$ $=(A(\delta_{1}+\mu_{1})+1)^{2}.$ (74)
The condition that $s_{1}$ and $s_{2}$ are affine functions of $X$ and $1/X$ respectively is equivalent to $D$ being a perfect square with respect to $X$, that is,
$d_{1}=\pm 2\sqrt{d_{2}d_{0}}.$ (75)
The equality (75) with negative sign has the solution $A=-1/\mu_{12}$ if $\mu_{12}\neq 0$. With the estimators considered in this paper, $\mu_{12}$ is either 00 or positive, and this results in no solution or a negative solution for $A$, which is not valid. On the other hand, (75) with positive sign is satisfied if and only if $A(\delta_{1}+\mu_{1})^{2}+2(\delta_{1}+\mu_{1})-\mu_{12}=0.$ Taking into account (62), this can be expressed as $c(A,r_{1},\xi)=0$ with
$\begin{split}c(A,r_{1},\xi)&=A\left(\left(\frac{1}{A}+r_{1}+\mu_{1}+\xi\right)\sqrt{\frac{(r_{1}-1)(r_{1}+\mu_{1}-\mu_{2}-1)}{r_{1}(r_{1}+\mu_{1}-\mu_{2})}}-\frac{1}{A}\right)^{2}\\ &\quad+2\left(\left(\frac{1}{A}+r_{1}+\mu_{1}+\xi\right)\sqrt{\frac{(r_{1}-1)(r_{1}+\mu_{1}-\mu_{2}-1)}{r_{1}(r_{1}+\mu_{1}-\mu_{2})}}-\frac{1}{A}\right)-\mu_{12}.\end{split}$ (76)
Furthermore, it is easy to see that $c(A,r_{1},\xi)$ being positive (negative) implies that $s_{1}$ and $s_{2}$ given by (64)–(66) are both convex (concave) functions of $X$ and $1/X$ respectively, which ensures that they are smaller (greater) than their approximations. Thus $c(A,r_{1},\xi)$ will be referred to as curvature function.
This characterization of the error allows transforming (70) and (71) into upper bounds. Since rounding $s_{1}$ and $s_{2}$ never increases their values by $1$ or more, taking $\xi=1$ and choosing $r_{1}$ such that $c(A,r_{1},\xi)\geq 0$ implies that approximations (43) and (44) become inequalities:
$\operatorname{E}[M_{1}+N_{1}]$ $<\frac{a_{1}+r_{1}+1}{p_{1}}+\frac{b_{1}}{p_{1}}\operatorname{E}[X],$ (77)
$\operatorname{E}[M_{2}+N_{2}]$ $<\frac{a_{2}+r_{2}+1}{p_{2}}+\frac{b_{2}}{p_{2}}\operatorname{E}\left[\frac{1}{X}\right].$ (78)
From (50) and (51) it is clear that $E[X]<r_{2}\theta/(r_{1}-1)$ and $\operatorname{E}[1/X]<r_{1}/((r_{2}-1)\theta)$. Substituting into (77) and (78) and using (39)–(42) and (60)–(62) yields
$\operatorname{E}[M_{1}+N_{1}]\phi$ $<\left(\frac{1}{A}+r_{1}+\mu_{1}+1\right)\left(\frac{1}{\sqrt{\lambda\theta}}+\sqrt{\lambda\theta}\right)\sqrt{\lambda},$ (79)
$\operatorname{E}[M_{2}+N_{2}]\phi$ $<\left(\frac{1}{A}+r_{1}+\mu_{1}+1\right)\left(\frac{1}{\sqrt{\lambda\theta}}+\sqrt{\lambda\theta}\right)\frac{1}{\sqrt{\lambda}}.$ (80)
These bounds hold for $\xi=1$ and for any $r_{1}$ such that $c(A,r_{1},1)\geq 0$. Considering the restriction (23), the best (i.e. lowest) upper bounds are obtained by choosing $r_{1}$ as
$r_{1}=\min\{r=3,4,5,\ldots\mid c(A,r,1)\geq 0\}.$ (81)
Thus the values $\xi=1$ and $r_{1}$ as in (81) will be used for the estimator.
Figure 3: Pairs $(A,r_{1})$ for which $c(A,r_{1},1)=0$
The condition $c(A,r_{1},\xi)=0$, with $c(A,r_{1},\xi)$ given by (76), can be written as
$\begin{split}&(r_{1}+\mu_{1}+\xi)^{2}(r_{1}-1)(r_{1}+\mu_{1}-\mu_{2}-1)A^{2}\\ &\quad+\left(2(r_{1}+\mu_{1}+\xi)(r_{1}-1)(r_{1}+\mu_{1}-\mu_{2}-1)-\mu_{12}r_{1}(r_{1}+\mu_{1}-\mu_{2})\right)A\\ &\quad+1-2r_{1}-\mu_{1}+\mu_{2}=0.\end{split}$ (82)
Considering $r_{1}$ as given, this is a quadratic equation in $A$ with a single positive solution. Figure 3 shows the resulting curve for RR, i.e. for $\mu_{1}$, $\mu_{2}$, $\mu_{12}$ as in (30), with $\xi=1$. (The figure also contains curves for other estimators, to be presented in following sections.) For the pairs $(A,r_{1})$ in this curve the approximations (37) and (38) are exact. Note that $r_{1}$ is considered as a continuous variable for clarity of the representation, but only the points with integer $r_{1}$ (marked with dots in the graph) are feasible. Furthermore, it is seen from (76) that $c(A,r_{1},1)$ increases with $r_{1}$, and therefore the region above (below) the curve corresponds to $c(A,r_{1},1)$ positive (negative). Thus, for a given $A$, once the value of $r_{1}$ for which $c(A,r_{1},1)=0$ is known, rounding this up gives the minimum integer $r_{1}$ such that $c(A,r_{1},1)\geq 0$. That is, (81) is the integer on or immediately above the curve in Figure 3, limited from below by $3$.
With the values selected for $\gamma$, $\delta_{1}$, $\delta_{2}$, $\epsilon_{1}$, $\epsilon_{2}$, $\xi$, $r_{1}$, $r_{2}$, the RR estimator is completely specified. The estimation procedure is summarized in Algorithm 1, together with a list of its properties, some of which will be derived in the remainder of this section. (The algorithm also includes the LRR case, to be presented in Section 3.)
Algorithm 1 Estimator of RR or LRR
  1. Input parameters:
    Target $A$, interpreted as relative MSE for RR, or MSE for LRR;
    Desired ratio of average sample sizes $\lambda$, or group sizes $l_{1}$, $l_{2}$.
  2. Estimation procedure:
  3. 1.
    Define $\mu_{1},\mu_{2},\mu_{12}$ as in (30) for RR, or as in (119) for LRR.
    Define $\epsilon_{1}=\epsilon_{2}=1/2$ and $\xi=1$.
    Define $\lambda=l_{1}/l_{2}$ if group sampling is applied.
  4. 2.
    Compute $r_{1}$ from (81) with $c(A,r_{1},\xi)$ given by (76) (for LRR the explicit expression (122) can equivalently be used), or obtain it by rounding up the value from Figure 3, with a minimum of $3$. Compute $r_{2}$ as $r_{1}+\mu_{1}-\mu_{2}$.
  5. 3.
    Compute $\gamma$ and $\delta_{1}$ from (61) and (62).
  6. 4.
    (First sampling stage): For each $i=1,2$, repeatedly sample from population $i$ until $r_{i}$ successes are obtained. Let $M_{i}$ be the number of samples.
  7. 5.
    Compute $X$ from (21). Compute $s_{1}$, $s_{2}$ from (64)–(66), and then round one of them up and the other down, or both up, to fulfill (31).
  8. 6.
    (Second sampling stage): For each $i=1,2$, repeatedly sample from population $i$ until $s_{i}$ successes are obtained. Let $N_{i}$ be the number of samples.
  9. 7.
    Compute $\hat{\theta}$ from (24), or $\hat{\Theta}$ from (117).
  10. Output:
    Estimation $\hat{\theta}$ or $\hat{\Theta}$.
  11. Properties:
    Unbiased, with relative MSE (RR) or MSE (LRR) less than $A$;
    Average sample sizes: (79), (80); approximate ratio $\lambda$;
    Efficiency with separate sampling: (92); approaches $1$ for $A$, $\phi$ small
    Average number of groups: (101), (112);
    Efficiency with group sampling: (115).
The performance of the RR estimator is evaluated by means of Monte Carlo simulations in the following. For each combination of parameters a simulation is run consisting of $10^{6}$ realizations of the estimator. The empirical MSE and average numbers of samples are computed from the simulation (i.e. expectation is replaced by sample mean), and then they are compared with the theoretical expressions.
Figure 4 shows the bounds (79) and (80) for the normalized average numbers of samples with $r_{1}$ given by (81) (lines), as well as simulation results for the same $r_{1}$ (bold, red dots) and for $r_{1}=3,\ldots,50$ (light, grey dots). The graphs consider several combinations of $\lambda$, $\theta$ and $\phi$, with variable $A$. As a reference, (81) gives $3,\ldots,31$ for the displayed range of $A$.
(a) $\lambda=1$, $\theta=1$, $\phi=0.01$ ($p_{1}=p_{2}=0.01$)
(b) $\lambda=1$, $\theta=16$, $\phi=0.01$ ($p_{1}=0.04$, $p_{2}=0.0025$)
(c) $\lambda=3$, $\theta=16$, $\phi=0.01$ ($p_{1}=0.04$, $p_{2}=0.0025$)
Line: bound. Bold dots: simulation. Light dots: simulation, other values of $r_{1}$.
Figure 4: Normalized average sample sizes for RR with separate sampling, varying $r_{1}$
(a) $\lambda=3$, $\theta=1/16$, $\phi=0.01$ ($p_{1}=0.0025$, $p_{2}=0.04$)
(b) $\lambda=1/3$, $\theta=1/16$, $\phi=0.01$ ($p_{1}=0.0025$, $p_{2}=0.04$)
(c) $\lambda=1/3$, $\theta=1/16$, $\phi=0.1$ ($p_{1}=0.025$, $p_{2}=0.4$)
Line: bound. Bold dots: simulation. Light dots: simulation, other values of $r_{1}$.
Figure 5: Normalized average sample sizes for RR with separate sampling, varying $r_{1}$
The following observations can be made from Figure 4. For $r_{1}$ as in (81), simulated values are close to the theoretical bounds (and always below). When $A$ is not too large they are in fact very close, whereas for large $A$ the difference increases because $r_{1}$ is limited below by $3$. In addition, comparing with simulations for other values of $r_{1}$, (81) achieves average sample sizes equal or very close to their minima with respect to $r_{1}$. For $r_{1}$ given by (81), the ratio $\operatorname{E}[M_{1}+N_{1}]/\operatorname{E}[M_{2}+N_{2}]$ in the simulations is very close to $\lambda$ in all cases (for reference, note that the theoretical curves are in the exact ratio $\lambda$); this will be analyzed with more detail later. The effect of $\phi$ on the normalized average sample sizes $\operatorname{E}[M_{i}+N_{i}]\phi$, $i=1,2$ is imperceptible (compare Figures 5(b) and 5(c)). The parameter $\theta$ has an impact on the average sample sizes, but not on their ratio, which remains close to $\lambda$ (compare Figures 4(a) and 4(b), or 4(c) and 5(a)). If both $\theta$ and $\lambda$ are replaced by their reciprocal values, $\operatorname{E}[M_{1}+N_{1}]$ and $\operatorname{E}[M_{2}+N_{2}]$ are simply swapped (see Figures 4(c) and 5(b)). This is in agreement with (70) and (71). Lastly, for $A$ small the average sample sizes are approximately inversely proportional to this parameter. Again, this can be observed in (70) and (71), where for $A$ small the term $1/A$ dominates the other summands in the first factor.
The theoretical curves in Figure 4 have small jump discontinuities (see for example Figure 4(c) near $A=0.07$), caused by the discrete character of $r_{1}$ and $r_{2}$. The jumps occur when the result of (81) changes by $1$. This effect is also present in the simulation results (bold dots), although less evident. Apart from this, for fixed $r_{1}$ the simulated results exhibit steep changes in certain locations, produced by the rounding applied to $s_{1}$ and $s_{2}$ (see for example the rightmost region for $\operatorname{E}[M_{2}+N_{2}]$ in Figure 4(c) with $r_{1}=3$, which in that region corresponds to the bold dots). These are only observable in the simulation results, because the theoretical curves do not explicitly model the rounding of $s_{1}$ and $s_{2}$; and they are not discontinuities, but short sections where the slope is large in absolute value, as will be discussed later.
Figure 6 compares the target $A$ with the relative MSE, $\operatorname{E}[(\hat{\theta}-\theta)^{2}]/\theta^{2}$, obtained from simulations in two specific cases, for $r_{1}$ given by (81). As seen, the relative MSE is always less than $A$, in accordance with (16). The difference between simulation and target is considerable when the latter is large. Again, this is a consequence of the limitation of (81) to values not smaller than $3$. In addition, the difference increases slightly with $\phi$ (compare Figures 6(a) and 6(b)). This is, at least in part, explained by the fact that the uniform bound (28) becomes less tight as $p_{1}$ or $p_{2}$ approach $1$, as can be seen comparing it with (27).
(a) $\lambda=1$, $\theta=16$, $\phi=0.1$
($p_{1}=0.025$, $p_{2}=0.4$)
(b) $\lambda=3$, $\theta=1/16$, $\phi=0.01$
($p_{1}=0.04$, $p_{2}=0.0025$)
Dashed line: target. Dots: simulation.
Figure 6: Relative MSE for RR with separate sampling
The relative MSE, like the average sample sizes, has steep changes near certain values of $A$, caused by rounding $s_{1}$ and $s_{2}$. The effect is noticeable in the rightmost region of Figure 6, and is explained as follows. For large $A$, at least one of the two parameters $s_{1}$ and $s_{2}$ is small, according to (29) and (33), so as to make the relative MSE similar to (but smaller than) the target. Consequently that parameter, $s_{j}$, has a very narrow distribution before rounding, which implies that rounding it almost always produces the same integer value. This will typically be the next greater integer, because for small $s_{j}$ rounding causes a large variation in the error function, and thus rounding down is not likely to satisfy (31). In these conditions, if $A$ is increased by a small amount, the change in the distribution of $s_{j}$ before rounding, even if correspondingly small, can be sufficient to cause the rounded value to decrease by $1$, producing a significant change in $\operatorname{E}[(\hat{\theta}-\theta)^{2}]/\theta^{2}$, as well as in $\operatorname{E}[M_{j}+N_{j}]$.
As an specific example, for $\lambda=1$, $\theta=16$ (Figure 6(a)), consider $A=0.562$ (leftmost dot in the highest “plateau” of simulated values), for which (81) gives $r_{1}=3$. The simulation yields an average value of $1.856$ for $s_{2}$ before rounding, and a standard deviation of $0.082$. The value after rounding is $2$ with probability $0.962$ (and greater with probability $0.038$). The parameter $s_{1}$ before rounding takes larger values, with average $109.2$ and standard deviation $76.6$. Reducing $A$ to $0.501$ (next simulated value towards the left), the average of $s_{2}$ before rounding becomes $2.087$, and the standard deviation is $0.094$. After rounding, $s_{2}$ now equals $3$ with probability $0.999$ (and is greater with probability $0.001$). Again, $s_{1}$ before rounding takes larger values, with mean and standard deviation $112.5$ and $77.9$ respectively. The change of the rounded value of $s_{2}$, from being $2$ to being $3$ with probabilities near $1$, causes the rightmost vertical gap in Figure 6(a), from a relative MSE approximately equal to $0.50$ down to $0.34$. In contrast, for small $A$ both $s_{1}$ and $s_{2}$ are large and the effect of rounding is less marked, because, on one one hand, the distribution of each parameter is wider, with many possible rounded values; and, on the other hand, rounding a large value only causes a small variation in the relative MSE.
The same behavior can be seen in the average sample sizes. Thus, in Figure 4(b), which also corresponds to $\lambda=1$, $\theta=16$, the graph of $\operatorname{E}[M_{2}+N_{2}]$ has a vertical gap at the same horizontal position discussed in the preceding paragraph for the relative MSE. The effect is not discernible in the graph of $\operatorname{E}[M_{1}+N_{1}]$ because $s_{1}$ is large.
It follows from the above that the effect of rounding $s_{1}$ and $s_{2}$ is not a discontinuity, but rather a short section with large slope in $\operatorname{E}[(\hat{\theta}-\theta)^{2}]/\theta^{2}$, or in $\operatorname{E}[M_{i}+N_{i}]$, as a function of $A$. Indeed, since (64) and (65) are continuous functions of $A$, the distributions of $s_{1}$ and $s_{2}$ also vary continuously with $A$, and so do the relative MSE and average sample sizes. This can be seen in Figure 7(a), which is a detailed view of the rightmost part of Figure 6(a) with smaller spacing along the horizontal axis, and in Figure 7(b), which shows the corresponding values of $\operatorname{E}[M_{2}+N_{2}]$.
(a) Relative MSE
(b) $\operatorname{E}[M_{2}+N_{2}]\phi$
Dashed line: target. Solid line: approximation. Dots: simulation.
Figure 7: Detail of relative MSE and normalized average sample size for RR with separate sampling, for large $A$; $\lambda=1$, $\theta=16$, $\phi=0.1$
The ratio $\operatorname{E}[M_{1}+N_{1}]/\operatorname{E}[M_{2}+N_{2}]$ obtained from simulation is represented in Figure 8. As can be seen, it is in general close to $\lambda$, and very close for small or moderate values of $A$. The relatively larger deviations in the rightmost region of the graphs can again be attributed to the fact that for large $A$, since $r_{1}$ is limited from below by $3$ in (81), the value of $r_{1}$ does not result in $c(A,r_{1},\xi)\approx 0$. In any case, the deviation between the actual ratio and $\lambda$ is less than $11\%$ for all combinations of parameters shown in the figure, and much lower for $A$ small.
(a) $\phi=0.01$
(b) $\phi=0.1$
$\lambda=1/3$, $1$, $3$ (bottom to top). $\theta=1/16$, $1$, $16$ (see legend).
Figure 8: Ratio of average sample sizes for RR with separate sampling
2.4 Efficiency with separate sampling
The efficiency of an unbiased estimator can be defined by comparing its variance against the minimum variance that can be achieved by any unbiased estimator with the same sample size, or with the same average sample size. For a fixed-size estimator based on independent observations from a single population, the minimum possible variance is given by the Cramér–Rao bound [11, 13]. This bound applies under certain regularity conditions on the distribution of the samples, which are satisfied when the observations are binary. For sequential estimators, there exists a generalization of the Cramér–Rao bound, obtained by Wolfowitz [22] see also 11, section 4.3. The actual bound is the same, except that the average sample size is used instead of a fixed sample size, and the regularity conditions are different.
For estimators based on independent observations from two populations, as considered in this paper, a variation of the Cramér–Rao bound can be applied in the fixed-size case [13, chapter 3]. However, to the author’s knowledge, there is no analogue of Wolfowitz’s result for sequential estimators when the observations are obtained from more than one population. Therefore, the efficiency of the RR estimator (and of the estimators to be presented in subsequent sections) will be defined by comparing its variance with the lowest variance that can be attained by any fixed-size estimator with the same average size for each population.
Consider fixed numbers $n_{1}$ and $n_{2}$ of independent binary samples taken from two populations with parameters $p_{1}$ and $p_{2}$. The number of successes observed from population $i$ follows a binomial distribution with parameters $n_{i}$ and $p_{i}$. Thus
$L(S_{1},S_{2};p_{1},p_{2})=\binom{n_{1}}{S_{1}}\binom{n_{2}}{S_{2}}p_{1}^{S_{1}}(1-p_{1})^{n_{1}-S_{1}}p_{2}^{S_{2}}(1-p_{2})^{n_{2}-S_{2}}$ (83)
is the probability of observing $S_{1}$ and $S_{2}$ successes from the two populations respectively. For a generic parameter $\zeta$ that is a function of $p_{1},p_{2}$, and an unbiased estimator $\hat{\zeta}$ of $\zeta$, the Cramér–Rao bound is [13, section 3.8]
$\operatorname{Var}[\hat{\zeta}]\geq\mathbf{J}\>\mathbf{F}^{-1}\,\mathbf{J}^{\intercal}$ (84)
where $\mathbf{J}$ is the $1\times 2$ Jacobian vector
$\mathbf{J}=\begin{bmatrix}\partial\zeta/\partial p_{1}&\partial\zeta/\partial p_{2}\end{bmatrix}$ (85)
and $\mathbf{F}$ is the $2\times 2$ Fisher information matrix, defined as
$F_{i,j}=-\operatorname{E}\left[\frac{\partial^{2}\log L(S_{1},S_{2};p_{1},p_{2})}{\partial p_{i}\partial p_{j}}\right],\quad i,j=1,2.$ (86)
This matrix is readily computed from (83) as
$F_{i,i}$ $=\frac{n_{i}}{p_{i}(1-p_{i})},$ (87)
$F_{i,j}$ $=0\qquad\qquad\text{ for }i\neq j$ (88)
(where (88) reflects the fact that observations from one population give no information about the other), and thus (84) becomes
$\operatorname{Var}[\hat{\zeta}]\geq\left(\frac{\partial\zeta}{\partial p_{1}}\right)^{2}\frac{p_{1}(1-p_{1})}{n_{1}}+\left(\frac{\partial\zeta}{\partial p_{2}}\right)^{2}\frac{p_{2}(1-p_{2})}{n_{2}}.$ (89)
Equating $n_{i}$ to the average number of observations of population $i$ used by the considered estimator, $\operatorname{E}[M_{i}+N_{i}]$, the efficiency with separate sampling $\eta_{\mathrm{sep}}$ is obtained as
$\eta_{\mathrm{sep}}=\frac{\displaystyle\left(\frac{\partial\zeta}{\partial p_{1}}\right)^{2}\frac{p_{1}(1-p_{1})}{\operatorname{E}[M_{1}+N_{1}]}+\left(\frac{\partial\zeta}{\partial p_{2}}\right)^{2}\frac{p_{2}(1-p_{2})}{\operatorname{E}[M_{2}+N_{2}]}}{\operatorname{Var}[\hat{\zeta}]}.$ (90)
Particularizing (90) for RR, that is $\zeta=\theta$, with $\partial\theta/\partial p_{1}=\theta/p_{1}$, $\partial\theta/\partial p_{2}=-\theta/p_{2}$,
$\eta_{\mathrm{sep}}=\frac{\displaystyle\frac{1-p_{1}}{\operatorname{E}[M_{1}+N_{1}]p_{1}}+\frac{1-p_{2}}{\operatorname{E}[M_{2}+N_{2}]p_{2}}}{\operatorname{Var}[\hat{\theta}]/\theta^{2}}.$ (91)
For $\xi=1$ and $r_{1}$ as in (81), substituting (16), (79) and (80) into (91) yields the following bound for the efficiency of the RR estimator with separate sampling:
$\begin{split}\eta_{\mathrm{sep}}>\frac{1}{1+A(r_{1}+\mu_{1}+1)}\left(1-\phi\frac{1/\sqrt{\lambda}+\sqrt{\lambda}}{1/\sqrt{\lambda\theta}+\sqrt{\lambda\theta}}\right),\end{split}$ (92)
where $\mu_{1}=2$.
Figure 9 shows the efficiency $\eta_{\mathrm{sep}}$ obtained from Monte Carlo simulation. Its value is computed using (91) with $\operatorname{E}[M_{i}+N_{i}]$, $i=1,2$ replaced by the corresponding sample means and $\operatorname{Var}[\hat{\theta}]$ replaced by the sample MSE. The bound (92) is also plotted. As seen in the figure, the efficiency is high for the values of the target $A$ commonly used in practice, and in particular it is close to $1$ for small $A$. For example, $A=0.04$, corresponding to a relative RMSE of $20\%$, gives values of $\eta_{\mathrm{sep}}$ around $80\%$. The efficiency as a function of $A$ exhibits the same discontinuities and steep changes that have been identified for the relative MSE and for $\operatorname{E}[M_{i}+N_{i}]$. The simulation results deviate more from the theoretical bound for large $A$ (rightmost part of the graphs) or for large $\phi$ (Figure 9(b)). This is explained by the fact that in these conditions the relative MSE is considerably smaller than the target, and the average sample sizes are also considerably smaller than their bounds, as discussed. On the other hand, the bound is quite tight for small or moderate values of $A$ and $\phi$, which is precisely when it is most important to characterize $\eta_{\mathrm{sep}}$ accurately, as sample sizes are large in that case.
It is easily seen that $\lim_{A\rightarrow 0}Ar_{1}^{2}=1$ for $A$ and $r_{1}$ related by (82), and thus also for $r_{1}$ obtained from (81). Combining this result with (92), it follows that $\eta_{\mathrm{sep}}$ tends to $1$ when $A$ and $\phi$ tend to 00. This is in consonance with the values shown in Figure 9.
(a) $\lambda=1$, $\theta=1$, $\phi=0.01$
($p_{1}=p_{2}=0.01$)
(b) $\lambda=1$, $\theta=16$, $\phi=0.1$
($p_{1}=0.4$, $p_{2}=0.025$)
(c) $\lambda=3$, $\theta=1$, $\phi=0.01$
($p_{1}=p_{2}=0.01$)
(d) $\lambda=3$, $\theta=1/16$, $\phi=0.01$
($p_{1}=0.0025$, $p_{2}=0.04$)
Line: bound. Dots: simulation.
Figure 9: Efficiency for RR with separate sampling
2.5 Group sampling. Estimation procedure
In group sampling, samples are taken in groups or batches of $l_{1}$ and $l_{2}$ samples from the two populations respectively. In consequence, the numbers of samples from both populations must be the same integer multiple of $l_{1}$ and $l_{2}$, i.e. $Gl_{1}$ and $Gl_{2}$, where $G$ is the number of groups.
Let $\lambda$ be defined for group sampling as
$\lambda=\frac{l_{1}}{l_{2}}.$ (93)
This definition is consistent with that used for separate sampling, because $\lambda$ still represents the ratio of average sample sizes. In addition, for group sampling $M_{i}$, $N_{i}$, $i=1,2$ will continue to refer to the numbers of samples that would result if the estimation procedure discussed thus far, using separate sampling, were applied with $\lambda$ given by (93) (and therefore these variables do not correspond to the numbers of samples actually required with group sampling).
The estimation process with group sampling is described next. Samples are used individually, following the same procedure as with separate sampling; but are in fact taken in groups. As individual samples from either population become necessary, groups of samples are taken, each group providing $l_{1}$ and $l_{2}$ samples of the two populations. During this process, a group may provide more samples than needed for one or the two populations. In that case the “surplus” samples are stored for later use. When a sample from a given population is required, a new group is taken only if the surplus samples from that population have been exhausted. Once the procedure has finished, there may remain some stored samples, which will be discarded (and at least for one of the two populations, with index $j$, the number of discarded samples will be less than $l_{j}$).
Estimation with group sampling thus proceeds as in Algorithm 1, but with an added “outer layer” that translates group sampling into separate sampling as described. With the above definitions of $M_{i}$ and $N_{i}$, it follows that the number of required groups is
$G=\max\left\{\left\lceil\frac{M_{1}+N_{1}}{l_{1}}\right\rceil,\left\lceil\frac{M_{2}+N_{2}}{l_{2}}\right\rceil\right\}.$ (94)
2.6 Average number of groups
The relevant measure of sample size with group sampling is the average number of groups, $\operatorname{E}[G]$. To make its characterization more tractable, it is helpful to approximate (94) by assuming $\phi$ small, which implies that $\operatorname{E}[M_{i}+N_{i}]$, $i=1,2$ are correspondingly large. In addition to simplifying computations, this is the most important case in practice, as argued previously. For large $M_{i}$ and $N_{i}$, the rounding operations in (94) can be removed with negligible error. Thus, defining
$\Delta=\frac{M_{1}+N_{1}}{l_{1}}-\frac{M_{2}+N_{2}}{l_{2}},$ (95)
and noting that $\max\{x,y\}=(x+y)/2+|x-y|/$2, $\operatorname{E}[G]$ can be expressed as
$\operatorname{E}[G]\approx\operatorname{E}\left[\max\left\{\frac{M_{1}+N_{1}}{l_{1}},\frac{M_{2}+N_{2}}{l_{2}}\right\}\right]=\frac{\operatorname{E}[M_{1}+N_{1}]}{2l_{1}}+\frac{\operatorname{E}[M_{2}+N_{2}]}{2l_{2}}+\frac{\operatorname{E}\left[|\Delta|\right]}{2}.$ (96)
The terms $\operatorname{E}[M_{1}+N_{1}]$ and $\operatorname{E}[M_{2}+N_{2}]$ in (96) are approximately given by (70) and (71). Regarding $\operatorname{E}[|\Delta|]$, it is shown in Appendix A.1 that, for $\phi$ small,
$\operatorname{E}\left[|\Delta|\right]\approx\operatorname{E}\left[\left|\frac{r_{1}+s_{1}}{l_{1}p_{1}}-\frac{r_{2}+s_{2}}{l_{2}p_{2}}\right|\right].$ (97)
Define
$\tilde{\Delta}=\frac{r_{1}+s_{1}}{l_{1}p_{1}}-\frac{r_{2}+s_{2}}{l_{2}p_{2}}.$ (98)
Using approximations (37) and (38) for $s_{1},s_{2}$ and including the rounding term $\xi$ as in Subsection 2.2 (this has a very small effect on the approximations obtained here, but it is done for consistency), (98) becomes
$\tilde{\Delta}\approx\frac{1}{\phi}\left(\frac{a_{1}+r_{1}+\xi+b_{1}X}{l_{1}\sqrt{\theta}}-\frac{a_{2}+r_{2}+\xi+b_{2}/X}{l_{2}}\sqrt{\theta}\right).$ (99)
Let $X_{0}$ denote the value of $X$ for which the right-hand side of (99) equals 00, and let $Y=X/\theta$. Taking into account (93), $X_{0}$ is seen to be the only positive solution of
$b_{1}X_{0}^{2}+(a_{1}+r_{1}+\xi-\lambda\theta(a_{2}+r_{2}+\xi))X_{0}-\lambda\theta b_{2}=0.$ (100)
Making use of (39)–(42) and (60)–(62), this solution is obtained as $X_{0}=\theta Y_{0}$ with
[…]
As established in Appendix A.1, for $\phi$ small the variable $Y$ approximately follows a beta prime distribution with parameters $r_{2}$, $r_{1}$. Thus, from (97)–(99),
$\operatorname{E}[|\Delta|]\phi\approx\operatorname{E}[|\tilde{\Delta}|]\phi\approx\operatorname{E}[\tilde{\Delta}\mid X\geq X_{0}]\Pr[X\geq X_{0}]\phi-\operatorname{E}[\tilde{\Delta}\mid X<X_{0}]\Pr[X<X_{0}]\phi\\ =\int_{Y_{0}}^{\infty}\left(\frac{a_{1}+r_{1}+\xi+b_{1}\theta y}{l_{1}\sqrt{\theta}}-\frac{a_{2}+r_{2}+\xi+b_{2}/(\theta y)}{l_{2}}\sqrt{\theta}\right)f(y;r_{2},r_{1})\,\mathrm{d}y\\ -\int_{0}^{Y_{0}}\left(\frac{a_{1}+r_{1}+\xi+b_{1}\theta y}{l_{1}\sqrt{\theta}}-\frac{a_{2}+r_{2}+\xi+b_{2}/(\theta y)}{l_{2}}\sqrt{\theta}\right)f(y;r_{2},r_{1})\,\mathrm{d}y,$ (102)
with $f(y;r_{2},r_{1})$ defined by (9). Applying (10),
$\int_{0}^{Y_{0}}f(y;r_{2},r_{1})\,\mathrm{d}y=I\left(\frac{Y_{0}}{Y_{0}+1};r_{2},r_{1}\right)=1-\int_{Y_{0}}^{\infty}f(y;r_{2},r_{1})\,\mathrm{d}y.$ (103)
Similarly, combining (9) and (10) with the identities (5) and (6) gives
[…]
Substituting (103)–(109) into (102),
$\begin{split}\operatorname{E}[|\Delta|]\phi&\approx\left(\frac{a_{1}+r_{1}+\xi}{l_{1}\sqrt{\theta}}-\frac{a_{2}+r_{2}+\xi}{l_{2}}\sqrt{\theta}\right)\left(1-2I\left(\frac{Y_{0}}{Y_{0}+1};r_{2},r_{1}\right)\right)\\ &\quad+\frac{b_{1}r_{2}\sqrt{\theta}}{(r_{1}-1)l_{1}}\left(1-2I\left(\frac{Y_{0}}{Y_{0}+1};r_{2}+1,r_{1}-1\right)\right)\\ &\quad-\frac{b_{2}r_{1}}{(r_{2}-1)l_{2}\sqrt{\theta}}\left(1-2I\left(\frac{Y_{0}}{Y_{0}+1};r_{2}-1,r_{1}+1\right)\right).\end{split}$ (110)
Using (39), (41), (56), (57), (60) and the identities (7) and (8) into (110) yields
$\begin{split}\operatorname{E}[|\Delta|]\phi&\approx 2\left(\frac{1}{A}+r_{1}+\mu_{1}+\xi\right)\frac{Y_{0}^{r_{1}+\mu_{1}-\mu_{2}-1}}{(Y_{0}+1)^{2r_{1}+\mu_{1}-\mu_{2}-1}\mathrm{B}(r_{1}+\mu_{1}-\mu_{2},r_{1})}\\ &\quad\cdot\left(\frac{1}{r_{1}l_{1}\sqrt{\theta}}+\frac{Y_{0}\sqrt{\theta}}{(r_{1}+\mu_{2}-\mu_{2})l_{2}}\right).\end{split}$ (111)
Substituting (70), (71) and (111) into (96), and using (93), the average number of required groups for the RR estimator is obtained as
$\operatorname{E}[G]\phi\approx\left(\frac{1}{A}+r_{1}+\mu_{1}+\xi\right)\left(\frac{1}{l_{1}\sqrt{\theta}}+\frac{\sqrt{\theta}}{l_{2}}\right.\\ \left.+\frac{Y_{0}^{r_{1}+\mu_{1}-\mu_{2}-1}}{(Y_{0}+1)^{2r_{1}+\mu_{1}-\mu_{2}-1}\mathrm{B}(r_{1}+\mu_{1}-\mu_{2},r_{1})}\left(\frac{1}{r_{1}l_{1}\sqrt{\theta}}+\frac{Y_{0}\sqrt{\theta}}{(r_{1}+\mu_{1}-\mu_{2})l_{2}}\right)\right),$ (112)
where $Y_{0}$ is given by (101), and with $\mu_{1}=2$, $\mu_{2}=0$.
Figure 10 represents the normalized average number of groups obtained from simulation, as well as its theoretical approximation (112). Overall, the graphs follow similar patterns to those observed for $\operatorname{E}[M_{i}+N_{i}]\phi$ with separate sampling. The approximation is accurate for small or moderate $A$, and less so for large $A$. Discontinuities due to the choice of $r_{1}$ are minimally visible, and steep changes due to rounding $s_{1}$ and $s_{2}$ can be observed in the simulation results for large $A$. Comparing $\operatorname{E}[G]$ for different values of $\theta$ only makes sense for $l_{1}$ and $l_{2}$ fixed; that is, between Figures 10(a) and 10(b) or between Figures 10(c) and 10(d). For $l_{1}$ and $l_{2}$ given, it can be seen from (112) that the minimum of $\operatorname{E}[G]$ with respect to $\theta$ does not necessarily occur when $\lambda\theta=1$, as was the case for separate sampling.
(a) $l_{1}=1$, $l_{2}=1$, $\theta=1$, $\phi=0.01$
($p_{1}=0.01$, $p_{2}=0.01$)
(b) $l_{1}=1$, $l_{2}=1$, $\theta=16$, $\phi=0.01$
($p_{1}=0.04$, $p_{2}=0.0025$)
(c) $l_{1}=2$, $l_{2}=5$, $\theta=16$, $\phi=0.01$
($p_{1}=0.04$, $p_{2}=0.0025$)
(d) $l_{1}=2$, $l_{2}=5$, $\theta=1/16$, $\phi=0.01$
($p_{1}=0.0025$, $p_{2}=0.04$)
Line: approximation. Dots: simulation.
Figure 10: Normalized average number of groups for RR with group sampling
2.7 Efficiency with group sampling
The efficiency of the estimator is defined, as in Subsection 2.4, by comparing the estimation variance with the lowest variance that could be achieved by a fixed-size estimator with the same average numbers of samples, as given by the Cramér–Rao bound. With group sampling, the average number of samples required from population $i$ is $\operatorname{E}[G]l_{i}$. Equating $n_{i}$ to $\operatorname{E}[G]l_{i}$ in (89), the efficiency $\eta_{\mathrm{gr}}$ for an unbiased estimator of a generic parameter $\zeta$ with group sampling is obtained as
$\eta_{\mathrm{gr}}=\frac{\displaystyle\left(\frac{\partial\zeta}{\partial p_{1}}\right)^{2}\frac{p_{1}(1-p_{1})}{l_{1}}+\left(\frac{\partial\zeta}{\partial p_{2}}\right)^{2}\frac{p_{2}(1-p_{2})}{l_{2}}}{\operatorname{E}[G]\operatorname{Var}[\hat{\zeta}]}.$ (113)
Particularizing for RR, i.e. for $\zeta=\theta$,
$\eta_{\mathrm{gr}}=\frac{\displaystyle\frac{1-p_{1}}{p_{1}l_{1}}+\frac{1-p_{2}}{p_{2}l_{2}}}{\operatorname{E}[G]\operatorname{Var}[\hat{\theta}]/\theta^{2}}.$ (114)
Approximating $\operatorname{Var}[\theta]/\theta^{2}\approx A$, the efficiency for RR with group sampling is expressed as
$\eta_{\mathrm{gr}}\approx\frac{(1-p_{1})/(l_{1}\sqrt{\theta})+(1-p_{2})\sqrt{\theta}/l_{2}}{A\operatorname{E}[G]\phi},$ (115)
where $\operatorname{E}[G]\phi$ is given by (112) with $\mu_{1}=2$, $\mu_{2}=0$.
Simulation results for the efficiency are shown in Figure 11. The values are obtained using (114) with $\operatorname{E}[G]$ and $\operatorname{Var}[\hat{\theta}]$ replaced by the sample mean and sample MSE. The approximation (115) is also displayed. The figure contains only two specific cases for brevity. The results are in general little sensitive to $\lambda$ and $\theta$; but, similarly to what was observed with separate sampling, the theoretical approximation becomes more conservative for large $A$ or $\phi$. Comparing with Figure 9, group sampling is seen to be less efficient than separate sampling, in accordance with the number of required groups being the maximum over the two populations, as seen in (94). This causes an efficiency loss of approximately $0.15$ for values of $\lambda$ in the range 0.010.010.010.010.10.10.10.1, and a less substantial loss for $\lambda$ small.
(a) $l_{1}=1$, $l_{2}=1$, $\theta=16$, $\phi=0.1$
($p_{1}=0.4$, $p_{2}=0.025$)
(b) $l_{1}=3$, $l_{2}=1$, $\theta=1/16$, $\phi=0.01$
($p_{1}=0.0025$, $p_{2}=0.04$)
Line: approximation. Dots: simulation.
Figure 11: Efficiency for RR with group sampling
3 Estimation of log relative risk
This section describes the LRR estimator and analyzes its properties. The presentation, as will also be the case for subsequent estimators, follows the same logical course as that in Section 2, but can be shorter, thanks to the similarities with RR. Subsection 3.1 addresses the estimation procedure with separate sampling, and Subsection 3.2 considers group sampling.
3.1 Separate sampling. Estimation procedure and results
The process for estimating the LRR, $\Theta=\log(p_{1}/p_{2})$, is analogous to that for RR: it consists of two stages, each of which applies IBS to each population. The first stage uses fixed IBS parameters, $r_{1}$, $r_{2}$, and the second uses IBS parameters $s_{1}$, $s_{2}$ computed from the results $M_{1}$, $M_{2}$ of the first stage by means of the variable $X$ defined in (21), with $\epsilon_{1}=\epsilon_{2}=1/2$. An unbiased estimation $\hat{\Theta}$ is computed from the second-stage results $N_{1}$, $N_{2}$, according to the expression given next; and the error requirement is in this case defined in terms of MSE (rather than relative MSE), or equivalently variance:
$\operatorname{Var}[\hat{\Theta}]\leq A.$ (116)
In the second stage, writing $\Theta=\log p_{1}-\log p_{2}$, the estimator for the logarithm of a probability described in Mendo [16, section 3] can be used for each of these two terms. Specifically, $-H_{N_{i}-1}+H_{s_{i}-1}$, where $H_{n}$ is the $n$-th harmonic number defined in (3), is an unbiased estimator of $\log p_{i}$, with variance less than $1/(s_{i}-1)$ for any $p_{i}\in(0,1)$ [16]. Therefore,
$\hat{\Theta}=-H_{N_{1}-1}+H_{N_{2}-1}+H_{s_{1}-1}-H_{s_{2}-1}$ (117)
is a conditionally unbiased estimator of $\Theta$ given $s_{1},s_{2}$, which implies that it is also unconditionally unbiased; and, since the observations of the two populations are independent,
$\operatorname{Var}[\hat{\Theta}\mid s_{1},s_{2}]<\frac{1}{s_{1}-1}+\frac{1}{s_{2}-1}.$ (118)
Defining the error function for LRR as the right-hand side of (118), i.e. as (29) with
$\mu_{1}=1,\qquad\mu_{2}=1,\qquad\mu_{12}=0,$ (119)
and assuming (31), it follows that
$\operatorname{Var}[\hat{\Theta}^{2}]=\operatorname{E}[\operatorname{E}[\hat{\Theta}^{2}\mid s_{1},s_{2}]]-\Theta^{2}<\operatorname{E}\left[e(s_{1},s_{2})+\Theta^{2}\right]-\Theta^{2}<A.$ (120)
That is, condition (31) ensures that (116) holds for any $p_{1},p_{2}\in(0,1)$.
As in the RR case, $s_{1}$ and $s_{2}$ are obtained from the equation system formed by (20) and (33), except with $\mu_{1},\mu_{2},\mu_{12}$ given by (119). The solution is (64)–(66). The values of $s_{1}$ and $s_{2}$ should then be rounded while maintaining (31). Using the first-order approximations (37)–(42), $\gamma$ and $\delta_{1}$ are obtained as as in (61) and (62); and (22) is satisfied. The average sample sizes are approximately given by (70) and (71).
The curvature function is defined as for RR, and the condition $c(A,r_{1},\xi)=0$ is expressed by (82). The fact that $\mu_{1}=\mu_{2}$ and $\mu_{12}=0$ for LRR implies that the positive solution of this equation has a simple expression,
$A=\frac{1}{(r_{1}+\mu_{1}+\xi)(r_{1}-1)}.$ (121)
In analogy with RR, the value $\xi=1$ is used for the LRR estimator, with $r_{1}$ given by (81), which in this case can be written explicitly as
$r_{1}=\max\left\{3,\left\lceil-1/2+\sqrt{(3/2)^{2}+1/A}\right\rceil\right\}.$ (122)
The pairs $(A,r_{1})$ determined by (121) for $\xi=1$ are shown in Figure 3. With $\xi=1$ and $r_{1}$ given by (122), the average sample sizes are bounded by (79) and (80).
Algorithm 1 describes the estimation procedure for LRR, as well as the estimator properties.
The simulation results for $\operatorname{E}[M_{i}+N_{i}]\phi$, $i=1,2$ with $\xi=1$ and varying $r_{1}$ are analogous to those shown for RR in Figure 4 (light dots), and are omitted for brevity. As in that case, with $r_{1}$ chosen as in (122) the bounds (79) and (80) are close to the actual average numbers of samples, and these approximately take their minimum values with respect to $r_{1}$. The simulated MSE, also omitted, has a similar behavior to that in Figure 6, with values always smaller than the target $A$, and very close to it unless $A$ is large.
Figure 12 shows, in two specific cases, the theoretical bounds (79) and (80) and simulated $\operatorname{E}[M_{i}+N_{i}]\phi$, for $\xi=1$ and with $r_{1}$ given by (122). The plotted values have a variation pattern similar to that observed for RR, with jump discontinuities and steep changes; again, the latter are only visible in the simulation results, and most apparent for large $A$. The values are slightly lower than those for RR in the rightmost region of the graphs (compare Figures 12(a) and 12(b) with 5(a) and 5(b) respectively), due to the fact that $\mu_{1}$ is $1$ for LRR and $2$ for RR. The effect of $\mu_{1}$ is only appreciable for large $A$, as it stems from (79) and (80). The ratio between average sample sizes obtained from simulation is close to $\lambda$, as can be seen from Figure 12 (note that the ratio between the bounds is exactly $\lambda$). More specifically, the achieved ratio deviates from $\lambda$ by small or very small percentages, very similar to those in RR (results not shown).
(a) $\lambda=3$, $\theta=1/16$, $\phi=0.01$ ($p_{1}=0.0025$, $p_{2}=0.04$)
(b) $\lambda=1/3$, $\theta=1/16$, $\phi=0.01$ ($p_{1}=0.0025$, $p_{2}=0.04$)
Line: bound. Dots: simulation.
Figure 12: Normalized average sample sizes for LRR with separate sampling
As for the efficiency with separate sampling, particularizing (90) for $\zeta=\Theta$, with $\partial\Theta/\partial p_{1}=1/p_{1}$, $\partial\Theta/\partial p_{2}=-1/p_{2}$ and substituting (79), (80) and (116) results in the same lower bound (92) as for RR (although the obtained values will be slightly different because $\mu_{1}$ and $r_{1}$ are), and the efficiency approaches $1$ when $A$ and $\phi$ tend to 00. Figure 13 compares this bound with simulation results, considering only two specific cases for brevity. The efficiency values from simulation are computed as explained in Subsection 2.4. Results are similar to those for RR (Figures 9(b) and 9(d)), except that for large $A$ the efficiency is larger in LRR, as is the difference between simulation and bound.
(a) $\lambda=1$, $\theta=16$, $\phi=0.1$
($p_{1}=0.4$, $p_{2}=0.025$)
(b) $\lambda=3$, $\theta=1/16$, $\phi=0.01$
($p_{1}=0.0025$, $p_{2}=0.04$)
Line: bound. Dots: simulation.
Figure 13: Efficiency for LRR with separate sampling
3.2 Group sampling. Estimation procedure and results
The estimation of LRR with group sampling uses, as in the RR case, a number of groups $G$ given by (94) in order to provide the necessary amounts of individual samples of the two populations. The procedure is the same as in Subsection 2.5: a new group is taken whenever a sample of either population is required and no surplus samples of that population are available from previous groups. At the end of the process, any leftover samples are discarded.
The analysis in Appendix A.1 shows that the approximation (97) is also valid for LRR; and $\operatorname{E}[G]$ is then given by (112), with $Y_{0}$ as in (101) and with $\mu_{1}=\mu_{2}=1$. Simulation results for the average number of groups, omitted for brevity, are very similar to those for RR (Figure 10), except that for large $A$ they are slightly lower than in RR, for the same reason as with separate sampling.
The efficiency with group sampling has the same approximate expression (115) as for RR, with $\operatorname{E}[G]$ computed as indicated above. The simulation results, shown in Figure 14, follow the same pattern observed for separate sampling: the theoretical approximation is accurate when $A$ and $\phi$ are small, and conservative otherwise; and for large $A$ the efficiency deviates more from the theoretical curve than in RR (compare with Figure 11).
(a) $l_{1}=1$, $l_{2}=1$, $\theta=16$, $\phi=0.1$
($p_{1}=0.4$, $p_{2}=0.025$)
(b) $l_{1}=3$, $l_{2}=1$, $\theta=1/16$, $\phi=0.01$
($p_{1}=0.0025$, $p_{2}=0.04$)
Line: approximation. Dots: simulation.
Figure 14: Efficiency for LRR with group sampling
4 Estimation of odds ratio
4.1 Separate sampling. Estimation procedure and results
Several approaches are conceivable to estimate the OR $\psi$ defined in (2). One method that could be employed is to estimate $p_{1}/p_{2}$ and $(1-p_{2})/(1-p_{1})$ separately, treating each as a RR and using the two-stage procedure described in Section 2. With this approach, the MSE in the estimation of $\psi$ depends on the errors in the two RR estimations; and the problem is how to distribute the target MSE between these two components so as to approximately achieve a desired ratio of average sample sizes. To this end, another sampling stage could be introduced before the RR estimations, but that would complicate the process. A better approach, which only requires two stages and results in good estimation efficiency, is based on estimating the odds $p_{1}/(1-p_{1})$ and $(1-p_{2})/p_{2}$ separately, using in each case the method in Mendo [16, section 2]. This is detailed next.
The estimation method consists, as in previous sections, of two stages with IBS. The second stage estimates $p_{1}/(1-p_{1})$ for the first population, and $(1-p_{2})/p_{2}$ for the second population. However, there are two differences with respect to the RR and LRR estimators. The first difference is that in the second stage two IBS procedures are used for each population. Given $s_{1}$, $s_{2}$ (which will be computed from the results of the first stage), IBS is applied to population $1$ to obtain $s_{1}$ successes, which requires $N^{\prime}_{1}$ samples. Then IBS is applied again to population $1$ to obtain $s_{1}-\alpha$ failures, with $\alpha=2$ (a different value of $\alpha$ will be used for LOR estimation). This requires $N^{\prime\prime}_{1}$ samples, for a total of $N_{1}=N^{\prime}_{1}+N^{\prime\prime}_{1}$ samples. For population $2$, IBS is applied to obtain $s_{2}-\alpha$ successes, which requires $N^{\prime}_{2}$ samples; and then to obtain $s_{2}$ failures, which requires $N^{\prime\prime}_{2}$ samples, for a total of $N_{2}=N^{\prime}_{2}+N^{\prime\prime}_{2}$ samples. The average numbers of samples used by the second stage are, for $\alpha=2$,
$\operatorname{E}[N_{1}]$ $=\frac{s_{1}}{p_{1}}+\frac{s_{1}-2}{1-p_{1}}=\frac{s_{1}-2p_{1}}{p_{1}(1-p_{1})}<\frac{s_{1}}{p_{1}(1-p_{1})},$ (123)
$\operatorname{E}[N_{2}]$ $=\frac{s_{2}-2}{p_{2}}+\frac{s_{2}}{1-p_{2}}=\frac{s_{2}-2(1-p_{2})}{p_{2}(1-p_{2})}<\frac{s_{2}}{p_{2}(1-p_{2})}.$ (124)
Thus $\operatorname{E}[N_{i}]$, $i=1,2$ is approximately inversely proportional to $\bar{p}_{i}$ defined as
$\bar{p}_{i}=p_{i}(1-p_{i}).$ (125)
Using the same ideas as for the RR estimator (Subsection 2.1), it can be seen that to approximately achieve a given ratio of average sample sizes, considering only the samples used by the second stage for the moment, $s_{1}/s_{2}$ should be roughly proportional to
$\bar{\theta}=\frac{\bar{p}_{1}}{\bar{p}_{2}}.$ (126)
In view of this, the second difference from previous estimators is that the first stage in this case needs to use samples with parameters $\bar{p}_{i}$, $i=1,2$, rather than $p_{i}$, so as to acquire information about $\bar{\theta}$. Specifically, for $i=1,2$, the first stage applies IBS with $r_{i}$ successes to a sequence of samples with parameter $\bar{p}_{i}$. Denoting the number of samples used from this sequence by $\bar{M}_{i}$, it is clear from (11) that $\operatorname{E}[\bar{M}_{i}]=r_{i}/\bar{p}_{i}$. The samples with parameter $\bar{p}_{i}$ must be generated from samples with parameter $\bar{p}_{i}$. A simple, efficient procedure for this will be given later, and it will be shown that with this procedure the total number of samples with parameter $p_{i}$ required by the first stage, $M_{i}$, has an average equal to $3r_{i}/(2\bar{p}_{i})$. Therefore, considering both sampling stages, the average numbers of samples from the two populations satisfy
$\frac{\operatorname{E}[M_{1}+N_{1}]}{\operatorname{E}[M_{2}+N_{2}]}=\frac{3r_{1}/2+s_{1}}{(3r_{2}/2+s_{2})\bar{\theta}}.$ (127)
This is analogous to (18), and in consequence $s_{1}$ and $s_{2}$ can be chosen using a similar approach as for RR, described next.
The first stage produces the variables $\bar{M}_{1}$ and $\bar{M}_{2}$, from which $X$ is defined as
$X=\frac{\bar{M}_{2}-\epsilon_{2}}{\bar{M}_{1}-\epsilon_{1}},$ (128)
where $\epsilon_{1}=\epsilon_{2}=1/2$ as in Section 2. For a target relative MSE given by $A$, the second-stage IBS parameters $s_{1}$ and $s_{2}$ are obtained from $X$ by solving the equation system formed by (20) and (33), where the values of the design parameters $\gamma$, $\delta_{1}$ and $\delta_{2}$ are yet to be specified; and the solutions will then have to be rounded to integer values, as usual. Once $s_{1}$, $s_{2}$ are known, the second stage is carried out, from which $N^{\prime}_{1}$, $N^{\prime\prime}_{1}$, $N^{\prime}_{2}$, $N^{\prime\prime}_{2}$ are obtained. According to Mendo [16, section 2], and taking into account that $\alpha=2$,
$\frac{(s_{1}-1)\,N^{\prime\prime}_{1}}{(s_{1}-2)\,(N^{\prime}_{1}-1)}$
is a conditionally unbiased estimator of $p_{1}/(1-p_{1})$ given $s_{1}$;
$\frac{(s_{2}-1)\,N^{\prime}_{2}}{(s_{2}-2)\,(N^{\prime\prime}_{2}-1)}$
is a conditionally unbiased estimator of $(1-p_{2})/p_{2}$ given $s_{2}$; and for $s_{1},s_{2}\geq 3$ the conditional variances of these estimators are respectively less than
$\frac{p_{1}^{2}}{(s_{1}-2)(1-p_{1})^{2}}\left(1-\frac{\bar{p}_{1}}{s_{1}-2+2p_{1}}\right)$
and
$\frac{(1-p_{2})^{2}}{(s_{2}-2)p_{2}^{2}}\left(1-\frac{\bar{p}_{2}}{s_{2}-2p_{2}}\right).$
Therefore, since the observations are independent,
$\hat{\psi}=\frac{(s_{1}-1)(s_{2}-1)\,N^{\prime\prime}_{1}N^{\prime}_{2}}{(s_{1}-2)(s_{2}-2)\,(N^{\prime}_{1}-1)(N^{\prime\prime}_{2}-1)}$ (129)
is an unbiased estimator of $\psi$, and for $s_{1},s_{2}\geq 3$
$\frac{\operatorname{E}\left[\hat{\psi}^{2}\mid s_{1},s_{2}\right]}{\psi^{2}}=\frac{p_{2}^{2}(1-p_{1})^{2}}{p_{1}^{2}(1-p_{2})^{2}}\left(\operatorname{Var}\left[\frac{(s_{1}-1)N^{\prime\prime}_{1}}{(s_{1}-2)(N^{\prime}_{1}-1)}\mathrel{}\middle|\mathrel{}s_{1}\right]+\frac{p_{1}^{2}}{(1-p_{1})^{2}}\right)\\ \quad\cdot\left(\operatorname{Var}\left[\frac{(s_{2}-1)N^{\prime}_{2}}{(s_{2}-2)(N^{\prime\prime}_{2}-1)}\mathrel{}\middle|\mathrel{}s_{2}\right]+\frac{(1-p_{2})^{2}}{p_{2}^{2}}\right)\\ \leq\left(\frac{1}{s_{1}-2}\left(1-\frac{\bar{p}_{1}}{s_{1}-2+2p_{1}}\right)+1\right)\left(\frac{1}{s_{2}-2}\left(1-\frac{\bar{p}_{2}}{s_{2}-2p_{2}}\right)+1\right),$ (130)
which implies that, for any $p_{1},p_{2}\in(0,1)$,
$\frac{\operatorname{E}\left[\hat{\psi}^{2}\mid s_{1},s_{2}\right]}{\psi^{2}}<\frac{1}{s_{1}-2}+\frac{1}{s_{2}-2}+\frac{1}{(s_{1}-2)(s_{2}-2)}+1.$ (131)
Based on (131), the error function $e(s_{1},s_{2})$ for OR is defined as in (29) with
$\mu_{1}=2,\qquad\mu_{2}=2,\qquad\mu_{12}=1;$ (132)
and then, by the same reasoning as in (32), if the rounded values of $s_{1}$ and $s_{2}$ satisfy (31) this guarantees that $\operatorname{Var}[\hat{\psi}]/\psi^{2}<A$ for any $p_{1},p_{2}\in(0,1)$.
The procedure to generate samples with parameter $\bar{p}_{i}$ is as follows. Taking two samples with parameter $p_{i}$ as inputs is clearly sufficient to produce a sample with parameter $\bar{p}_{i}$. Namely, one possible criterion is (a) to output success if and only if the first input is a success and the second is a failure. However, if the first input happens to be a failure the second input need not be observed. This occurs with probability $1-p_{i}$. Alternatively, (b) the output can be defined to be success if the first input is a failure and the second is a success. In this case, the second input is not needed if the first is a success, which occurs with probability $p_{i}$. If criteria (a) or (b) are randomly chosen with equal probabilities, the average number of inputs required to produce an output is $21+(1-p_{i})/2+p_{i}/2=3/$ 2. This is an instance of a Bernoulli factory [14].
The $\bar{M}_{i}$ samples with parameter $\bar{p}_{i}$, $i=1,2$ required by the first stage are generated with the above method, requiring a total of $M_{i}$ samples of population $i$ as inputs. For the subsequent analysis it is necessary to characterize the relationship between $\operatorname{E}[M_{i}]$ and $\operatorname{E}[\bar{M}_{i}]$. It cannot be directly concluded from the preceding paragraph that $\operatorname{E}[M_{i}]/\operatorname{E}[\bar{M}_{i}]=3/2$, because the IBS stopping rule could conceivably introduce some deviation in this ratio. However, this equality turns out to be true. More generally, if $\beta$ denotes the average number of inputs used by the Bernoulli factory to produce an output,
$\operatorname{E}[M_{i}]=\beta\operatorname{E}[\bar{M}_{i}]=\frac{\beta r_{i}}{\bar{p}_{i}},$ (133)
where $\beta=3/2$ for the described method. To see this, let $\beta_{\mathrm{s}}$ and $\beta_{\mathrm{f}}$ be the average number of inputs required to produce an output, conditioned on the output being success or failure respectively. Then
$\beta=\beta_{\mathrm{s}}\bar{p}_{i}+\beta_{\mathrm{f}}(1-\bar{p}_{i}).$ (134)
IBS with parameter $r_{i}$ is applied to the outputs of the Bernoulli factory, and consumes $\bar{M}_{i}$ of those outputs, of which $r_{i}$ are successes and $\bar{M}_{i}-r_{i}$ are failures. Therefore, using (11),
$\operatorname{E}[M_{i}]=\beta_{\mathrm{s}}r_{i}+\beta_{\mathrm{f}}\left(\operatorname{E}[\bar{M}_{i}]-r_{i}\right)=\beta_{\mathrm{s}}r_{i}+\frac{\beta_{\mathrm{f}}r_{i}(1-\bar{p}_{i})}{\bar{p}_{i}},$ (135)
which combined with (134) gives (133).
The characterization of $\operatorname{E}[M_{i}+N_{i}]$, $i=1,2$, and the ensuing selection of $\gamma$, $\delta_{1}$ and $\delta_{2}$, relies, as in the RR case, on using first-order approximations for $s_{1}$ and $s_{2}$ as functions of $X$ and $1/X$ respectively. Expressions (37)–(42) remain valid for OR, and (52) and (53) hold with $\theta$ replaced by $\bar{\theta}$:
$\operatorname{E}[X]$ $\approx\frac{r_{2}\bar{\theta}}{r_{1}-1},$ (136)
$\operatorname{E}\left[\frac{1}{X}\right]$ $\approx\frac{r_{1}}{(r_{2}-1)\bar{\theta}}.$ (137)
In addition, (54) and (55) have to be modified to take into account that the first stage uses samples with parameter $\bar{p}_{i}$, obtained by transforming samples with parameter $p_{i}$, $i=1,2$. Specifically, from (37), (38), (123), (124), (133), (136) and (137), and introducing the term $\xi$ to account for the effect of rounding $s_{1}$ and $s_{2}$,
[…]
and analogously
$\operatorname{E}[M_{2}+N_{2}]=\frac{3r_{2}/2+s_{2}-2(1-2p_{2})}{\bar{p}_{2}}\approx\frac{a_{2}+3r_{2}/2+\xi}{\bar{p}_{2}}+\frac{b_{2}r_{1}}{(r_{2}-1)\bar{p}_{1}}.$ (139)
According to (39)–(42), (138) and (139), the condition (17) for the ratio of average sample sizes will be satisfied regardless of $p_{1},p_{2}$ if
$\frac{1}{A}+\frac{3r_{1}}{2}+\mu_{1}+\xi$ $=\frac{\lambda r_{1}}{\gamma(r_{2}-1)}\left(\frac{1}{A}+\delta_{1}+\mu_{1}\right),$ (140)
$\frac{\lambda}{\gamma}\left(\frac{1}{A}+\frac{3r_{2}}{2}+\mu_{2}+\xi\right)$ $=\frac{r_{2}}{r_{1}-1}\left(\frac{1}{A}+\delta_{2}+\mu_{2}\right).$ (141)
By analogy with (58) and (59), a simple solution to (140) and (141) is obtained if $3r_{1}/2+\mu_{1}=3r_{2}/2+\mu_{2}$ and $\delta_{1}+\mu_{1}=\delta_{2}+\mu_{2}$. The former condition is compatible with $r_{1},r_{2}\in\mathbb{N}$ only if $2(\mu_{1}-\mu_{2})/3\in\mathbb{Z}$. In particular, this holds if
$\mu_{1}=\mu_{2}.$ (142)
This is the case for OR estimation, and will also be true for LOR estimation (Section 5). Thus, in the sequel (142) will be assumed to hold. The two indicated conditions then reduce to
$r_{2}=r_{1},\qquad\delta_{2}=\delta_{1},$ (143)
and the solution to (140) and (141)
$\gamma$ $=\lambda,$ (144)
$\delta_{1}$ $=\left(\frac{1}{A}+\frac{3r_{1}}{2}+\mu_{1}+\xi\right)\frac{r_{1}-1}{r_{1}}-\frac{1}{A}-\mu_{1}.$ (145)
An argument analogous to that used in Subsection 2.2 shows that (22) is satisfied.
The values of $s_{1}$, $s_{2}$ before rounding, obtained by solving (20) and (33) with $\delta_{1}$ and $\delta_{2}$ as in (143), are given by (64)–(66) as usual. The curvature function is defined as in (76) with $r_{1}$ replaced by $23r_{1}/$ 2 and with $\mu_{1}=\mu_{2}$:
$\begin{split}c(A,r_{1},\xi)&=A\left(\left(\frac{1}{A}+\frac{3r_{1}}{2}+\mu_{1}+\xi\right)\frac{r_{1}-1}{r_{1}}-\frac{1}{A}\right)^{2}\\ &\quad+2\left(\left(\frac{1}{A}+\frac{3r_{1}}{2}+\mu_{1}+\xi\right)\frac{r_{1}-1}{r_{1}}-\frac{1}{A}\right)-\mu_{12},\end{split}$ (146)
from which the condition $c(A,r_{1},\xi)=0$ is
$\begin{split}&\left(\frac{3r_{1}}{2}+\mu_{1}+\xi\right)^{2}(r_{1}-1)^{2}A^{2}\\ &\quad+\left(2\left(\frac{3r_{1}}{2}+\mu_{1}+\xi\right)(r_{1}-1)^{2}-\mu_{12}r_{1}^{2}\right)A+1-2r_{1}=0.\end{split}$ (147)
Similarly to previous estimators, $\xi$ is set to $1$ and $r_{i}$ is chosen as in (81), where $c(A,r_{1},\xi)$ is given by (146). Equivalently, the curve formed by the pairs $(A,r_{1})$ that solve (147) can be plotted, as shown in Figure 3; and then, for a given $A$, (81) corresponds to rounding up the value obtained from the curve, with a minimum of $3$. Defining
$\bar{\phi}=\sqrt{\bar{p}_{1}\bar{p}_{2}},$ (148)
this choice of $\xi$ and $r_{1}$ results, by analogy with Subsection 2.3, in the upper bounds
$\operatorname{E}[M_{1}+N_{1}]\bar{\phi}$ $<\left(\frac{1}{A}+\frac{3r_{1}}{2}+\mu_{1}+1\right)\left(\frac{1}{\sqrt{\lambda\bar{\theta}}}+\sqrt{\lambda\bar{\theta}}\right)\sqrt{\lambda},$ (149)
$\operatorname{E}[M_{2}+N_{2}]\bar{\phi}$ $<\left(\frac{1}{A}+\frac{3r_{1}}{2}+\mu_{1}+1\right)\left(\frac{1}{\sqrt{\lambda\bar{\theta}}}+\sqrt{\lambda\bar{\theta}}\right)\frac{1}{\sqrt{\lambda}}.$ (150)
The estimation procedure for OR is specified in Algorithm 2, where the properties of the estimator are also indicated. (The algorithm covers the LOR case as well, to be presented in Section 5.)
Algorithm 2 Estimator of OR or LOR
  1. Input parameters:
    Target $A$, interpreted as relative MSE for OR, or MSE for LOR;
    Desired ratio of average sample sizes $\lambda$, or group sizes $l_{1}$, $l_{2}$.
  2. Estimation procedure:
  3. 1.
    Define $\mu_{1},\mu_{2},\mu_{12}$ as in (132) for OR, or as in (159) for LOR.
    Define $\epsilon_{1}=\epsilon_{2}=1/2$ and $\xi=1$.
    Define $\alpha=2$ for OR, or $\alpha=0$ for LOR.
    Define $\lambda=l_{1}/l_{2}$ if group sampling is applied.
  4. 2.
    Compute $r_{1}$ from (81) with $c(A,r_{1},\xi)$ given by (146) (for LOR the explicit expression (162) can equivalently be used), or obtain it by rounding up the value from Figure 3, with a minimum of $3$. Set $r_{2}=r_{1}$.
  5. 3.
    Set $\gamma=\lambda$ and compute $\delta_{1}$ from (145).
  6. 4.
    (First sampling stage): For each $i=1,2$, using the method below, generate samples with parameter $\bar{p}_{i}$ until $r_{i}$ successes are obtained. Let $\bar{M}_{i}$ be the number of generated samples.
    To generate a sample with parameter $\bar{p}_{i}$, choose one of these options equally likely:
    • (a)
      Take a sample from population $i$. If failure, output failure. Else, take another sample from population $i$ and output the opposite of its value.
    • (b)
      Take a sample from population $i$. If success, output failure. Else, take another sample from population $i$ and output its value.
    Let $M_{i}$ be the total number of samples used from population $i$.
  7. 5.
    Compute $X$ from (128). Compute $s_{1}$, $s_{2}$ from (64)–(66), and then round one of them up and the other down, or both up, to fulfill (31).
  8. 6.
    (Second sampling stage): From population $1$, take as many samples as needed, $N^{\prime}_{1}$, to obtain $s_{1}$ successes; then as many as needed, $N^{\prime\prime}_{1}$, to obtain ${s_{1}-\alpha}$ failures. From population $2$, take as many samples as needed, ${N^{\prime}_{2}}$, to obtain ${s_{2}-\alpha}$ successes; then as many as needed, $N^{\prime\prime}_{2}$, to obtain $s_{2}$ failures. Let $N_{1}=N_{1}^{\prime}+N_{1}^{\prime\prime}$, $N_{2}=N_{2}^{\prime}+N_{2}^{\prime\prime}$.
  9. 7.
    Compute $\hat{\psi}$ from (129), or $\hat{\Psi}$ from (157).
  10. Output:
    Estimation $\hat{\psi}$ or $\hat{\Psi}$.
  11. Properties:
    Unbiased, with relative MSE (OR) or MSE (LOR) less than $A$;
    Average sample sizes: (149), (150); approximate ratio $\lambda$;
    Efficiency with separate sampling: (152); approaches $1$ for $A$ small
    Average number of groups: (154), (155);
    Efficiency with group sampling: (156).
The relative MSE obtained from simulations is compared with the target $A$ in Figure 15. The simulation consists, as with previous estimators, of $10^{6}$ realizations for each combination of parameters. The difference between simulation results and target is seen to increase with $A$ and with $\phi$, as for RR. However, in OR the difference observed for large $\phi$ vanishes when $A$ is small, unlike in RR (compare the leftmost parts of Figures 6(a) and 15(a)). This is related to the fact that the terms $1-\bar{p}_{1}/(s_{1}-2+2p_{1})$ and $1-\bar{p}_{2}/(s_{2}-2p_{2})$ in (130), which are replaced by $1$ in the uniform bound (131), approach $1$ for large $s_{1}$, $s_{2}$. Thus when $A$ is small, which gives large values of $s_{1}$ and $s_{2}$, the uniform bound (131) is almost as good as (130). In RR, on the other hand, the corresponding factors are $(s_{1}-2)(1-p_{1})/(s_{1}-2+2p_{1})<1-p_{1}$ and $1-p_{2}$, as it is seen comparing (27) and (28), and these do not tend to $1$ for large $s_{1}$, $s_{2}$. (In LRR the factors are more cumbersome, see Mendo [16, theorem 2]; but for large $s_{1}$, $s_{2}$ they tend to the same values $1-p_{1}$, $1-p_{2}$ as in RR.)
(a) $\lambda=1$, $\theta=16$, $\phi=0.1$
($p_{1}=0.025$, $p_{2}=0.4$)
(b) $\lambda=3$, $\theta=1/16$, $\phi=0.01$
($p_{1}=0.04$, $p_{2}=0.0025$)
Dashed line: target. Dots: simulation.
Figure 15: Relative MSE for OR with separate sampling
Figure 16 shows simulation results for $\operatorname{E}[M_{i}+N_{i}]\bar{\phi}$, $i=1,2$. The values are very similar to those for $\operatorname{E}[M_{i}+N_{i}]\phi$ in RR (Figures 5(a) and 5(b)) except that for OR they are somewhat larger in the rightmost part of the graphs. This is due to the $3/2$ factor in OR, whose effect on the average sample sizes is only significant for large $A$.
(a) $\lambda=3$, $\theta=1/16$, $\phi=0.01$ ($p_{1}=0.0025$, $p_{2}=0.04$)
(b) $\lambda=1/3$, $\theta=1/16$, $\phi=0.01$ ($p_{1}=0.0025$, $p_{2}=0.04$)
Line: bound. Dots: simulation.
Figure 16: Normalized average sample sizes for OR with separate sampling
The efficiency of the OR estimator with separate sampling results from particularizing (90) to $\zeta=\psi$, with $\partial\psi/\partial p_{1}=\psi/\bar{p}_{1}$, $\partial\psi/\partial p_{2}=-\psi/\bar{p}_{2}$:
$\eta_{\mathrm{sep}}=\frac{\displaystyle\frac{1}{\operatorname{E}[M_{1}+N_{1}]\bar{p}_{1}}+\frac{1}{\operatorname{E}[M_{2}+N_{2}]\bar{p}_{2}}}{\operatorname{Var}[\psi]/\psi^{2}}.$ (151)
Then, using (149) and (150) and considering that $\operatorname{Var}[\hat{\psi}]/\psi^{2}<A$, the efficiency can be bounded as
$\eta_{\mathrm{sep}}>\frac{1}{1+A(3r_{1}/2+\mu_{1}+1)},$ (152)
where $\mu_{1}=2$. This is similar to the bound (92) for RR and LRR, the only differences being the $3/2$ factor in (152), arising from the Bernoulli factory, and the last factor in (92) (but note that the latter is approximately $1$ for small $\phi$). By the same arguments used for RR and LRR, $\eta_{\mathrm{sep}}$ for OR approaches $1$ when $A$ tends to 00.
Figure 17 shows simulation results for $\eta_{\mathrm{sep}}$, and compares them with the bound (152). In the simulation, $\eta_{\mathrm{sep}}$ is computed as in previous sections, i.e. using sample averages in (151). The values are seen to be similar to those for RR and LRR, except that with large $A$ the efficiency is slightly lower for OR. Again, this is a consequence of the $3/2$ factor. As in RR and LRR, the difference between simulation and bound tends to be larger when $\phi$ is increased. However, in OR that difference vanishes for small $A$, unlike in the other cases (compare Figure 17(a) with Figure 9(b) or 13(a)). This agrees with the behavior of the MSE discussed earlier.
(a) $\lambda=1$, $\theta=16$, $\phi=0.1$
($p_{1}=0.4$, $p_{2}=0.025$)
(b) $\lambda=3$, $\theta=1/16$, $\phi=0.01$
($p_{1}=0.0025$, $p_{2}=0.04$)
Line: bound. Dots: simulation.
Figure 17: Efficiency for OR with separate sampling
4.2 Group sampling. Estimation procedure and results
Group sampling for OR estimation consumes a number of groups $G$ given by (94), as with previous estimators, in order to provide the required amounts of samples $M_{i}+N_{i}$ of each population $i=1,2$; and $\operatorname{E}[G]$ can be approximately computed from (96). It is shown in Appendix A.2 that, for $\bar{\phi}$ small, the term $\operatorname{E}[|\Delta|]$ in (96) can be expressed as
$\operatorname{E}\left[|\Delta|\right]\approx\operatorname{E}\left[\left|\frac{3r_{1}/2+s_{1}}{l_{1}p_{1}}-\frac{3r_{2}/2+s_{2}}{l_{2}p_{2}}\right|\right].$ (153)
Thus, proceeding as in Subsection 2.6 and taking into account (142),
$\begin{split}\operatorname{E}[G]\bar{\phi}&\approx\left(\frac{1}{A}+\frac{3r_{1}}{2}+\mu_{1}+\xi\right)\left(\frac{1}{l_{1}\sqrt{\bar{\theta}}}+\frac{\sqrt{\bar{\theta}}}{l_{2}}\right.\\ &\quad\left.+\frac{Y_{0}^{r_{1}-1}}{(Y_{0}+1)^{2r_{1}-1}\mathrm{B}(r_{1},r_{1})r_{1}}\left(\frac{1}{l_{1}\sqrt{\bar{\theta}}}+\frac{Y_{0}\sqrt{\bar{\theta}}}{l_{2}}\right)\right),\end{split}$ (154)
where
$Y_{0}=\frac{r_{1}}{2\lambda\bar{\theta}(r_{1}-1)}\left(\lambda\bar{\theta}-1+\sqrt{(\lambda\bar{\theta}-1)^{2}+\frac{4\lambda\bar{\theta}(r_{1}-1)^{2}}{r_{1}^{2}}}\right)$ (155)
and $\mu_{1}=2$. The simulation results for $\operatorname{E}[G]\bar{\phi}$, not shown, are similar to those for $\operatorname{E}[G]\phi$ in RR and LRR, with the difference that for large $A$ the values are slightly greater in OR compared with those two cases, as it has been observed with separate sampling.
The efficiency with group sampling is obtained particularizing (113) for $\zeta=\psi$ and approximating $\operatorname{Var}[\hat{\psi}]/\psi^{2}\approx A$:
$\eta_{\mathrm{gr}}\approx\frac{1/(l_{1}\sqrt{\bar{\theta}})+\sqrt{\bar{\theta}}/l_{2}}{A\operatorname{E}[G]\bar{\phi}},$ (156)
where $\operatorname{E}[G]\bar{\phi}$ is given by (154) and (155). Figure 18 represents the theoretical approximation (156), as well as results from simulation. As with separate sampling, the efficiency for large $A$ is slightly lower than in RR and LRR (Figures 11 and 14), and for small $A$ there is almost no difference between simulation results and theoretical approximation even if $\bar{\phi}$ is large.
(a) $\lambda=1$, $\theta=16$, $\phi=0.1$
($p_{1}=0.4$, $p_{2}=0.025$)
(b) $\lambda=3$, $\theta=1/16$, $\phi=0.01$
($p_{1}=0.0025$, $p_{2}=0.04$)
Line: approximation. Dots: simulation.
Figure 18: Efficiency for OR with group sampling
5 Estimation of log odds ratio
5.1 Separate sampling. Estimation procedure and results
The estimation method for the LOR, $\Psi=\log(p_{1}(1-p_{2})/(p_{2}(1-p_{1})))$, has few differences compared to that presented for OR in Section 4. The second stage estimates $\log(p_{1}/(1-p_{1}))$ and $\log(p_{2}/(1-p_{2}))$ separately, using for each the method described in Mendo [16, section 3], which consists of two IBS procedures with the same parameter. Specifically, to estimate $\log(p_{i}/(1-p_{i}))$, $i=1,2$, IBS is applied until $s_{i}$ successes are obtained, which requires a random number $N_{i}^{\prime}$ of observations from population $i$, and then IBS is applied until $s_{i}$ failures are obtained, which requires $N_{i}^{\prime\prime}$ additional observations. These are the same steps as for OR, but with $\alpha=0$. Then, conditioned on $s_{1}$ and $s_{2}$, which are obtained in the first stage, $-H_{N_{i}^{\prime}-1}+H_{N_{i}^{\prime\prime}-1}$ is a conditionally unbiased estimator of $\log(p_{i}/(1-p_{i}))$ with conditional variance less than $1/(s_{i}-5/4)$ for any $p_{i}\in(0,1)$ [16, theorem 3]. Therefore,
$\hat{\Psi}=-H_{N_{1}^{\prime}-1}+H_{N_{1}^{\prime\prime}-1}+H_{N_{2}^{\prime}-1}-H_{N_{2}^{\prime\prime}-1}$ (157)
is a conditionally unbiased estimator of $\Psi$ with
$\operatorname{Var}[\hat{\Psi}^{2}\mid s_{1},s_{2}]<\frac{1}{s_{1}-5/4}+\frac{1}{s_{2}-5/4}.$ (158)
It follows that $\hat{\Psi}$ is unconditionally unbiased, and, defining its error function as in (29) with
$\mu_{1}=5/4,\qquad\mu_{2}=5/4,\qquad\mu_{12}=0,$ (159)
inequality (31) guarantees that $\operatorname{Var}[\hat{\Psi}]<A$ for any $p_{1},p_{2}\in(0,1)$.
The number of samples from population $i=1,2$ used in the second stage, $N_{i}=N_{i}^{\prime}+N_{i}^{\prime\prime}$, has an average
$\operatorname{E}[N_{i}]=\frac{s_{i}}{p_{i}}+\frac{s_{i}}{1-p_{i}}=\frac{s_{i}}{\bar{p}_{i}},$ (160)
where $\bar{p}_{i}$ is defined by (125). Thus, for the same reason as in OR estimation, the first stage for LOR estimation must be based on samples with parameters $\bar{p}_{i}$, $i=1,2$, generated from observations with parameters $p_{i}$. For each $i=1,2$, given $r_{i}$, a random number $\bar{M}_{i}$ of samples with parameter $\bar{p}_{i}$ are generated until $r_{i}$ successes are obtained, which in turn requires $M_{i}$ observations from population $i$. Then $X$ is computed from $\bar{M}_{1}$ and $\bar{M}_{2}$ as in (128). Since (142) holds, imposing the additional condition (143), $\gamma$ and $\delta_{1}$ are obtained from (144) and (145) as for OR, and (22) is satisfied. Once $X$, $\gamma$ and $\delta_{1}$ are known, $s_{1}$ and $s_{2}$ are computed using (64)–(66), as for the other estimators, and then rounded with the restriction (31). The resulting average numbers of samples are bounded by (149) and (150).
The condition $c(A,r_{1},\xi)=0$, from which $r_{1}$ and $r_{2}$ are obtained, is expressed by (147). Because $\mu_{1}=\mu_{2}$ and $\mu_{12}=0$, this simplifies in the same way as for LRR, and has the positive solution
$A=\frac{1}{(3r_{1}/2+\mu_{1}+\xi)(r_{1}-1)}.$ (161)
The LOR estimator uses $\xi=1$ and $r_{1}$ given by (81), which in this case is written as
$r_{1}=\displaystyle\max\left\{3,\left\lceil-1/4+\sqrt{(5/4)^{2}+2/(3A)}\right\rceil\right\}.$ (162)
As with previous estimators, this choice of $r_{1}$ yields average numbers of samples close to their minimum values with respect to this parameter (results not shown). The curve $c(A,r_{1},1)=0$, plotted in Figure 3, is almost indistinguishable from that for OR; and (162) corresponds to rounding up the ordinate values of this curve, with a minimum of $3$.
The estimation procedure for LOR is summarized in Algorithm 2, which also lists the properties of the estimator.
Particularizing (90) for $\zeta=\Psi$, with $\partial\Psi/\partial p_{1}=1/\bar{p}_{1}$, $\partial\Psi/\partial p_{2}=-1/\bar{p}_{2}$, and using (149), (150) and the fact that $\operatorname{Var}[\hat{\Psi}]<A$, it follows that the efficiency with separate sampling is bounded by (152) with $\mu_{1}=5/4$. Figure 19 shows the results, which are very similar to those for OR.
(a) $\lambda=1$, $\theta=16$, $\phi=0.1$
($p_{1}=0.4$, $p_{2}=0.025$)
(b) $\lambda=3$, $\theta=1/16$, $\phi=0.01$
($p_{1}=0.0025$, $p_{2}=0.04$)
Line: bound. Dots: simulation.
Figure 19: Efficiency for LOR with separate sampling
5.2 Group sampling. Estimation procedure and results
The group sampling procedure is analogous to that described for the other estimators. The average number of groups $\operatorname{E}[G]$ is computed from (96), where $\operatorname{E}[|\Delta|]$ satisfies (153), as shown in Appendix A.2. Thus $\operatorname{E}[G]$ is approximately given by (154) and (155), as for OR, but with $\mu_{1}=5/4$ and with a possibly different computed value for $r_{1}$. Simulation results are omitted.
The efficiency with group sampling is expressed by (156), as for OR. The results, plotted in Figure 20, are not the same as in that case because of the differences in $\mu_{1}$ and $r_{1}$, which affect $\operatorname{E}[G]$ and thus the efficiency; but they are seen to be very similar.
(a) $\lambda=1$, $\theta=16$, $\phi=0.1$
($p_{1}=0.4$, $p_{2}=0.025$)
(b) $\lambda=3$, $\theta=1/16$, $\phi=0.01$
($p_{1}=0.0025$, $p_{2}=0.04$)
Line: approximation. Dots: simulation.
Figure 20: Efficiency for LOR with group sampling
6 Conclusions
Two-stage sequential methods have been presented to estimate the RR $p_{1}/p_{2}$, the OR $p_{1}(1-p_{2})/(p_{2}(1-p_{1}))$ or their logarithmic versions using independent binary observations from two populations with parameters $p_{1}$ and $p_{2}$. The estimators are unbiased; guarantee that the relative mean square error, or the mean square error for the logarithmic versions, is less than a target value $A$ irrespective of $p_{1}$ and $p_{2}$; and approximately achieve a prescribed ratio of average sample sizes when samples are taken from each population separately. The estimators can also be used with group sampling. In this case, samples are taken simultaneously from the two populations in fixed-size groups, and individual samples are extracted from those groups as needed, with a number of samples possibly left unused at the end of the process. The properties of unbiasedness and guaranteed accuracy are maintained with group sampling (and an exact sample size ratio is imposed by the sampling). Bounds and approximate expressions have been derived for the average sample sizes and the average number of sample groups, respectively. The estimation efficiency, defined in terms of the Cramér–Rao bound, has been characterized in the same way. The efficiency is generally good, both with separate sampling and with group sampling; and is close to $1$ for small $A$. Algorithms 1 and 2 specify the estimation procedure and the properties of the estimators.
The described method can be extended to estimate other functions of $p_{1}$ and $p_{2}$, provided that an error function can be defined as in (29) and (63), and that a Bernoulli factory can be found, if needed, to generate samples with parameters equal to the probabilities to which $\operatorname{E}[N_{1}]$ and $\operatorname{E}[N_{2}]$ are approximately inversely proportional (as was done for OR and LOR), where $N_{1}$ and $N_{2}$ are the numbers of observations required in the second stage. In addition, the Bernoulli factory needs to have an average number of inputs per output equal to a constant $\beta$.
As an example, consider the estimation of $p_{1}p_{2}$. A straightforward approach would be to generate samples with parameter $p_{1}p_{2}$ from samples with parameters $p_{1}$ and $p_{2}$, using a procedure analogous to the Bernoulli factory described in Section 4, and then apply IBS to the generated samples. However, this does not give any control on the proportion of samples from the two populations, and the average number of samples required from each population scales with the inverse of $p_{1}p_{2}$. Instead, the method presented for RR and LRR can be used: two IBS processes with parameters $s_{1}$ and $s_{2}$ are respectively applied to the two populations (second stage), which requires $N_{1}$ and $N_{2}$ samples, where $s_{1}$ and $s_{2}$ are obtained from a previous pair of IBS processes with fixed parameters $r_{1}$ and $r_{2}$ (first stage). Then,
$\frac{(s_{1}-1)(s_{2}-1)}{(N_{1}-1)(N_{2}-1)}$
is an unbiased estimator of $p_{1}p_{2}$, and by means of (14) a target relative error can be guaranteed. The same expressions as in RR and LRR apply for the average sample sizes, average number of groups and efficiency. In particular, the average sample sizes with separate sampling are approximately in the desired ratio, and, for $p_{1}/p_{2}$ fixed, they scale with the inverse of $\sqrt{p_{1}p_{2}}$.
Appendix A Approximation of $\operatorname{E}[|\Delta|]$
A.1 For relative risk and log relative risk
The random variable $M_{1}$ in RR and LRR has a negative binomial distribution with parameters $r_{1}$ and $p_{1}=\phi\sqrt{\theta}$. For $\phi\rightarrow 0$, the distribution of $M_{1}\phi$ tends to a gamma distribution with location parameter $r_{1}$ and scale parameter $1/\sqrt{\theta}$, because
$\begin{split}\lim_{\phi\rightarrow 0}\Pr[M_{1}\phi\leq x]&=\lim_{\phi\rightarrow 0}\sum_{k=r_{1}}^{\lfloor x/\phi\rfloor}\binom{k-1}{r_{1}-1}\left(\phi\sqrt{\theta}\right)^{r_{1}}\left(1-\phi\sqrt{\theta}\right)^{k-r_{1}}\\ &=\frac{\theta^{r_{1}/2}}{(r_{1}-1)!}\int_{0}^{x}t^{r_{1}-1}\exp(-t\sqrt{\theta})\,\mathrm{d}t.\end{split}$ (163)
Likewise, the distribution of $M_{2}\phi$ tends to that of a gamma random variable with location parameter $r_{2}$ and scale parameter $\sqrt{\theta}$. By the continuous mapping theorem [21, theorem 2.3], the variable $X$ defined in (21) converges in distribution to the ratio of these two gamma random variables, and therefore $Y=X/\theta$ and $1/Y$ converge in distribution to beta prime random variables with parameters $r_{2}$, $r_{1}$ and $r_{1}$, $r_{2}$ respectively [8, section 4.4]. It is easy to see that $Y^{2}$ and $1/Y^{2}$ are uniformly integrable as $\phi\rightarrow 0$, which implies [6, theorem 25.12] that the variances of $Y$ and $1/Y$ converge to those of the referred beta prime distributions. Thus, approximating $\operatorname{Var}[X]$ and $\operatorname{Var}[1/X]$ by their values for $\phi\rightarrow 0$,
$\operatorname{Var}[X]$ $\approx\frac{r_{2}(r_{1}+r_{2}-1)\theta^{2}}{(r_{1}-2)(r_{1}-1)^{2}},$ (164)
$\operatorname{Var}[1/X]$ $\approx\frac{r_{1}(r_{1}+r_{2}-1)}{(r_{2}-2)(r_{2}-1)^{2}\theta^{2}}.$ (165)
[…]
$\begin{split}\textstyle\operatorname{Var}\left[\Delta\right]&=\operatorname{Var}\left[\frac{M_{1}}{l_{1}}\right]+\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{1}}{l_{1}}\mathrel{}\middle|\mathrel{}s_{1}\right]\right]+\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}\mathrel{}\middle|\mathrel{}s_{1}\right]\right]\\ &\quad+\operatorname{Var}\left[\frac{M_{2}}{l_{2}}\right]+\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{2}\right]\right]+\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{2}\right]\right],\end{split}$ (166)
where (i) corresponds to the first and fourth summands, (ii) to the second and fifth, and (iii) to the third and sixth. The claim is then that (166) is well approximated by the sum of the third and sixth terms on the right-hand side. This is substantiated next, considering the small-$\phi$ regime.
The first and fourth summands in (166) are obtained from (13), (60) and (93) as
$\operatorname{Var}\left[\frac{M_{1}}{l_{1}}\right]\phi^{2}l_{1}l_{2}$ $=\frac{r_{1}(1-p_{1})}{\lambda\theta}\approx\frac{r_{1}}{\lambda\theta},$ (167)
$\operatorname{Var}\left[\frac{M_{2}}{l_{2}}\right]\phi^{2}l_{1}l_{2}$ $=r_{2}(1-p_{2})\lambda\theta\approx r_{2}\lambda\theta=(r_{1}+\mu_{1}-\mu_{2})\lambda\theta.$ (168)
Using (11), (13) and (37), and including the rounding term $\xi$ when computing $\operatorname{E}[s_{1}]$, the second and third summands in (166) can be expressed as
$\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{1}}{l_{1}}\mathrel{}\middle|\mathrel{}s_{1}\right]\right]\phi^{2}l_{1}l_{2}$ $=\frac{\operatorname{E}\left[s_{1}(1-p_{1})\right]}{\lambda\theta}\approx\frac{a_{1}+\xi+b_{1}\operatorname{E}[X]}{\lambda\theta},$ (169)
$\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}\mathrel{}\middle|\mathrel{}s_{1}\right]\right]\phi^{2}l_{1}l_{2}$ $=\frac{\operatorname{Var}\left[s_{1}\right]}{\lambda\theta}\approx\frac{b_{1}^{2}\operatorname{Var}[X]}{\lambda\theta}.$ (170)
Substituting (52) and (164) into (169) and (170), and using (39), (41), (57) and (60),
$\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{1}}{l_{1}}\mathrel{}\middle|\mathrel{}s_{1}\right]\right]\phi^{2}l_{1}l_{2}$ $\approx\frac{1}{\lambda\theta}\left(\frac{1}{A}+\mu_{1}+\xi\right)+\frac{1}{A}+r_{1}+\mu_{1}+\xi,$ (171)
$\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}\mathrel{}\middle|\mathrel{}s_{1}\right]\right]\phi^{2}l_{1}l_{2}$ $\approx\lambda\theta\left(\frac{1}{A}+r_{1}+\mu_{1}+\xi\right)^{\!2}\frac{2r_{1}+\mu_{1}-\mu_{2}-1}{(r_{1}-2)(r_{1}+\mu_{2}-\mu_{1})}.$ (172)
By an analogous procedure, making use of (11), (13), (38) and then (39), (41), (53), (56), (60) and (165), the last two terms in (166) can be written as
$\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{2}\right]\right]\phi^{2}l_{1}l_{2}$ $\approx\lambda\theta\left(\frac{1}{A}+\mu_{2}+\xi\right)+\frac{1}{A}+r_{1}+\mu_{1}+\xi,$ (173)
$\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{2}\right]\right]\phi^{2}l_{1}l_{2}$ $\approx\frac{1}{\lambda\theta}\left(\frac{1}{A}+r_{1}+\mu_{1}+\xi\right)^{\!2}\frac{2r_{1}+\mu_{1}-\mu_{2}-1}{r_{1}(r_{1}+\mu_{1}-\mu_{2}-2)}.$ (174)
In practice, the target $A$ will typically be smaller, or much smaller, than $1$. For example, for RR a relative RMSE of $20\%$ corresponds to $A=0.04$. On the other hand, $\xi$ equals $1$, and the values of $r_{1}$ obtained from (81) are small compared to $1/A$, as seen in Figure 3. In the example, $A=0.04$ gives $r_{1}=5$. Thus, from (167), (168), (171)–(174) the following behavior is observed, both for RR ($\mu_{1}=2$, $\mu_{1}=0$) and for LRR ($\mu_{1}=\mu_{2}=1$): if $\lambda\theta\approx 1$ the four terms $\operatorname{Var}[M_{1}/l_{1}]$, $\operatorname{Var}[M_{2}/l_{2}]$, $\operatorname{E}[\operatorname{Var}[N_{1}/l_{1}\mid s_{1}]]$ and $\operatorname{E}[\operatorname{Var}[N_{2}/l_{2}\mid s_{2}]]$ are significantly smaller than $\operatorname{Var}[\operatorname{E}[N_{1}/l_{1}\mid s_{1}]]$ and $\operatorname{Var}[\operatorname{E}[N_{2}/l_{2}\mid s_{2}]]$; if $\lambda\theta>1$ the term $\operatorname{Var}[\operatorname{E}[N_{1}/l_{1}\mid s_{1}]]$ is much greater than the other five; and if $\lambda\theta<1$ the term $\operatorname{Var}[\operatorname{E}[N_{2}/l_{2}\mid s_{2}]]$ is much greater than the others. The conclusion is that, regardless of $\lambda\theta$, one or both of the terms $\operatorname{Var}[\operatorname{E}[N_{i}/l_{i}\mid s_{i}]]$, $i=1,2$ in (166) dominate. This means that to account for the variability of $\Delta$ it is sufficient, with small approximation error, to only consider the variability of $s_{1}$, $s_{2}$, and disregard that of $M_{1}$, $M_{2}$ and that of $N_{1}$, $N_{2}$ conditioned on $s_{1}$, $s_{2}$. Therefore, to compute $\operatorname{E}[|\Delta|]$, the variables $M_{1}$, $M_{2}$, $N_{1}$, $N_{2}$ in (95) can be replaced by their conditional means given $s_{1}$, $s_{2}$, which yields (97).
A.2 For odds ratio and log odds ratio
The decomposition (166) obtained for $\operatorname{Var}[\Delta]$ in RR and LRR also applies in OR and LOR, but the expressions for the six summands are different. To compute $\operatorname{Var}[M_{i}/l_{i}]$, $i=1,2$ the law of total variance [5, theorem 12.2.6] is used:
$\operatorname{Var}\left[M_{i}\right]=\operatorname{E}\left[\operatorname{Var}\left[M_{i}\mid\bar{M}_{i}\right]\right]+\operatorname{Var}\left[\operatorname{E}\left[M_{i}\mid\bar{M}_{i}\right]\right].$ (175)
The first stage applies IBS to samples with parameter $\bar{p}_{i}$, using a number $\bar{M}_{i}$ of those samples, of which $r_{i}$ are successes. Each sample with parameter $\bar{p}_{i}$ is generated by the Bernoulli factory described in Subsection 4.1, taking samples with parameter $p_{i}$ as inputs. With this factory, the average number of inputs needed to produce an output is $3/2$; a success output always uses $2$ inputs, and a failure output uses either $1$ or $2$ inputs. Let $\pi_{i}$ denote the probability that $2$ inputs are used, conditioned on the output being a failure. Then $3/2=2\bar{p}_{i}+(1+\pi_{i})(1-\bar{p}_{i})$, which yields
$\pi_{i}=1-\frac{1}{2(1-\bar{p}_{i})}=\frac{1-2\bar{p}_{i}}{2(1-\bar{p}_{i})}.$ (176)
Thus, for a given $\bar{M}_{i}$, the average number of required inputs is
$\operatorname{E}[M_{i}\mid\bar{M}_{i}]=2r_{i}+(1+\pi_{i})(\bar{M}_{i}-r_{i})=\frac{(3-4\bar{p}_{i})\bar{M}_{i}+r_{i}}{2(1-\bar{p}_{i})}.$ (177)
The term $\operatorname{Var}\left[M_{i}\mid\bar{M}_{i}\right]$ in (175) equals the variance of a binomial random variable with parameters $\bar{M}_{i}-r_{i}$ and $\pi_{i}$. Therefore, computing $\operatorname{E}[\bar{M}_{i}]$ from (11) and using (176),
$\operatorname{E}\left[\operatorname{Var}\left[M_{i}\mid\bar{M}_{i}\right]\right]=\pi_{i}(1-\pi_{i})\operatorname{E}[\bar{M}_{i}-r_{i}]=\frac{(1-2\bar{p}_{i})r_{i}}{4\bar{p}_{i}(1-\bar{p}_{i})}.$ (178)
On the other hand, using (177) and computing $\operatorname{Var}[\bar{M}_{i}]$ from (13),
$\operatorname{Var}\left[\operatorname{E}\left[M_{i}\mid\bar{M}_{i}\right]\right]=\frac{(3-4\bar{p}_{i})^{2}}{4(1-\bar{p}_{i})^{2}}\operatorname{Var}\left[\bar{M}_{i}\right]=\frac{(3-4\bar{p}_{i})^{2}r_{i}}{4\bar{p}_{i}^{2}(1-\bar{p}_{i})}.$ (179)
Substituting (178) and (179) into (175),
$\operatorname{Var}\left[M_{i}\right]=\frac{(14\bar{p}_{i}^{2}-23\bar{p}_{i}+9)r_{i}}{4\bar{p}_{i}^{2}(1-\bar{p}_{i})}=\frac{(9-14\bar{p}_{i})r_{i}}{4\bar{p}_{i}^{2}},$ (180)
and therefore, taking into account (143),
$\operatorname{Var}\left[\frac{M_{1}}{l_{1}}\right]\bar{\phi}^{2}l_{1}l_{2}$ $=\frac{(9-14\bar{p}_{1})r_{1}}{4\lambda\bar{\theta}},$ (181)
$\operatorname{Var}\left[\frac{M_{2}}{l_{2}}\right]\bar{\phi}^{2}l_{1}l_{2}$ $=\frac{\lambda\bar{\theta}(9-14\bar{p}_{2})r_{1}}{4}.$ (182)
For the other four summands in (166) it is helpful, as in Appendix A.1, to approximate the variances of $X$ and $1/X$ by their values for small $\bar{\phi}$, which in this case yields
$\operatorname{Var}[X]$ $\approx\frac{r_{2}(r_{1}+r_{2}-1)\bar{\theta}^{2}}{(r_{1}-2)(r_{1}-1)^{2}},$ (183)
$\operatorname{Var}[1/X]$ $\approx\frac{r_{1}(r_{1}+r_{2}-1)}{(r_{2}-2)(r_{2}-1)^{2}\bar{\theta}^{2}}.$ (184)
The term $\operatorname{E}[\operatorname{Var}[N_{1}/l_{1}\mid X]]$ is obtained, making use of (13) again, as
$\begin{split}&\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{1}}{l_{1}}\mathrel{}\middle|\mathrel{}s_{1}\right]\right]\bar{\phi}^{2}l_{1}l_{2}=\frac{\operatorname{E}[\operatorname{Var}[N_{1}^{\prime}\mid s_{1}]]\bar{\phi}^{2}}{\lambda}+\frac{\operatorname{E}[\operatorname{Var}[N_{1}^{\prime\prime}\mid s_{1}]]\bar{\phi}^{2}}{\lambda}\\ &\quad=\frac{\operatorname{E}[s_{1}](1-p_{1})\bar{\phi}^{2}}{\lambda p_{1}^{2}}+\frac{\operatorname{E}[(s_{1}-\alpha)]p_{1}\bar{\phi}^{2}}{\lambda(1-p_{1})^{2}}\\ &=\frac{(p_{1}^{3}+(1-p_{1})^{3})\operatorname{E}[s_{1}]\bar{\phi}^{2}-\alpha p_{1}^{3}\bar{\phi}^{2}}{\lambda\bar{p}_{1}^{2}}=\frac{\operatorname{E}[s_{1}](1-3\bar{p}_{1})-\alpha p_{1}^{3}}{\lambda\bar{\theta}},\end{split}$ (185)
with $\alpha$ equal to $2$ for OR and 00 for LOR. Using (37), including the rounding term $\xi$, and then substituting (39), (40), (136) and (142)–(145),
$\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{1}}{l_{1}}\mathrel{}\middle|\mathrel{}s_{1}\right]\right]\bar{\phi}^{2}l_{1}l_{2}\approx\frac{(a_{1}+\xi+b_{1}\operatorname{E}[X])(1-3\bar{p}_{1})-\alpha p_{1}^{3}}{\lambda\bar{\theta}}\\ \approx\left(\frac{1}{\lambda\bar{\theta}}\left(\frac{1}{A}+\mu_{1}+\xi\right)+\frac{1}{A}+\frac{3r_{1}}{2}+\mu_{1}+\xi\right)(1-3\bar{p}_{1})-\frac{\alpha p_{1}^{3}}{\lambda\bar{\theta}}.$ (186)
As for $\operatorname{Var}[\operatorname{E}[N_{1}/l_{1}\mid s_{1}]]$, from (11), (37), (40), (142)–(145), and (183),
[…]
Analogously, using (11), (13), (38), (41), (42), (137), (142)–(145) and (184),
$\begin{split}&\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{2}\right]\right]\bar{\phi}^{2}l_{1}l_{2}\approx\lambda\bar{\theta}\left((a_{2}+\xi+b_{2}\operatorname{E}[1/X])(1-3\bar{p}_{2})-\alpha(1-p_{2})^{3}\right)\\ &\quad\approx\left(\lambda\bar{\theta}\left(\frac{1}{A}+\mu_{1}+\xi\right)+\frac{1}{A}+\frac{3r_{1}}{2}+\mu_{1}+\xi\right)(1-3\bar{p}_{2})-\alpha\lambda\bar{\theta}(1-p_{2})^{3},\end{split}$ (188)
$\begin{split}&\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{2}\right]\right]\bar{\phi}^{2}l_{1}l_{2}\approx\lambda\bar{\theta}b_{2}^{2}\operatorname{Var}[1/X]\approx\frac{1}{\lambda\bar{\theta}}\left(\frac{1}{A}+\frac{3r_{1}}{2}+\mu_{1}+\xi\right)^{\!2}\,\frac{2r_{1}-1}{r_{1}(r_{1}-2)}.\end{split}$ (189)
From (181), (182), (186)–(189), a similar assessment as in Appendix A.1 can be made, and the conclusion is the same: for typical values of $A$, both for OR ($\mu_{1}=\mu_{2}=2$, $\alpha=2$) and LOR ($\mu_{1}=\mu_{2}=5/4$, $\alpha=0$), and for any $\lambda$ and $\bar{\theta}$, one or both of the terms $\operatorname{Var}[\operatorname{E}[N_{i}/l_{i}\mid s_{i}]]$ in (166) dominate the sum, which implies that (153) is a good approximation for small $\bar{\phi}$.