### v4

# Estimation of relative risk, odds ratio and their logarithms with guaranteed accuracy and controlled sample size ratio

Abstract
Given two populations from which independent binary observations are taken with parameters $p_{1}$ and $p_{2}$ respectively, estimators are proposed for the relative risk $p_{1}/p_{2}$, the odds ratio $p_{1}(1-p_{2})/(p_{2}(1-p_{1}))$ and their logarithms. The sampling strategy used by the estimators is based on two-stage sequential sampling applied to each population, where the sample sizes of the second stage depend on the results observed in the first stage. The estimators guarantee that the relative mean-square error, or the mean-square error for the logarithmic versions, is less than a target value for any $p_{1},p_{2}\in(0,1)$, and the ratio of average sample sizes from the two populations is close to a prescribed value. The estimators can also be used with group sampling, whereby samples are taken in batches of fixed size from the two populations simultaneously, each batch containing samples from the two populations. The efficiency of the estimators with respect to the Cramér–Rao bound is good, and in particular it is close to $1$ for small values of the target error.
keywords:
Estimation, sequential sampling, group sampling, relative risk, odds ratio, log odds ratio, mean-square error, efficiency.
pacs:
MSC2010 Classification]62F10, 62L12
1 Introduction
Let $p_{1},p_{2}\in(0,1)$ denote the probabilities of occurrence of a given dichotomous attribute in two different populations. The problem of estimating the relative risk (RR) or risk ratio,
$\theta=\frac{p_{1}}{p_{2}},$ (1)
from binary observations of the two populations arises frequently in medical and social sciences, as well as in other fields. For example, in a phase-III clinical trial of a vaccine [4, chapter 18] the relevant attribute is the presence of a disease, and the two populations are vaccinated and non-vaccinated people. The odds ratio (OR),
$\psi=\frac{p_{1}(1-p_{2})}{p_{2}(1-p_{1})},$ (2)
is sometimes used instead of RR, as are their logarithmic versions: log relative risk (LRR), $\Theta=\log\theta$, and log odds ratio (LOR), $\Psi=\log\psi$. The latter is especially important in connection with logistic regression [3], which is a prevalent tool in machine learning [7].
When estimating any of these parameters, it is crucial to have knowledge about the accuracy of the estimation. Ideally, the estimation should guarantee a target accuracy, regardless of the unknown $p_{1}$ and $p_{2}$. A common measure of accuracy is the mean-square error (MSE), or its square root, known as root-mean-square error (RMSE), which for unbiased estimators reduce to variance or standard deviation respectively. For non-logarithmic parameters such as RR and OR, these error measurements are meaningful in a relative sense [16], because the significance of a given estimation error can only be assessed by comparing it with the true value of the parameter. Thus, it is natural to require that the RMSE be proportional to the true value of the parameter. On the other hand, for LRR and LOR the estimation error is meaningful by itself, without comparing with the true value, as the logarithm transforms ratios into differences.
A second desirable feature, along with guaranteed accuracy, is to have control on the proportion of sample sizes of the two populations. Consider first the case that the two populations are sampled individually, i.e. with element sampling, meaning that samples from any population can be taken one by one as needed. In many use cases, it may be required that the two sample sizes be similar, or that they approximately satisfy a given ratio. Another possible sampling procedure is group sampling, whereby samples are collected in groups or batches, each containing $l_{1}$ samples from population 11 and $l_{2}$ from population $2$. This imposes a strict sample size ratio of $l_{1}/l_{2}$. Either with element sampling or with group sampling, these conditions only refer to the sample size ratio; the actual sample sizes should be chosen to fulfill the target accuracy.
Note that the difference of the group sampling scheme with respect to element sampling is not only that samples are taken in groups (as opposed to one by one), but also that each group simultaneously contains samples from both populations (instead of each population being sampled separately).
It will be assumed that each population is infinite, and observations are statistically independent. This implies that the observations can be modeled as Bernoulli trials. The requirement that the target accuracy, as defined earlier, be satisfied for all $p_{1},p_{2}\in(0,1)$ makes it necessary to use sequential sampling, because any fixed sample size will fail to satisfy that requirement for low enough $p_{1},p_{2}$. More specifically, this work will make extensive use of inverse binomial sampling (IBS) 12; 15, chapter 2. Denoting the presence or absence of the attribute of interest in a sample as “success” or “failure”, IBS consists in observing samples until a predefined number $r$ of successes is obtained. The number $r$ will be referred to as the parameter of the IBS procedure.
This paper presents unbiased estimators of the four parameters RR, LRR, OR and LOR, that guarantee a target accuracy and provide control on the sample size ratio, irrespective of $p_{1}$ and $p_{2}$. The estimators are based on the two-stage sampling procedure suggested in a previous work [16, section 4]. As argued above, the target accuracy, $A$, is defined as relative MSE for RR and OR, or as MSE for LRR and LOR. The control on the sample size ratio means that, for element sampling of each population, average sample sizes will approximately satisfy a specified proportion. The estimators can also be applied with group sequential sampling, using groups of $l_{1}$ and $l_{2}$ samples from each population. This incurs, as will be seen, a small increase in the average sample size compared to element sampling, but ensures an exact ratio of the (random) sample sizes. In both cases the estimation efficiency, defined in terms of the Cramér–Rao bound, is considerably large, and is close to $1$ for small $A$.
Estimation of the RR, OR or their logarithmic versions is an important problem in statistics. For a review of results in this area, see Mendo [16, section 1]. Variants of the problem considering a desired ratio of sample sizes have been addressed in existing works (often assuming the ratio equal to $1$); see for example Siegmund [20], Agresti [2], Cho [9], Cho [10]. However, to the author’s knowledge, no previously proposed estimators for these parameters guarantee a target accuracy as defined above, i.e. relative error for RR and OR or absolute error for LRR and LOR, while offering control on the proportion of sample sizes.
The following notation and basic identities will be used. The function $\log x$ represents the natural logarithm of $x$. The $n$-th harmonic number is denoted as
$H_{n}=\sum_{k=1}^{n}\frac{1}{k}.$ (3)
Matrices are written in boldface letters, and $\mathbf{Q}^{\intercal}$ represents the transpose of a matrix $\mathbf{Q}$. The regularized incomplete beta function is defined as
$I(x;u,v)=\frac{1}{\mathrm{B}(u,v)}\int_{0}^{x}t^{u-1}(1-t)^{v-1}\,\mathrm{d}\mkern 0.8mut,\quad 0<x<1;\ u,v>0,$ (4)
where $\mathrm{B}(u,v)$ is the beta function; and from Abramowitz and Stegun [1, equations (6.1.15), (6.2.2), (26.5.15)] it follows that
$(v-1)\mathrm{B}(u+1,v-1)$ $=u\,\mathrm{B}(u,v),$ (5)
$(u-1)\mathrm{B}(u-1,v+1)$ $=v\,\mathrm{B}(u,v),$ (6)
$I(x;u,v)-I(x;u+1,v-1)$ $=\frac{x^{u}(1-x)^{v-1}}{u\,\mathrm{B}(u,v)},$ (7)
$I(x;u,v)-I(x;u-1,v+1)$ $=-\frac{x^{u-1}(1-x)^{v}}{v\,\mathrm{B}(u,v)}.$ (8)
The probability density function of a beta prime distribution with parameters $u$, $v$ is denoted as $f(y;u,v)$:
$f(y;u,v)=\frac{y^{u-1}}{\mathrm{B}(u,v)(1+y)^{u+v}},\quad y>0;\ u,v>0.$ (9)
For a random variable $Y$ with this distribution [8, section 4.4],
$\Pr[Y\leq y]=\int_{0}^{y}f(t;u,v)\,\mathrm{d}\mkern 0.8mut=I\left(\frac{y}{y+1};u,v\right).$ (10)
In IBS with success probability $p$, the number $N$ of samples needed to obtain $r$ successes has a negative binomial distribution with parameters $r$ and $p$. Then, for $p\in(0,1)$ 18, equation (3.1); 19, section 4.8.2,
$\operatorname{E}[N]$ $=\frac{r}{p},$ (11)
$\operatorname{E}\left[\frac{1}{N-1}\right]$ $=\frac{p}{r-1}\quad\text{for }r\geq 2,$ (12)
$\operatorname{Var}[N]$ $=\frac{r(1-p)}{p^{2}},$ (13)
$\operatorname{Var}\left[\frac{1}{N-1}\right]$ $\leq\frac{p^{2}(1-p)}{(r-1)^{2}(r-2+2p)}<\frac{p^{2}(1-p)}{(r-1)^{2}(r-2)}\quad\text{for }r\geq 3.$ (14)
The rest of the paper is organized as follows. Section 2 describes the estimation procedure for RR, derives approximate expressions and bounds for the average sample sizes and estimation efficiency, and compares these with values obtained from Monte Carlo simulations. The estimation procedure and the results for LRR, OR and LOR are to some extent analogous, and are presented in Sections 3, 4 and 5. Concluding remarks are given in Section 6.
2 Estimation of relative risk
The estimation procedure for RR with element sampling of the two populations is considered in Section 2.1. First, the general estimation approach is motivated and described (Section 2.1.1). The precise definition of the estimator is then completed (Section 2.1.2). Lastly, theoretical bounds are obtained for the average sample sizes and estimation efficiency, and these are compared with simulation results (Sections 2.1.3 and 2.1.4).
Group sampling is addressed in Section 2.2. The estimation procedure used in this case is described (Section 2.2.1), and then approximations for the average number of groups and efficiency are obtained and compared with simulation results (Sections 2.2.2 and 2.2.3).
2.1 Element sampling
2.1.1 Estimation procedure
The estimator to be presented is unbiased, and uses two-stage sampling. Each of the two sampling stages is comprised of two independent IBS procedures, one for each population. Before explaining the purpose of each stage, it is necessary to define several parameters and variables. The IBS parameters of the first stage are denoted as $r_{1}$ and $r_{2}$ for the two populations respectively. The resulting numbers of samples are negative binomial random variables, $M_{1}$ and $M_{2}$. Similarly, the second-stage IBS procedures have parameters $s_{1}$ and $s_{2}$, and the numbers of samples are $N_{1}$ and $N_{2}$. The parameters $s_{1}$ and $s_{2}$ are obtained from $M_{1}$ and $M_{2}$, as will be seen, and are thus random variables. The total number of samples used from each population $i=1,2$ is then $M_{i}+N_{i}$. From (11) it stems that $\operatorname{E}[M_{i}]=r_{i}/p_{1}$, $\operatorname{E}[N_{i}\mid s_{i}]=s_{i}/p_{i}$, and
$\operatorname{E}[M_{i}+N_{i}]=\frac{r_{i}+\operatorname{E}[s_{i}]}{p_{i}}.$ (15)
Consider a target relative MSE equal to $A$, and a desired ratio $\lambda$ of average sample sizes. Let $\hat{\theta}$ denote the estimation of $\theta$. Since MSE reduces to variance for an unbiased estimator, the conditions that $\hat{\theta}$ must satisfy are
$\frac{\operatorname{Var}[\hat{\theta}]}{\theta^{2}}$ $\leq A,$ (16)
$\frac{\operatorname{E}[M_{1}+N_{1}]}{\operatorname{E}[M_{2}+N_{2}]}$ $\approx\lambda$ (17)
for any $p_{1},p_{2}\in(0,1)$.
The purpose of the first sampling stage is to obtain two pilot sets of samples, one from each population, using predefined values for the IBS parameters $r_{1}$ and $r_{2}$; and from those acquire some knowledge about $\theta$. With this knowledge, suitable values for the second-stage IBS parameters $s_{1}$ and $s_{2}$ are computed such that (16) and (17) are satisfied. The results of the second stage, i.e. $N_{1}$ and $N_{2}$, are then used to produce the final estimate $\hat{\theta}$. The rationale is as follows. According to (15),
$\frac{\operatorname{E}[M_{1}+N_{1}]}{\operatorname{E}[M_{2}+N_{2}]}=\frac{r_{1}+\operatorname{E}[s_{1}]}{(r_{2}+\operatorname{E}[s_{2}])\theta}.$ (18)
Each of the IBS procedures in the second stage provides information about one of the two probabilities $p_{1}$ and $p_{2}$. Given a target accuracy for the second-stage estimate $\hat{\theta}$, there is a trade-off between the parameters $s_{1}$ and $s_{2}$: decreasing one of them causes the information on the corresponding probability to be less accurate, which can be compensated for by increasing the other parameter. In view of (18), this can be exploited for balancing $\operatorname{E}[M_{1}+N_{1}]$ and $\operatorname{E}[M_{2}+N_{2}]$; doing so requires knowledge about $\theta$, which is provided by the first stage.
An initial idea to obtain information about $\theta$ from the first-stage variables $M_{1}$ and $M_{2}$ is to compute an estimate of it using a generalization of the method described in Mendo [16, section 2] for estimating p/(1p)p/(1-p). Namely, from (11) and (12) it follows that $(r_{1}-1)/(M_{1}-1)$ is an unbiased estimator of $p_{1}$ and $M_{2}/r_{2}$ is an unbiased estimator of $1/p_{2}$. Therefore, since the observations used by those estimators are independent,
$\frac{(r_{1}-1)\,M_{2}}{r_{2}\,(M_{1}-1)}$
is an unbiased estimator of $\theta$. Replacing $\theta$ by this estimate and $\operatorname{E}[s_{i}]$ by $s_{i}$, $i=1,2$ in (18) suggests that the condition (17) will be roughly satisfied if $s_{1}$ and $s_{2}$ are chosen so that
$\frac{r_{1}+s_{1}}{r_{2}+s_{2}}\approx\lambda\,\frac{(r_{1}-1)\,M_{2}}{r_{2}\,(M_{1}-1)}.$ (19)
It is more convenient, however, to substitute the requirement (19) by a generalized version:
$\frac{s_{1}+\delta_{1}}{s_{2}+\delta_{2}}=\gamma\,X$ (20)
with $X$ defined as
$X=\frac{M_{2}-\epsilon_{2}}{M_{1}-\epsilon_{1}},$ (21)
where $\gamma>0$, $\delta_{i}\mathbb{\in}\mathbb{R}$, $\epsilon_{i}\in(0,1)$ for $i=1,2$ are design parameters, whose values can be selected to facilitate meeting (17) with good approximation. (Observe that (20) indeed reduces to (19) for $\gamma=\lambda(r_{1}-1)/r_{2}$, $\delta_{i}=r_{i}$, $\epsilon_{1}=1$, $\epsilon_{2}=0$.) In addition, $s_{i}$ and $\delta_{i}$ must satisfy
$s_{i}+\delta_{i}>0,\quad i=1,2,$ (22)
to ensure that the left-hand side of (20) does not involve division by zero and remains positive. It should be noted that (20) may give non-integer values for $s_{1}$ and $s_{2}$, which thus have to be rounded. This will introduce a small additional error in (17).
For $r_{1}\geq 3$, a simple analysis based on (13) and (14) shows that the relative variance of $X$ is bounded uniformly on $p_{1}$, $p_{2}$. On the other hand, for $r_{1}=2$ it is easy to see that the relative variance takes arbitrarily large values as $p_{1}\rightarrow 0$. Thus, to ensure that the variability of $X$ is not too large, the following additional requirement is imposed:
$r_{i}\geq 3,\quad i=1,2.$ (23)
The second sampling stage uses the parameters $s_{1}$ and $s_{2}$, determined in the first stage, to obtain $N_{1}$ and $N_{2}$. By the same reasoning applied earlier,
$\hat{\theta}=\frac{(s_{1}-1)\,N_{2}}{s_{2}\,(N_{1}-1)}$ (24)
is a conditionally unbiased estimation of $\theta$ given $s_{1},s_{2}$; and thus it is also unconditionally unbiased. For $s_{1}\geq 3$ the conditional variance of $(s_{1}-1)/(N_{1}-1)$ can be bounded, using (14), as
$\operatorname{Var}\left[\frac{s_{1}-1}{N_{1}-1}\mathrel{}\middle|\mathrel{}s_{1}\right]\leq\frac{p_{1}^{2}(1-p_{1})}{s_{1}-2+2p_{1}};$ (25)
and for $s_{2}\geq 1$ the conditional variance of $N_{2}/s_{2}$ is, according to (13),
$\operatorname{Var}\left[\frac{N_{2}}{s_{2}}\mathrel{}\middle|\mathrel{}s_{2}\right]=\frac{1-p_{2}}{s_{2}p_{2}^{2}}.$ (26)
Therefore,
$\begin{split}\frac{\operatorname{E}\left[\hat{\theta}^{2}\mid s_{1},s_{2}\right]}{\theta^{2}}&=\frac{p_{2}^{2}}{p_{1}^{2}}\left(\operatorname{Var}\left[\frac{s_{1}-1}{N_{1}-1}\mathrel{}\middle|\mathrel{}s_{1}\right]+p_{1}^{2}\right)\left(\operatorname{Var}\left[\frac{N_{2}}{s_{2}}\mathrel{}\middle|\mathrel{}s_{2}\right]+\frac{1}{p_{2}^{2}}\right)\\ &\leq\left(\frac{1-p_{1}}{s_{1}-2+2p_{1}}+1\right)\left(\frac{1-p_{2}}{s_{2}}+1\right).\end{split}$ (27)
This implies that, for all $p_{1},p_{2}\in(0,1)$,
$\frac{\operatorname{E}\left[\hat{\theta}^{2}\mid s_{1},s_{2}\right]}{\theta^{2}}<\frac{1}{s_{1}-2}+\frac{1}{s_{2}}+\frac{1}{(s_{1}-2)s_{2}}+1.$ (28)
In view of (28), let the function $e(s_{1},s_{2})$ be defined as
$e(s_{1},s_{2})=\frac{1}{s_{1}-\mu_{1}}+\frac{1}{s_{2}-\mu_{2}}+\frac{\mu_{12}}{(s_{1}-\mu_{1})(s_{2}-\mu_{2})},\quad s_{1}>\mu_{1},\ s_{2}>\mu_{2},$ (29)
with
$\mu_{1}=2,\qquad\mu_{2}=0,\qquad\mu_{12}=1.$ (30)
This will be referred to as error function. The parameters $\mu_{1}$, $\mu_{2}$ and $\mu_{12}$ are introduced for convenience; this way the expression of $e(s_{1},s_{2})$ for other estimators will be the same as (29), only with different values of these parameters. Then, requiring
$e(s_{1},s_{2})\leq A$ (31)
guarantees that condition (16) holds; in fact with strict inequality. Namely, from (28)–(31),
$\frac{\operatorname{Var}\left[\hat{\theta}\right]}{\theta^{2}}=\frac{\operatorname{E}\left[\hat{\theta}^{2}\right]}{\theta^{2}}-\frac{\left(\operatorname{E}\left[\hat{\theta}\right]\right)^{2}}{\theta^{2}}=\frac{\operatorname{E}\left[\operatorname{E}\left[\hat{\theta}^{2}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]}{\theta^{2}}-1<\operatorname{E}\left[e(s_{1},s_{2})\right]\leq A.$ (32)
2.1.2 Estimator parameters and approximate average sample sizes
According to (15), to achieve small average sample sizes, $r_{i}$ and $s_{i}$, $i=1,2$ should be as small as possible. In this section, approximate expressions are first derived for the average sample sizes, from which the choice of $s_{1}$, $s_{2}$ is addressed and the values of the estimator parameters $\gamma$, $\delta_{1}$, $\delta_{2}$, $\epsilon_{1}$, $\epsilon_{2}$ are determined. (Section 2.1.3 will discuss how to select the values of $r_{1}$, $r_{2}$.)
Since $e(s_{1},s_{2})$ is a decreasing function of $s_{1}$ and $s_{2}$, to minimize these parameters (31) should be treated as an equality. Thus, $s_{1}$ and $s_{2}$ are determined by
$e(s_{1},s_{2})=A$ (33)
together with (20) (where the values of the design parameters $\gamma$, $\delta_{1}$ and $\delta_{2}$ are yet to be defined). Solving this quadratic equation system yields
$s_{1}$ $=\frac{\gamma X(A(\delta_{2}+\mu_{2})+1)-A(\delta_{1}-\mu_{1})+1+\sqrt{D}}{2A},$ (34)
$s_{2}$ $=\frac{s_{1}+\delta_{1}}{\gamma X}-\delta_{2},$ (35)
where the discriminant $D$ is
$\begin{split}D&=(\gamma X(A(\delta_{2}+\mu_{2})+1)-A(\delta_{1}-\mu_{1})+1)^{2}\\ &\quad-4A\left(\gamma X((A\mu_{1}+1)(\delta_{2}+\mu_{2})+\mu_{1}-\mu_{12})-(A\mu_{1}+1)\delta_{1}\right).\end{split}$ (36)
There would be another pair of solutions where the square root in (34) has a negative sign, but that pair is not valid because it does not satisfy $s_{1}>\mu_{1}$, $s_{2}>\mu_{2}$ as required by (29). This is illustrated in Figure 1, which makes it clear that there is only one solution pair in the valid range; and therefore it corresponds to the positive sign.
Figure 1: Solutions $s_{1},s_{2}$ to (20) and (33) (example with $A=0.05$, $\gamma=0.5$, $\delta_{1}=1$, $\delta_{2}=3$, $X=2.5$, $\mu_{1}=2$, $\mu_{2}=0$, $\mu_{12}=1$)
As indicated in Section 2.1.1, the solutions (34) and (35) have to be rounded, because only integer numbers can be used as IBS parameters. Depending on their specific values it may be necessary to round both of them up, or it may be sufficient to round one up and the other down, if that satisfies (31). A simple criterion, which will be assumed in the rest of the paper, is as follows. First, randomly choose $(i,j)=(1,2)$ or $(2,1)$ with equal probability; then round $s_{i}$ up and $s_{j}$ down, and check if those values satisfy (31). If not, try rounding $s_{i}$ down and $s_{j}$ up. If not valid either, round both values up, which necessarily satisfies (31).
Expressions (34) and (35) give $s_{1}$ and $s_{2}$ as functions of $X$, up to the required rounding, but are difficult to deal with. A natural simplification, which will be helpful in fulfilling (17), is to replace them by the first-order approximations
$s_{1}$ $\approx a_{1}+b_{1}X,$ (37)
$s_{2}$ $\approx a_{2}+\frac{b_{2}}{X},$ (38)
where the coefficients $a_{1}$, $b_{1}$, $a_{2}$, $b_{2}$ are obtained from (34)–(36) as
$a_{1}$ $=\lim_{X\rightarrow 0}s_{1}$ $=1/A+\mu_{1},$ (39)
$b_{1}$ $=\lim_{X\rightarrow\infty}\frac{s_{1}}{X}$ $=\gamma(1/A+\delta_{2}+\mu_{2}),$ (40)
$a_{2}$ $=\lim_{X\rightarrow\infty}s_{2}$ $=1/A+\mu_{2},$ (41)
$b_{2}$ $=\lim_{X\rightarrow 0}s_{2}X$ $=\frac{1/A+\delta_{1}+\mu_{1}}{\gamma}.$ (42)
Figure 2 represents these approximations using the example values (same as in Figure 1) $A=0.05$, $\gamma=0.5$, $\delta_{1}=1$, $\delta_{2}=3$, and with $\mu_{1}$, $\mu_{2}$, $\mu_{12}$ given by (30) as specified for RR. The figure illustrates that the accuracy of the approximations depends on the curvature of $s_{1}$ and $s_{2}$ as functions of $X$ and $1/X$. In fact, for certain values of $A$, $\delta_{1}$ and $\delta_{2}$ the approximations (37) and (38) are exact. This will be analyzed in Section 2.1.3.
(a) $s_{1}$
(b) $s_{2}$
Figure 2: First-order approximations of $s_{1}$ and $s_{2}$ (example with $A=0.05$, $\gamma=0.5$, $\delta_{1}=1$, $\delta_{2}=3$, $\mu_{1}=2$, $\mu_{2}=0$, $\mu_{12}=1$)
Substituting (37) and (38) into (15), and introducing an additional term $\xi$ to account for the effect of rounding $s_{1}$ and $s_{2}$, the average sample sizes are approximated as
$\operatorname{E}[M_{1}+N_{1}]$ $\approx\frac{a_{1}+r_{1}+\xi}{p_{1}}+\frac{b_{1}}{p_{1}}\operatorname{E}[X],$ (43)
$\operatorname{E}[M_{2}+N_{2}]$ $\approx\frac{a_{2}+r_{2}+\xi}{p_{2}}+\frac{b_{2}}{p_{2}}\operatorname{E}\left[\frac{1}{X}\right].$ (44)
The term $\xi$ models the average increase in $s_{1}$ and $s_{2}$ due to rounding. The impact of this on $\operatorname{E}[M_{i}+N_{1}]$, $i=1,2$ will usually be negligible, and thus $\xi$ could be taken as 0 with good approximation. However, choosing $\xi=1$ (together with appropriate values of $r_{1}$ and $r_{2}$) will be useful in Section 2.1.3 to obtain upper bounds on $\operatorname{E}[M_{i}+N_{i}]$.
The terms $\operatorname{E}[X]$ and $\operatorname{E}[1/X]$ in (43) and (44) can be obtained as follows. Since $M_{1}$ and $M_{2}$ in (21) are independent,
$\operatorname{E}[X]=\operatorname{E}[M_{2}-\epsilon_{2}]\operatorname{E}\left[\frac{1}{M_{1}-\epsilon_{1}}\right].$ (45)
From (11),
$\operatorname{E}[M_{2}-\epsilon_{2}]=\frac{r_{2}}{p_{2}}-\epsilon_{2}=\frac{r_{2}}{p_{2}}\left(1-\frac{\epsilon_{2}p_{2}}{r_{2}}\right).$ (46)
As regards $\operatorname{E}[1/(M_{1}-\epsilon_{1})]$, it is easy to see that for $\epsilon_{1}\in(0,1)$
$\frac{1}{M_{1}}<\frac{1}{M_{1}-\epsilon_{1}}<\frac{\epsilon_{1}}{M_{1}-1}+\frac{1-\epsilon_{1}}{M_{1}},$ (47)
where $\operatorname{E}[1/(M_{1}-1)]$ is given by (12) and $\operatorname{E}[1/M_{1}]$ is bounded as [17, section II]
$\frac{p_{1}}{r_{1}-1}\left(1-\frac{p_{1}}{r_{1}-2}\right)<\operatorname{E}\left[\frac{1}{M_{1}}\right]<\frac{p_{1}}{r_{1}-1}\left(1-\frac{p_{1}}{r_{1}-1+p_{1}}\right).$ (48)
Therefore,
$\frac{p_{1}}{r_{1}-1}\left(1-\frac{p_{1}}{r_{1}-2}\right)<\operatorname{E}\left[\frac{1}{M_{1}-\epsilon_{1}}\right]<\frac{p_{1}}{r_{1}-1}\left(1-\frac{(1-\epsilon_{1})p_{1}}{r_{1}-1+p_{1}}\right).$ (49)
It follows from (45), (46) and (49) that $\operatorname{E}[X]$ is bounded for $\epsilon_{1},\epsilon_{2}\in(0,1)$ as
$\left(1-\frac{\epsilon_{2}p_{2}}{r_{2}}\right)\left(1-\frac{p_{1}}{r_{1}-2}\right)<\frac{r_{1}-1}{r_{2}\theta}\operatorname{E}[X]<\left(1-\frac{\epsilon_{2}p_{2}}{r_{2}}\right)\left(1-\frac{(1-\epsilon_{1})p_{1}}{r_{1}-1+p_{1}}\right).$ (50)
By analogous arguments, $\operatorname{E}[1/X]$ satisfies the following bound for $\epsilon_{1},\epsilon_{2}\in(0,1)$:
$\left(1-\frac{\epsilon_{1}p_{1}}{r_{1}}\right)\left(1-\frac{p_{2}}{r_{2}-2}\right)<\frac{(r_{2}-1)\theta}{r_{1}}\operatorname{E}\left[\frac{1}{X}\right]<\left(1-\frac{\epsilon_{1}p_{1}}{r_{1}}\right)\left(1-\frac{(1-\epsilon_{2})p_{2}}{r_{2}-1+p_{2}}\right).$ (51)
A convenient choice for $\epsilon_{1}$ and $\epsilon_{2}$, which will be assumed in the sequel, is $\epsilon_{1}=\epsilon_{2}=1/2$. Then, in view of (50) and (51), $\operatorname{E}[X]$ and $\operatorname{E}[1/X]$ are well approximated, especially for small $p_{1}$, $p_{2}$, as
$\operatorname{E}[X]$ $\approx\frac{r_{2}\theta}{r_{1}-1},$ (52)
$\operatorname{E}\left[\frac{1}{X}\right]$ $\approx\frac{r_{1}}{(r_{2}-1)\theta}.$ (53)
Substituting (52) and (53) into (43) and (44) yields
$\operatorname{E}[M_{1}+N_{1}]$ $\approx\frac{a_{1}+r_{1}+\xi}{p_{1}}+\frac{b_{1}r_{2}}{(r_{1}-1)p_{2}},$ (54)
$\operatorname{E}[M_{2}+N_{2}]$ $\approx\frac{a_{2}+r_{2}+\xi}{p_{2}}+\frac{b_{2}r_{1}}{(r_{2}-1)p_{1}}.$ (55)
By means of (54) and (55), a desired ratio of average sample sizes can be approximately achieved. Specifically, (17) will hold for all $p_{1}$, $p_{2}$ if
$a_{1}+r_{1}+\xi$ $=\frac{\lambda b_{2}r_{1}}{r_{2}-1},$ (56)
$\lambda(a_{2}+r_{2}+\xi)$ $=\frac{b_{1}r_{2}}{r_{1}-1}.$ (57)
This equation system can be expressed, making use of (39)–(42), as
$\frac{1}{A}+r_{1}+\mu_{1}+\xi$ $=\frac{\lambda r_{1}}{\gamma(r_{2}-1)}\left(\frac{1}{A}+\delta_{1}+\mu_{1}\right),$ (58)
$\frac{\lambda}{\gamma}\left(\frac{1}{A}+r_{2}+\mu_{2}+\xi\right)$ $=\frac{r_{2}}{r_{1}-1}\left(\frac{1}{A}+\delta_{2}+\mu_{2}\right).$ (59)
Given $A$, $\lambda$, $\xi$, and with $\mu_{1}$, $\mu_{2}$, $\mu_{12}$ known, the design parameters $r_{1}$, $r_{2}$, $\delta_{1}$, $\delta_{2}$, $\gamma$ provide degrees of freedom that help meet (58) and (59). Namely, imposing the relationships
$r_{2}=r_{1}+\mu_{1}-\mu_{2},\qquad\delta_{2}=\delta_{1}+\mu_{1}-\mu_{2},$ (60)
it is straightforward to solve for $\gamma$ and $\delta_{1}$ in (58) and (59):
$\gamma$ $=\lambda\sqrt{\frac{r_{1}(r_{1}-1)}{r_{2}(r_{2}-1)}}=\lambda\sqrt{\frac{r_{1}(r_{1}-1)}{(r_{1}+\mu_{1}-\mu_{2})(r_{1}+\mu_{1}-\mu_{2}-1)}},$ (61)
$\begin{split}\delta_{1}&=\left(\frac{1}{A}+r_{1}+\mu_{1}+\xi\right)\frac{\lambda(r_{1}-1)}{\gamma\,(r_{1}+\mu_{1}-\mu_{2})}-\frac{1}{A}-\mu_{1}\\ &=\left(\frac{1}{A}+r_{1}+\mu_{1}+\xi\right)\sqrt{\frac{(r_{1}-1)(r_{1}+\mu_{1}-\mu_{2}-1)}{r_{1}(r_{1}+\mu_{1}-\mu_{2})}}-\frac{1}{A}-\mu_{1}.\end{split}$ (62)
The parameters $\mu_{1}$, $\mu_{2}$ and $\mu_{12}$ for the RR estimator, given by (30), satisfy
$0\leq\mu_{1},\mu_{2}\leq 2,\quad 0\leq\mu_{12},$ (63)
and this will also be the case for the other estimators to be described later. Combining (63) with (23), (60) and (62), it follows that $\delta_{i}>-1/A-\mu_{i}$, $i=1,2$. On the other hand, (29) and (31) imply that $s_{i}-\mu_{i}>1/A$. From these two inequalities it stems that $\delta_{1}$ and $\delta_{2}$ as computed from (60) and (62) fulfill the requirement (22), and are thus valid. Making use of (60), the expressions (34)–(36) for $s_{1}$, $s_{2}$ are more conveniently written as
$s_{1}$ $=\frac{\gamma X(A(\delta_{1}+\mu_{1})+1)-A(\delta_{1}-\mu_{1})+1+\sqrt{D}}{2A},$ (64)
$s_{2}$ $=\frac{s_{1}+\delta_{1}}{\gamma X}-\delta_{1}-\mu_{1}+\mu_{2},$ (65)
$\begin{split}D&=(\gamma X(A(\delta_{1}+\mu_{1})+1)-A(\delta_{1}-\mu_{1})+1)^{2}\\ &\quad-4A\left(\gamma X((A\mu_{1}+1)(\delta_{1}+\mu_{1})+\mu_{1}-\mu_{12})-(A\mu_{1}+1)\delta_{1}\right).\end{split}$ (66)
Substituting (39)–(42) and (60)–(62) into (54) and (55) yields
$\operatorname{E}[M_{1}+N_{1}]$ $\approx\left(\frac{1}{A}+r_{1}+\mu_{1}+\xi\right)\left(\frac{1}{p_{1}}+\frac{\lambda}{p_{2}}\right),$ (67)
$\operatorname{E}[M_{2}+N_{2}]$ $\approx\left(\frac{1}{A}+r_{1}+\mu_{1}+\xi\right)\left(\frac{1}{\lambda p_{1}}+\frac{1}{p_{2}}\right).$ (68)
It will be beneficial to use normalized versions of the average sample sizes, $\operatorname{E}[M_{i}+N_{i}]\sqrt{p_{1}p_{2}}$, so that the expressions depend on $p_{1}$ and $p_{2}$ only through their ratio $\theta$. Defining
$\phi=\sqrt{p_{1}p_{2}},$ (69)
the approximations (67) and (68) can be written as
$\operatorname{E}[M_{1}+N_{1}]\phi$ $\approx\left(\frac{1}{A}+r_{1}+\mu_{1}+\xi\right)\left(\frac{1}{\sqrt{\lambda\theta}}+\sqrt{\lambda\theta}\right)\sqrt{\lambda},$ (70)
$\operatorname{E}[M_{2}+N_{2}]\phi$ $\approx\left(\frac{1}{A}+r_{1}+\mu_{1}+\xi\right)\left(\frac{1}{\sqrt{\lambda\theta}}+\sqrt{\lambda\theta}\right)\frac{1}{\sqrt{\lambda}}.$ (71)
The average number of samples $\operatorname{E}[M_{i}+N_{i}]$, $i=1,2$ is the sum of two terms inversely proportional to $p_{1}$ and $p_{2}$, according to (67) and (68); and it is $\lambda$ times more sensitive to $p_{2}$ than to $p_{1}$. That is, the parameter of the population for which a smaller average sample size is desired (as specified by (17)) has a stronger influence on both average sample sizes. It is also noteworthy that, for $\lambda$ fixed, (70) and (71) are minimized when $\lambda\theta=1$, which according to (17) and (18) corresponds to $r_{1}+\operatorname{E}[s_{1}]\approx r_{2}+\operatorname{E}[s_{2}]$.
2.1.3 Analysis of the approximation and bounds on average sample sizes
Based on an analysis of the approximation error in (37) and (38), this section obtains bounds on the average sample sizes and selects appropriate values for $r_{1}$ and $r_{2}$.
The error in approximating (34) and (35) by (37) and (38) vanishes when $s_{1}$ is an affine function of $X$ and $s_{2}$ is an affine function of $1/X$ (see Figure 2). This happens when $A$, $r_{1}$ and $\xi$ satisfy a certain relationship, as discussed next. Expanding (66) and collecting terms, $D$ can be written as $d_{2}X^{2}+d_{1}X+d_{0}$ with
$d_{2}$ $=\gamma^{2}(A(\delta_{1}+\mu_{1})+1)^{2},$ (72)
$d_{1}$ $=2\gamma(-A^{2}(\delta_{1}+\mu_{1})^{2}-2A(\delta_{1}+\mu_{1}-\mu_{12})+1),$ (73)
$d_{0}$ $=(A(\delta_{1}+\mu_{1})+1)^{2}.$ (74)
The condition that $s_{1}$ and $s_{2}$ are affine functions of $X$ and $1/X$ is equivalent to $D$ being a perfect square with respect to $X$, that is,
$d_{1}=\pm 2\sqrt{d_{2}d_{0}}.$ (75)
The equality (75) with negative sign has the solution $A=-1/\mu_{12}$ if $\mu_{12}\neq 0$. With the estimators considered in this paper, $\mu_{12}$ is either 0 or positive, and this results in no solution or a negative solution for $A$, which is not valid. On the other hand, (75) with positive sign is satisfied if and only if $A(\delta_{1}+\mu_{1})^{2}+2(\delta_{1}+\mu_{1})-\mu_{12}=0.$ Taking into account (62), this can be expressed as $c(A,r_{1},\xi)=0$ with
$\begin{split}c(A,r_{1},\xi)&=A\left(\left(\frac{1}{A}+r_{1}+\mu_{1}+\xi\right)\sqrt{\frac{(r_{1}-1)(r_{1}+\mu_{1}-\mu_{2}-1)}{r_{1}(r_{1}+\mu_{1}-\mu_{2})}}-\frac{1}{A}\right)^{2}\\ &\quad+2\left(\left(\frac{1}{A}+r_{1}+\mu_{1}+\xi\right)\sqrt{\frac{(r_{1}-1)(r_{1}+\mu_{1}-\mu_{2}-1)}{r_{1}(r_{1}+\mu_{1}-\mu_{2})}}-\frac{1}{A}\right)-\mu_{12}.\end{split}$ (76)
Furthermore, it is easy to see that $c(A,r_{1},\xi)$ being positive (negative) implies that $s_{1}$ and $s_{2}$ given by (64)–(66) are both convex (concave) functions of $X$ and $1/X$ respectively, which ensures that they are smaller (greater) than their approximations. Thus $c(A,r_{1},\xi)$ will be referred to as curvature function.
This characterization of the error allows transforming (70) and (71) into upper bounds. Since rounding $s_{1}$ and $s_{2}$ never increases their values by $1$ or more, taking $\xi=1$ and choosing $r_{1}$ such that $c(A,r_{1},\xi)\geq 0$ implies that approximations (43) and (44) become inequalities:
$\operatorname{E}[M_{1}+N_{1}]$ $<\frac{a_{1}+r_{1}+1}{p_{1}}+\frac{b_{1}}{p_{1}}\operatorname{E}[X],$ (77)
$\operatorname{E}[M_{2}+N_{2}]$ $<\frac{a_{2}+r_{2}+1}{p_{2}}+\frac{b_{2}}{p_{2}}\operatorname{E}\left[\frac{1}{X}\right].$ (78)
From (50) and (51) it is clear that $\operatorname{E}[X]<r_{2}\theta/(r_{1}-1)$ and $\operatorname{E}[1/X]<r_{1}/((r_{2}-1)\theta)$. Substituting into (77) and (78) and using (39)–(42) and (60)–(62) yields
$\operatorname{E}[M_{1}+N_{1}]\phi$ $<\left(\frac{1}{A}+r_{1}+\mu_{1}+1\right)\left(\frac{1}{\sqrt{\lambda\theta}}+\sqrt{\lambda\theta}\right)\sqrt{\lambda},$ (79)
$\operatorname{E}[M_{2}+N_{2}]\phi$ $<\left(\frac{1}{A}+r_{1}+\mu_{1}+1\right)\left(\frac{1}{\sqrt{\lambda\theta}}+\sqrt{\lambda\theta}\right)\frac{1}{\sqrt{\lambda}}.$ (80)
These bounds hold for $\xi=1$ and for any $r_{1}$ such that $c(A,r_{1},1)\geq 0$. Considering the restriction (23), the best (i.e. lowest) upper bounds are obtained by choosing $r_{1}$ as
$r_{1}=\min\{r=3,4,5,\ldots\mid c(A,r,1)\geq 0\}.$ (81)
Thus the values $\xi=1$ and $r_{1}$ as in (81) will be used for the estimator.
Figure 3: Pairs $(A,r_{1})$ for which $c(A,r_{1},1)=0$
The condition $c(A,r_{1},\xi)=0$ can be written, from (76), as
$\begin{split}&(r_{1}+\mu_{1}+\xi)^{2}(r_{1}-1)(r_{1}+\mu_{1}-\mu_{2}-1)A^{2}\\ &\quad+\left(2(r_{1}+\mu_{1}+\xi)(r_{1}-1)(r_{1}+\mu_{1}-\mu_{2}-1)-\mu_{12}r_{1}(r_{1}+\mu_{1}-\mu_{2})\right)A\\ &\quad+1-2r_{1}-\mu_{1}+\mu_{2}=0.\end{split}$ (82)
Considering $r_{1}$ as given, this is a quadratic equation in $A$ with a single positive solution. Figure 3 shows the resulting curve for RR, i.e. for $\mu_{1}$, $\mu_{2}$, $\mu_{12}$ as in (30), with $\xi=1$. (The figure also contains curves for other estimators, to be presented in following sections.) For the pairs $(A,r_{1})$ in this curve the approximations (37) and (38) are exact. Note that $r_{1}$ is considered as a continuous variable for clarity of the representation, but only the points with integer $r_{1}$ (marked with dots in the graph) are feasible. Furthermore, it is seen from (76) that $c(A,r_{1},1)$ increases with $r_{1}$, and therefore the region above (below) the curve corresponds to $c(A,r_{1},1)$ positive (negative). Thus, for a given $A$, once the value of $r_{1}$ for which $c(A,r_{1},1)=0$ is known, rounding this up gives the minimum integer $r_{1}$ such that $c(A,r_{1},1)\geq 0$. That is, (81) is the integer on or immediately above the curve in Figure 3, limited from below by $3$.
With the values selected for $\gamma$, $\delta_{1}$, $\delta_{2}$, $\epsilon_{1}$, $\epsilon_{2}$, $\xi$, $r_{1}$, $r_{2}$, the RR estimator is completely specified. The estimation procedure is summarized in Algorithm 1 (see Appendix A), together with a list of its properties, some of which will be derived in the remainder of this section. (The algorithm also includes the LRR case, to be presented in Section 3.)
The performance of the RR estimator is evaluated by means of Monte Carlo simulations in the following. For each combination of parameters a simulation is run consisting of $10^{6}$ realizations of the estimator. The empirical MSE and average numbers of samples are computed from the simulation (i.e. expectation is replaced by sample mean), and then they are compared with the theoretical expressions.
Figure 4 shows the bounds (79) and (80) for the normalized average numbers of samples with $r_{1}$ given by (81) (lines), as well as simulation results for the same $r_{1}$ (bold, red dots) and for all $r_{1}=3,\ldots,50$ (light, grey dots). The graphs consider several combinations of $\lambda$, $\theta$ and $\phi$, with variable $A$. As a reference, (81) gives values $3,\ldots,31$ for the displayed range of $A$.
(a) $\lambda=1$, $\theta=1$, $\phi=0.01$ ($p_{1}=p_{2}=0.01$)
(b) $\lambda=1$, $\theta=16$, $\phi=0.01$ ($p_{1}=0.04$, $p_{2}=0.0025$)
(c) $\lambda=3$, $\theta=16$, $\phi=0.01$ ($p_{1}=0.04$, $p_{2}=0.0025$)
Line: bound. Bold dots: simulation. Light dots: simulation, other values of $r_{1}$.
Figure 4: Normalized average sample sizes for RR with element sampling, varying $r_{1}$
(a) $\lambda=3$, $\theta=1/16$, $\phi=0.01$ ($p_{1}=0.0025$, $p_{2}=0.04$)
(b) $\lambda=1/3$, $\theta=1/16$, $\phi=0.01$ ($p_{1}=0.0025$, $p_{2}=0.04$)
(c) $\lambda=1/3$, $\theta=1/16$, $\phi=0.1$ ($p_{1}=0.025$, $p_{2}=0.4$)
Line: bound. Bold dots: simulation. Light dots: simulation, other values of $r_{1}$.
Figure 5: Normalized average sample sizes for RR with element sampling, varying $r_{1}$
The following observations can be made from Figure 4. For $r_{1}$ as in (81), simulated values are close to the theoretical bounds (and always below). When $A$ is not too large they are in fact very close, whereas for large $A$ the difference increases because $r_{1}$ is limited below by $3$. In addition, comparing with simulations for other values of $r_{1}$, (81) achieves average sample sizes equal or very close to their minima with respect to $r_{1}$. For $r_{1}$ given by (81), the ratio $\operatorname{E}[M_{1}+N_{1}]/\operatorname{E}[M_{2}+N_{2}]$ in the simulations is very close to $\lambda$ in all cases (for reference, note that the theoretical curves (79) and (80) are in the exact ratio $\lambda$); this will be analyzed with more detail later. The effect of $\phi$ on the normalized average sample sizes $\operatorname{E}[M_{i}+N_{i}]\phi$, $i=1,2$ is imperceptible (compare Figures 5(b) and 5(c)). The parameter $\theta$ has an impact on the average sample sizes, but not on their ratio, which remains close to $\lambda$ (compare Figures 4(a) and 4(b), or 4(c) and 5(a)). If both $\theta$ and $\lambda$ are replaced by their reciprocal values, $\operatorname{E}[M_{1}+N_{1}]$ and $\operatorname{E}[M_{2}+N_{2}]$ are simply swapped (see Figures 4(c) and 5(b)). This is in agreement with (70) and (71). Lastly, for $A$ small the average sample sizes are approximately inversely proportional to this parameter. Again, this can be observed in (70) and (71), where for $A$ small the term $1/A$ dominates the other summands in the first factor.
The theoretical curves in Figure 4 have small jump discontinuities (see for example Figure 4(c) near $A=0.07$), caused by the discrete character of $r_{1}$ and $r_{2}$. The jumps occur when the result of (81) changes by $1$. This effect is also present in the simulation results (bold dots), although less evident. Apart from this, for fixed $r_{1}$ the simulated results exhibit steep changes in certain locations, produced by the rounding applied to $s_{1}$ and $s_{2}$ (see for example the rightmost region for $\operatorname{E}[M_{2}+N_{2}]$ in Figure 4(c) with $r_{1}=3$, which in that region corresponds to the bold dots). These are only observable in the simulation results, because the theoretical curves do not explicitly model the rounding of $s_{1}$ and $s_{2}$; and they are not discontinuities, but short sections where the slope is large in absolute value, as will be discussed later.
Figure 6 compares the target $A$ with the relative MSE, $\operatorname{E}[(\hat{\theta}-\theta)^{2}]/\theta^{2}$, obtained from simulations in two specific cases, for $r_{1}$ given by (81). As seen, the relative MSE is always less than $A$, in accordance with (16). The difference between simulation and target is considerable when the latter is large. Again, this is a consequence of the limitation of $r_{1}$ in (81) to values not smaller than $3$. In addition, the difference increases slightly with $\phi$ (compare Figures 6(a) and 6(b)). This is, at least in part, explained by the fact that the uniform bound (28) becomes less tight as $p_{1}$ or $p_{2}$ approach $1$, as can be seen by comparing it with (27).
(a) $\lambda=1$, $\theta=16$, $\phi=0.1$
($p_{1}=0.4$, $p_{2}=0.025$)
(b) $\lambda=3$, $\theta=1/16$, $\phi=0.01$
($p_{1}=0.0025$, $p_{2}=0.04$)
Dashed line: target. Dots: simulation.
Figure 6: Relative MSE for RR with element sampling
The relative MSE, like the average sample sizes, has steep changes near certain values of $A$, caused by rounding $s_{1}$ and $s_{2}$. The effect is noticeable in the rightmost region of Figure 6, and is explained as follows. For large $A$, at least one of the two parameters $s_{1}$ and $s_{2}$ is small, according to (29) and (33), so as to make the relative MSE similar to (but smaller than) the target. Consequently that parameter, $s_{j}$, has a very narrow distribution before rounding, which implies that rounding it almost always produces the same integer value. This will typically be the next greater integer, because for small $s_{j}$ rounding causes a large variation in the error function, and thus rounding down is not likely to satisfy (31). In these conditions, if $A$ is increased by a small amount, the change in the distribution of $s_{j}$ before rounding, even if correspondingly small, can be sufficient to cause the rounded value to decrease by $1$, producing a significant change in $\operatorname{E}[(\hat{\theta}-\theta)^{2}]/\theta^{2}$, as well as in $\operatorname{E}[M_{j}+N_{j}]$.
As a specific example, for $\lambda=1$, $\theta=16$ (Figure 6(a)), consider $A=0.562$ (leftmost dot in the highest “plateau” of simulated values), for which (81) gives $r_{1}=3$. The simulation yields an average value of $1.856$ for $s_{2}$ before rounding, and a standard deviation of $0.082$. The value after rounding is $2$ with probability $0.962$ (and greater with probability $0.038$). The parameter $s_{1}$ before rounding takes larger values, with average $109.2$ and standard deviation $76.6$. Reducing $A$ to $0.501$ (next simulated value towards the left), which still corresponds to $r_{1}=3$, the average of $s_{2}$ before rounding becomes $2.087$, and the standard deviation is $0.094$. After rounding, $s_{2}$ now equals $3$ with probability $0.999$ (and is greater with probability $0.001$). Again, $s_{1}$ before rounding takes larger values, with mean and standard deviation $112.5$ and $77.9$ respectively. The change of the rounded value of $s_{2}$, from being $2$ to being $3$ with probabilities near $1$, causes the rightmost vertical gap in Figure 6(a), from a relative MSE approximately equal to $0.50$ down to $0.34$. In contrast, for small $A$ both $s_{1}$ and $s_{2}$ are large and the effect of rounding is less marked, because, on one hand, the distribution of each parameter is wider, with many possible rounded values; and, on the other hand, rounding a large value only causes a small variation in the relative MSE.
The same behavior can be seen in the average sample sizes. Thus, in Figure 4(b), which also corresponds to $\lambda=1$, $\theta=16$, the simulation results for $\operatorname{E}[M_{2}+N_{2}]$ with $r_{1}$ as in (81) (bold dots) show a vertical gap at the same horizontal position discussed in the preceding paragraph for the relative MSE. The effect is not discernible in the graph of $\operatorname{E}[M_{1}+N_{1}]$ because $s_{1}$ is large.
It follows from the above that the effect of rounding $s_{1}$ and $s_{2}$ is not a discontinuity, but rather a short section with large slope in $\operatorname{E}[(\hat{\theta}-\theta)^{2}]/\theta^{2}$, or in $\operatorname{E}[M_{i}+N_{i}]$, as a function of $A$. Indeed, since (64) and (65) are continuous functions of $A$, the distributions of $s_{1}$ and $s_{2}$ also vary continuously with $A$, and so do the relative MSE and average sample sizes. This can be seen in Figure 7(a), which is a detailed view of the rightmost part of Figure 6(a) with smaller spacing along the horizontal axis, and in Figure 7(b), which shows the corresponding values of $\operatorname{E}[M_{2}+N_{2}]$.
(a) Relative MSE
(b) $\operatorname{E}[M_{2}+N_{2}]\phi$
Dashed line: target. Solid line: approximation. Dots: simulation.
Figure 7: Detail of relative MSE and normalized average sample size for RR with element sampling, for large AA; $\lambda=1$, $\theta=16$, $\phi=0.1$ ($p_{1}=0.4$, $p_{2}=0.025$)
The ratio $\operatorname{E}[M_{1}+N_{1}]/\operatorname{E}[M_{2}+N_{2}]$ obtained from simulation is represented in Figure 8. As can be seen, it is in general close to $\lambda$, and very close for small or moderate values of $A$. The relatively larger deviations in the rightmost region of the graphs can again be attributed to the fact that for large $A$, since $r_{1}$ is limited from below by $3$ in (81), the value of $r_{1}$ does not result in $c(A,r_{1},\xi)$ close to 0. In any case, the deviation between the actual ratio and $\lambda$ is less than $11\%$ for all combinations of parameters shown in the figure, and much lower for $A$ small.
(a) $\phi=0.01$
(b) $\phi=0.1$
$\lambda=1/3$, $1$, 33 (bottom to top). $\theta=1/16$, $1$, $16$ (see legend).
Figure 8: Ratio of average sample sizes for RR with element sampling
2.1.4 Estimation efficiency
The efficiency of an unbiased estimator can be defined by comparing its variance against the minimum variance that can be achieved by any unbiased estimator with the same sample size, or with the same average sample size. For a fixed-size estimator based on independent observations from a single population, the minimum possible variance is given by the Cramér–Rao bound [11, 13]. This bound applies under certain regularity conditions on the distribution of the samples, which are satisfied when the observations are binary. For sequential estimators, there exists a generalization of the Cramér–Rao bound, obtained by Wolfowitz [22] see also 11, section 4.3. The actual bound is the same, except that the average sample size is used instead of a fixed sample size, and the regularity conditions are different.
For estimators based on independent observations from two populations, as considered in this paper, a variation of the Cramér–Rao bound can be applied in the fixed-size case [13, chapter 3]. However, to the author’s knowledge, there is no analogue of Wolfowitz’s result for sequential estimators when the observations are obtained from more than one population. Therefore, the efficiency of the RR estimator (and of the estimators to be presented in subsequent sections) will be defined by comparing its variance with the lowest variance that can be attained by any fixed-size estimator with the same average size for each population.
Consider fixed numbers $n_{1}$ and $n_{2}$ of independent binary samples taken from two populations with parameters $p_{1}$ and $p_{2}$. The number of successes observed from population $i=1,2$ follows a binomial distribution with parameters $n_{i}$ and $p_{i}$. Thus
$L(S_{1},S_{2};p_{1},p_{2})=\binom{n_{1}}{S_{1}}\binom{n_{2}}{S_{2}}p_{1}^{S_{1}}(1-p_{1})^{n_{1}-S_{1}}p_{2}^{S_{2}}(1-p_{2})^{n_{2}-S_{2}}$ (83)
is the probability of observing $S_{1}$ and $S_{2}$ successes from the two populations respectively. For a generic parameter $\zeta$ that is a function of $p_{1},p_{2}$, and an unbiased estimator $\hat{\zeta}$ of $\zeta$, the Cramér–Rao bound is [13, section 3.8]
$\operatorname{Var}[\hat{\zeta}]\geq\mathbf{J}\>\mathbf{F}^{-1}\,\mathbf{J}^{\intercal}$ (84)
where $\mathbf{J}=\left[\partial\zeta/\partial p_{1}\;\;\partial\zeta/\partial p_{2}\right]$ is the $1\times 2$ Jacobian vector and $\mathbf{F}$ is the $2\times 2$ Fisher information matrix, defined as
$F_{i,j}=-\operatorname{E}\left[\frac{\partial^{2}\log L(S_{1},S_{2};p_{1},p_{2})}{\partial p_{i}\partial p_{j}}\right],\quad i,j=1,2.$ (85)
This matrix is readily computed from (83) as $F_{i,i}=n_{i}/(p_{i}(1-p_{i}))$, $F_{i,j}=0$ for $i\neq j$ (which reflects the fact that observations from one population give no information about the other), and therefore (84) becomes
$\operatorname{Var}[\hat{\zeta}]\geq\left(\frac{\partial\zeta}{\partial p_{1}}\right)^{2}\frac{p_{1}(1-p_{1})}{n_{1}}+\left(\frac{\partial\zeta}{\partial p_{2}}\right)^{2}\frac{p_{2}(1-p_{2})}{n_{2}}.$ (86)
Equating $n_{i}$ to the average number of observations of population $i$ used by the considered estimator, $\operatorname{E}[M_{i}+N_{i}]$, the efficiency with element sampling $\eta_{\mathrm{el}}$ is obtained as
$\eta_{\mathrm{el}}=\frac{\displaystyle\left(\frac{\partial\zeta}{\partial p_{1}}\right)^{2}\frac{p_{1}(1-p_{1})}{\operatorname{E}[M_{1}+N_{1}]}+\left(\frac{\partial\zeta}{\partial p_{2}}\right)^{2}\frac{p_{2}(1-p_{2})}{\operatorname{E}[M_{2}+N_{2}]}}{\operatorname{Var}[\hat{\zeta}]}$. (87)
Particularizing (87) for RR, that is $\zeta=\theta$, with $\partial\theta/\partial p_{1}=\theta/p_{1}$, $\partial\theta/\partial p_{2}=-\theta/p_{2}$,
$\eta_{\mathrm{el}}=\frac{\displaystyle\frac{1-p_{1}}{\operatorname{E}[M_{1}+N_{1}]p_{1}}+\frac{1-p_{2}}{\operatorname{E}[M_{2}+N_{2}]p_{2}}}{\operatorname{Var}[\hat{\theta}]/\theta^{2}}.$ (88)
For $\xi=1$ and $r_{1}$ as in (81), substituting (16), (79) and (80) into (88) yields the following bound for the efficiency of the RR estimator with element sampling:
$\begin{split}\eta_{\mathrm{el}}>\frac{1}{1+A(r_{1}+\mu_{1}+1)}\left(1-\phi\frac{1/\sqrt{\lambda}+\sqrt{\lambda}}{1/\sqrt{\lambda\theta}+\sqrt{\lambda\theta}}\right).\end{split}$ (89)
Figure 9 shows the efficiency $\eta_{\mathrm{el}}$ obtained from Monte Carlo simulation. Its value is computed using (88) with $\operatorname{E}[M_{i}+N_{i}]$, $i=1,2$ replaced by the corresponding sample means and $\operatorname{Var}[\hat{\theta}$] replaced by the sample MSE. The bound (89) is also plotted. As seen in the figure, the efficiency is high for the values of the target $A$ commonly used in practice, and in particular it is close to $1$ for small $A$. For example, $A=0.04$, corresponding to a relative RMSE of $20\%$, gives values of $\eta_{\mathrm{el}}$ around $80\%$. The efficiency as a function of $A$ exhibits the same discontinuities and steep changes that have been identified for the relative MSE and for $\operatorname{E}[M_{i}+N_{i}]$. The simulation results deviate more from the theoretical bound for large $A$ (rightmost part of the graphs) or for large $\phi$ (Figure 9(b)). This is explained by the fact that in these conditions the relative MSE is considerably smaller than the target, and the average sample sizes are also considerably smaller than their bounds, as discussed previously. On the other hand, the bound is quite tight for small or moderate values of $A$ and $\phi$, which is precisely when it is most important to characterize $\eta_{\mathrm{el}}$ accurately, as sample sizes are large in that case.
It is easily seen that $\lim_{A\rightarrow 0}Ar_{1}^{2}=1$ for $A$ and $r_{1}$ related by (82), and thus also for $r_{1}$ obtained from (81). Combining this result with (89), it follows that $\eta_{\mathrm{el}}$ tends to $1$ when $A$ and $\phi$ tend to 0. This is in consonance with the values shown in Figure 9.
(a) $\lambda=1$, $\theta=1$, $\phi=0.01$
($p_{1}=p_{2}=0.01$)
(b) $\lambda=1$, $\theta=16$, $\phi=0.1$
($p_{1}=0.4$, $p_{2}=0.025$)
(c) $\lambda=3$, $\theta=1$, $\phi=0.01$
($p_{1}=p_{2}=0.01$)
(d) $\lambda=3$, $\theta=1/16$, $\phi=0.01$
($p_{1}=0.0025$, $p_{2}=0.04$)
Line: bound. Dots: simulation.
Figure 9: Efficiency for RR with element sampling
2.2 Group sampling
2.2.1 Estimation procedure
In group sampling, samples are taken in groups or batches, each of which contains $l_{1}$ and $l_{2}$ samples from the two populations respectively. In consequence, the numbers of samples from both populations must be the same integer multiple of $l_{1}$ and $l_{2}$, i.e. $Gl_{1}$ and $Gl_{2}$, where $G$ is the number of groups.
Let $\lambda$ be defined for group sampling as
$\lambda=\frac{l_{1}}{l_{2}}.$ (90)
This definition is consistent with that used for element sampling, because $\lambda$ still represents the ratio of average sample sizes. In addition, for group sampling $M_{i}$, $N_{i}$, $i=1,2$ will continue to refer to the numbers of samples that would result if the estimation procedure discussed thus far, using element sampling, were applied with $\lambda$ given by (90) (and therefore these variables do not correspond to the numbers of samples actually required with group sampling).
The estimation process with group sampling is as follows. Samples are used individually, following the same procedure as with element sampling; but are in fact taken in groups. As individual samples from either population become necessary, groups of samples are taken, each group providing $l_{1}$ and $l_{2}$ samples of the two populations. During this process, a group may provide more samples than needed for one or the two populations. In that case the “surplus” samples are stored for later use. When a sample from a given population is required, a new group is taken only if the surplus samples from that population have been exhausted. Once the procedure has finished, there may remain some stored samples, which will be discarded (and at least for one of the two populations, with index $j$, the number of discarded samples will be less than $l_{j}$).
Estimation with group sampling thus proceeds as in Section 2.1, but with an added “outer layer” that translates group sampling into element sampling as described. With the above definition of $M_{i}$ and $N_{i}$, it follows that the number of required groups is
$G=\max\left\{\left\lceil\frac{M_{1}+N_{1}}{l_{1}}\right\rceil,\left\lceil\frac{M_{2}+N_{2}}{l_{2}}\right\rceil\right\}.$ (91)
2.2.2 Average number of groups
The relevant measure of sample size with group sampling is the average number of groups, $\operatorname{E}[G]$. To make its characterization more tractable, it is helpful to approximate (91) by assuming $\phi$ small, which implies that $\operatorname{E}[M_{i}+N_{i}]$, $i=1,2$ are correspondingly large. In addition to simplifying computations, this is the most important case in practice, as argued previously. For large $M_{i}$ and $N_{i}$, the rounding operations in (91) can be removed with negligible error. Thus, defining
$\Delta=\frac{M_{1}+N_{1}}{l_{1}}-\frac{M_{2}+N_{2}}{l_{2}},$ (92)
and noting that $\max\{x,y\}=(x+y)/2+|x-y|/$2, it is possible to express $\operatorname{E}[G]$ as
$\operatorname{E}[G]\approx\operatorname{E}\left[\max\left\{\frac{M_{1}+N_{1}}{l_{1}},\frac{M_{2}+N_{2}}{l_{2}}\right\}\right]=\frac{\operatorname{E}[M_{1}+N_{1}]}{2l_{1}}+\frac{\operatorname{E}[M_{2}+N_{2}]}{2l_{2}}+\frac{\operatorname{E}\left[|\Delta|\right]}{2}.$ (93)
The terms $\operatorname{E}[M_{1}+N_{1}]$ and $\operatorname{E}[M_{2}+N_{2}]$ in (93) are approximately given by (70) and (71). Regarding $\operatorname{E}[|\Delta|]$, it is shown in Appendix B.1 that, for $\phi$ small,
$\operatorname{E}\left[|\Delta|\right]\approx\operatorname{E}\left[\left|\frac{r_{1}+s_{1}}{l_{1}p_{1}}-\frac{r_{2}+s_{2}}{l_{2}p_{2}}\right|\right].$ (94)
Define
$\tilde{\Delta}=\frac{r_{1}+s_{1}}{l_{1}p_{1}}-\frac{r_{2}+s_{2}}{l_{2}p_{2}}.$ (95)
Using approximations (37) and (38) for $s_{1},s_{2}$ and including the rounding term $\xi$ as in Section 2.1.2 (this has very little effect on the approximations obtained here, but it is done for consistency), (95) becomes
$\tilde{\Delta}\phi\approx\frac{a_{1}+r_{1}+\xi+b_{1}X}{l_{1}\sqrt{\theta}}-\frac{a_{2}+r_{2}+\xi+b_{2}/X}{l_{2}}\sqrt{\theta}.$ (96)
Let $X_{0}$ denote the value of $X$ for which the right-hand side of (96) equals 0, and let $Y=X/\theta$. Taking into account (90), $X_{0}$ is seen to be the only positive solution of
$b_{1}X_{0}^{2}+(a_{1}+r_{1}+\xi-\lambda\theta(a_{2}+r_{2}+\xi))X_{0}-\lambda\theta b_{2}=0.$ (97)
Making use of (39)–(42) and (60)–(62), this solution is obtained as $X_{0}=\theta Y_{0}$ with
$Y_{0}=\frac{\textstyle r_{1}+\mu_{1}-\mu_{2}}{\textstyle 2\lambda\theta(r_{1}-1)}\left(\lambda\theta-1+\sqrt{(\lambda\theta-1)^{2}+\frac{4\lambda\theta(r_{1}-1)(r_{1}+\mu_{1}-\mu_{2}-1)}{r_{1}(r_{1}+\mu_{1}-\mu_{2})}}\right).\rule$[-16.00008pt]{0.0pt}{0.0pt} (98)
As established in Appendix B.1, for $\phi$ small the variable $Y$ approximately follows a beta prime distribution with parameters $r_{2}$, $r_{1}$. Thus, from (94)–(96),
$\operatorname{E}[|\Delta|]\phi\approx\operatorname{E}[|\tilde{\Delta}|]\phi=\\ \int_{Y_{0}}^{\infty}\left(\frac{a_{1}+r_{1}+\xi+b_{1}\theta y}{l_{1}\sqrt{\theta}}-\frac{a_{2}+r_{2}+\xi+b_{2}/(\theta y)}{l_{2}}\sqrt{\theta}\right)f(y;r_{2},r_{1})\,\mathrm{d}y\\ -\int_{0}^{Y_{0}}\left(\frac{a_{1}+r_{1}+\xi+b_{1}\theta y}{l_{1}\sqrt{\theta}}-\frac{a_{2}+r_{2}+\xi+b_{2}/(\theta y)}{l_{2}}\sqrt{\theta}\right)f(y;r_{2},r_{1})\,\mathrm{d}y,$ (99)
with $f(y;r_{2},r_{1})$ defined by (9). Applying (10),
$\int_{0}^{Y_{0}}f(y;r_{2},r_{1})\,\mathrm{d}y=I\left(\frac{Y_{0}}{Y_{0}+1};r_{2},r_{1}\right)=1-\int_{Y_{0}}^{\infty}f(y;r_{2},r_{1})\,\mathrm{d}y.$ (100)
Similarly, combining (9) and (10) with the identities (5) and (6) gives
[…]
From (99)–(106),
$\begin{split}\operatorname{E}[|\Delta|]\phi&\approx\left(\frac{a_{1}+r_{1}+\xi}{l_{1}\sqrt{\theta}}-\frac{a_{2}+r_{2}+\xi}{l_{2}}\sqrt{\theta}\right)\left(1-2I\left(\frac{Y_{0}}{Y_{0}+1};r_{2},r_{1}\right)\right)\\ &\quad+\frac{b_{1}r_{2}\sqrt{\theta}}{(r_{1}-1)l_{1}}\left(1-2I\left(\frac{Y_{0}}{Y_{0}+1};r_{2}+1,r_{1}-1\right)\right)\\ &\quad-\frac{b_{2}r_{1}}{(r_{2}-1)l_{2}\sqrt{\theta}}\left(1-2I\left(\frac{Y_{0}}{Y_{0}+1};r_{2}-1,r_{1}+1\right)\right).\end{split}$ (107)
Using (39), (41), (56), (57), (60) and the identities (7) and (8) into (107) yields
$\begin{split}\operatorname{E}[|\Delta|]\phi&\approx 2\left(\frac{1}{A}+r_{1}+\mu_{1}+\xi\right)\frac{Y_{0}^{r_{1}+\mu_{1}-\mu_{2}-1}}{(Y_{0}+1)^{2r_{1}+\mu_{1}-\mu_{2}-1}\mathrm{B}(r_{1}+\mu_{1}-\mu_{2},r_{1})}\\ &\quad\cdot\left(\frac{1}{r_{1}l_{1}\sqrt{\theta}}+\frac{Y_{0}\sqrt{\theta}}{(r_{1}+\mu_{2}-\mu_{2})l_{2}}\right).\end{split}$ (108)
Substituting (70), (71) and (108) into (93), and using (90), the average number of required groups for the RR estimator is obtained as
$\operatorname{E}[G]\phi\approx\left(\frac{1}{A}+r_{1}+\mu_{1}+\xi\right)\left(\frac{1}{l_{1}\sqrt{\theta}}+\frac{\sqrt{\theta}}{l_{2}}\right.\\ \left.+\frac{Y_{0}^{r_{1}+\mu_{1}-\mu_{2}-1}}{(Y_{0}+1)^{2r_{1}+\mu_{1}-\mu_{2}-1}\mathrm{B}(r_{1}+\mu_{1}-\mu_{2},r_{1})}\left(\frac{1}{r_{1}l_{1}\sqrt{\theta}}+\frac{Y_{0}\sqrt{\theta}}{(r_{1}+\mu_{1}-\mu_{2})l_{2}}\right)\right),$ (109)
where $Y_{0}$ is given by (98).
Figure 10 represents the normalized average number of groups obtained from simulation, as well as its theoretical approximation (109). Overall, the graphs follow similar patterns to those observed for $\operatorname{E}[M_{i}+N_{i}]\phi$ with element sampling. The approximation is accurate for small or moderate $A$, and less so for large $A$. Discontinuities due to the choice of $r_{1}$ are minimally visible, and steep changes due to rounding $s_{1}$ and $s_{2}$ can be observed in the simulation results for large $A$. Comparing $\operatorname{E}[G]$ for different values of $\theta$ only makes sense when $l_{1}$ and $l_{2}$ are fixed; that is, between Figures 10(a) and 10(b) or between Figures 10(c) and 10(d). For $l_{1}$ and $l_{2}$ given, it can be seen from (109) that the minimum of $\operatorname{E}[G]$ with respect to $\theta$ does not necessarily occur when $\lambda\theta=1$, as was the case for element sampling.
(a) $l_{1}=1$, $l_{2}=1$, $\theta=1$, $\phi=0.01$
($p_{1}=0.01$, $p_{2}=0.01$)
(b) $l_{1}=1$, $l_{2}=1$, $\theta=16$, $\phi=0.01$
($p_{1}=0.04$, $p_{2}=0.0025$)
(c) $l_{1}=2$, $l_{2}=5$, $\theta=16$, $\phi=0.01$
($p_{1}=0.04$, $p_{2}=0.0025$)
(d) $l_{1}=2$, $l_{2}=5$, $\theta=1/16$, $\phi=0.01$
($p_{1}=0.0025$, $p_{2}=0.04$)
Line: approximation. Dots: simulation.
Figure 10: Normalized average number of groups for RR with group sampling
2.2.3 Estimation efficiency
The efficiency of the estimator is defined, as in Section 2.1.4, by comparing the estimation variance with the lowest variance that could be achieved by a fixed-size estimator with the same average numbers of samples, as given by the Cramér–Rao bound. With group sampling, the average number of samples required from population $i$ is $\operatorname{E}[G]l_{i}$, $i=1,2$. Equating $n_{i}$ to $\operatorname{E}[G]l_{i}$ in (86), the efficiency $\eta_{\mathrm{gr}}$ for an unbiased estimator of a generic parameter $\zeta$ with group sampling is obtained as
$\eta_{\mathrm{gr}}=\frac{\displaystyle\left(\frac{\partial\zeta}{\partial p_{1}}\right)^{2}\frac{p_{1}(1-p_{1})}{l_{1}}+\left(\frac{\partial\zeta}{\partial p_{2}}\right)^{2}\frac{p_{2}(1-p_{2})}{l_{2}}}{\operatorname{E}[G]\operatorname{Var}[\hat{\zeta}]}$. (110)
Particularizing for RR, i.e. for $\zeta=\theta$,
$\eta_{\mathrm{gr}}=\frac{(1-p_{1})/(l_{1}p_{1})+(1-p_{2})/(l_{2}p_{2})}{\operatorname{E}[G]\operatorname{Var}[\hat{\theta}]/\theta^{2}}.$ (111)
Approximating $\operatorname{Var}[\theta]/\theta^{2}\approx A$, the efficiency for RR with group sampling is expressed as
$\eta_{\mathrm{gr}}\approx\frac{(1-p_{1})/(l_{1}\sqrt{\theta})+(1-p_{2})\sqrt{\theta}/l_{2}}{A\operatorname{E}[G]\phi},$ (112)
where $\operatorname{E}[G]\phi$ is given by (109).
Simulation results for the efficiency are shown in Figure 11. The values are obtained using (111) with $\operatorname{E}[G]$ and $\operatorname{Var}[\hat{\theta}$] replaced by the sample mean and sample MSE. The approximation (112) is also displayed. The figure contains only two specific cases for brevity. The results are in general little sensitive to $\lambda$ and $\theta$; but, similarly to what was observed with element sampling, the theoretical approximation becomes more conservative for large $A$ or $\phi$. Comparing with Figure 9, group sampling is seen to be less efficient than element sampling, in accordance with the number of required groups being the maximum over the two populations, as given by (91). This causes an efficiency loss of approximately $0.15$ for values of $A$ in the range 0.010.010.10.1, and a less substantial loss for $A$ small.
(a) $l_{1}=1$, $l_{2}=1$, $\theta=16$, $\phi=0.1$
($p_{1}=0.4$, $p_{2}=0.025$)
(b) $l_{1}=3$, $l_{2}=1$, $\theta=1/16$, $\phi=0.01$
($p_{1}=0.0025$, $p_{2}=0.04$)
Line: approximation. Dots: simulation.
Figure 11: Efficiency for RR with group sampling
3 Estimation of log relative risk
This section describes the LRR estimator and analyzes its properties. The presentation, as will also be the case for subsequent estimators, follows the same logical course as in Section 2, but can be shorter, thanks to the similarities with RR. Section 3.1 addresses the estimation procedure with element sampling, and Section 3.2 considers group sampling.
3.1 Element sampling
The process for estimating the LRR, $\Theta=\log(p_{1}/p_{2})$, is analogous to that for RR: it consists of two stages, each of which applies IBS to each population. The first stage uses fixed IBS parameters, $r_{1}$, $r_{2}$, and the second uses IBS parameters $s_{1}$, $s_{2}$ computed from the results $M_{1}$, $M_{2}$ of the first stage by means of the variable $X$ defined in (21), with $\epsilon_{1}=\epsilon_{2}=1/2$. An unbiased estimation $\hat{\Theta}$ is computed from the second-stage results $N_{1}$, $N_{2}$, according to the expression given next; and the error requirement is in this case defined in terms of MSE (rather than relative MSE), or equivalently variance:
$\operatorname{Var}[\hat{\Theta}]\leq A.$ (113)
In the second stage, writing $\Theta=\log p_{1}-\log p_{2}$, the estimator for the logarithm of a probability described in Mendo [16, section 3] can be used for each of these two terms. Specifically, $H_{N_{i}-1}+H_{s_{i}-1}$, where $H_{n}$ is the $n$-th harmonic number defined in (3), is an unbiased estimator of $\log p_{i}$, with variance less than $1/(s_{i}-1)$ for any $p_{i}\in(0,1)$ [16]. Therefore,
$\hat{\Theta}=-H_{N_{1}-1}+H_{N_{2}-1}+H_{s_{1}-1}-H_{s_{2}-1}$ (114)
is a conditionally unbiased estimator of $\Theta$ given $s_{1},s_{2}$, which implies that it is also unconditionally unbiased; and, since the observations of the two populations are independent,
$\operatorname{Var}\left[\hat{\Theta}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]<\frac{1}{s_{1}-1}+\frac{1}{s_{2}-1}.$ (115)
Defining the error function for LRR as the right-hand side of (115), i.e. as (29) with
$\mu_{1}=1,\qquad\mu_{2}=1,\qquad\mu_{12}=0,$ (116)
and assuming (31), it follows that
$\operatorname{Var}\left[\hat{\Theta}^{2}\right]=\operatorname{E}\left[\operatorname{E}\left[\hat{\Theta}^{2}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]-\Theta^{2}<\operatorname{E}\left[e(s_{1},s_{2})+\Theta^{2}\right]-\Theta^{2}\leq A.$ (117)
That is, condition (31) ensures that (113) holds for any $p_{1},p_{2}\in(0,1)$.
As in the RR case, $s_{1}$ and $s_{2}$ are obtained from the equation system formed by (20) and (33), except with $\mu_{1},\mu_{2},\mu_{12}$ given by (116). The solution is (64)–(66). The values of $s_{1}$ and $s_{2}$ should then be rounded while satisfying (31). Using the first-order approximations (37)–(42), $\gamma$ and $\delta_{1}$ are obtained as in (61) and (62); and (22) holds. The average sample sizes are approximately given by (70) and (71).
The curvature function is defined in the same way as for RR, and the condition $c(A,r_{1},\xi)=0$ is expressed by (82). The fact that $\mu_{1}=\mu_{2}$ and $\mu_{12}=0$ for LRR implies that the positive solution of this equation has a simple expression,
$A=\frac{1}{(r_{1}+\mu_{1}+\xi)(r_{1}-1)}.$ (118)
In analogy with RR, the value $\xi=1$ is used for the LRR estimator, with $r_{1}$ given by (81), which in this case can be written explicitly as
$r_{1}=\max\left\{3,\left\lceil-1/2+\sqrt{(3/2)^{2}+1/A}\right\rceil\right\}.$ (119)
The pairs $(A,r_{1})$ determined by (118) for $\xi=1$ are shown in Figure 3. With $\xi=1$ and $r_{1}$ given by (119), the average sample sizes are bounded by (79) and (80).
Algorithm 1 (see Appendix A) describes the estimation procedure for LRR, as well as the properties of the estimator.
The simulation results for $\operatorname{E}[M_{i}+N_{i}]\phi$, $i=1,2$ with $\xi=1$ and varying $r_{1}$ are analogous to those shown for RR in Figure 4 (light dots), and are omitted for brevity. As in that case, with $r_{1}$ chosen as in (119) the bounds (79) and (80) are close to the actual average numbers of samples, and these approximately take their minimum values with respect to $r_{1}$. The simulated MSE, also omitted, has a similar behavior to that in Figure 6, with values always smaller than the target $A$, and very close to it unless $A$ is large.
Figure 12 shows, in two specific cases, the theoretical bounds (79) and (80) and simulated $\operatorname{E}[M_{i}+N_{i}]\phi$ for $\xi=1$ and with $r_{1}$ given by (119). The plotted values have a variation pattern similar to that observed for RR (Section 2.1.3), with jump discontinuities and steep changes; again, the latter are only visible in the simulation results, and most apparent for large $A$. The values are slightly lower than those for RR in the rightmost region of the graphs (compare Figures 12(a) and 12(b) with 5(a) and 5(b) respectively), due to the fact that $\mu_{1}$ is $1$ for LRR and $2$ for RR. The effect of $\mu_{1}$ is only appreciable for large $A$, as it stems from (79) and (80). The ratio between average sample sizes obtained from simulation is close to $\lambda$, as can be seen in Figure 12 (note that the ratio between the bounds is exactly $\lambda$). More specifically, the achieved ratio deviates from $\lambda$ by small or very small percentages, very similar to those in RR (results not shown).
(a) $\lambda=3$, $\theta=1/16$, $\phi=0.01$ ($p_{1}=0.0025$, $p_{2}=0.04$)
(b) $\lambda=1/3$, $\theta=1/16$, $\phi=0.01$ ($p_{1}=0.0025$, $p_{2}=0.04$)
Line: bound. Dots: simulation.
Figure 12: Normalized average sample sizes for LRR with element sampling
As for the efficiency with element sampling, particularizing (87) for $\zeta=\Theta$, with $\partial\Theta/\partial p_{1}=1/p_{1}$, $\partial\Theta/\partial p_{2}=-1/p_{2}$ and substituting (79), (80) and (113) results in the same lower bound (89) as for RR (although the obtained values will be slightly different because $\mu_{1}$ and $r_{1}$ are), and the efficiency approaches $1$ when $A$ and $\phi$ tend to 0. Figure 13 compares this bound with simulation results, considering only two specific cases for brevity. The efficiency values from simulation are computed as explained in Section 2.1.4. Results are similar to those for RR (Figures 9(b) and 9(d)), except that for large $A$ the efficiency is larger in LRR, as is the difference between simulation and bound.
(a) $\lambda=1$, $\theta=16$, $\phi=0.1$
($p_{1}=0.4$, $p_{2}=0.025$)
(b) $\lambda=3$, $\theta=1/16$, $\phi=0.01$
($p_{1}=0.0025$, $p_{2}=0.04$)
Line: bound. Dots: simulation.
Figure 13: Efficiency for LRR with element sampling
3.2 Group sampling
The estimation of LRR with group sampling uses, as in the RR case, a number of groups $G$ given by (91) in order to provide the necessary amounts of individual samples of the two populations. The procedure is the same as in Section 2.2.1: a new group is taken whenever a sample of either population is required and no surplus samples of that population are available from previous groups. At the end of the process, any leftover samples are discarded.
The analysis in Appendix B.1 shows that the approximation (94) is also valid for LRR; and $\operatorname{E}[G]$ is then given by (109), with $Y_{0}$ as in (98) (and with the values of $\mu_{1}$, $\mu_{2}$ corresponding to LRR). Simulation results for the average number of groups, omitted for brevity, are very similar to those for RR (Figure 10), except that for large $A$ they are slightly lower than in RR, for the same reason as with element sampling.
The efficiency with group sampling has the same approximate expression (112) as for RR, with $\operatorname{E}[G]$ computed as indicated above. The simulation results, shown in Figure 14, follow the same pattern observed for element sampling: the theoretical approximation is accurate when $A$ and $\phi$ are small, and conservative otherwise; and for large $A$ the efficiency deviates more from the theoretical curve than in RR (compare with Figure 11).
(a) $l_{1}=1$, $l_{2}=1$, $\theta=16$, $\phi=0.1$
($p_{1}=0.4$, $p_{2}=0.025$)
(b) $l_{1}=3$, $l_{2}=1$, $\theta=1/16$, $\phi=0.01$
($p_{1}=0.0025$, $p_{2}=0.04$)
Line: approximation. Dots: simulation.
Figure 14: Efficiency for LRR with group sampling
4 Estimation of odds ratio
4.1 Element sampling
Several approaches are conceivable to estimate the OR $\psi$ defined in (2). One method that could be employed is to estimate $p_{1}/p_{2}$ and $(1-p_{2})/(1-p_{1})$ separately, treating each as a RR and using the two-stage procedure described in Section 2. With this approach, the MSE in the estimation of $\psi$ depends on the errors in the two RR estimations; and the problem is how to distribute the target MSE between these two components so as to approximately achieve a desired ratio of average sample sizes. To this end, another sampling stage could be introduced before the RR estimations, but that would complicate the process. A better approach, which only requires two stages and results in good estimation efficiency, is based on estimating the odds $p_{1}/(1-p_{1})$ and $(1-p_{2})/p_{2}$ separately, using in each case the method in Mendo [16, section 2]. This is detailed next.
The estimation method consists, as in previous sections, of two stages with IBS. The second stage estimates $p_{1}/(1-p_{1})$ for the first population, and $(1-p_{2})/p_{2}$ for the second population. However, there are two differences with respect to the RR and LRR estimators. The first difference is that in the second stage two IBS procedures are used for each population. Given $s_{1}$, $s_{2}$ (which will be computed from the results of the first stage), IBS is applied to population $1$ to obtain $s_{1}$ successes, which requires $N^{\prime}_{1}$ samples. Then IBS is applied again to population $1$ to obtain $s_{1}-\alpha$ failures, with $\alpha=2$ (a different value of $\alpha$ will be used for LOR estimation). This requires $N^{\prime\prime}_{1}$ samples, for a total of $N_{1}=N^{\prime}_{1}+N^{\prime\prime}_{1}$ samples. For population $2$, IBS is applied to obtain $s_{2}-\alpha$ successes, which requires $N^{\prime}_{2}$ samples; and then to obtain $s_{2}$ failures, which requires $N^{\prime\prime}_{2}$ samples, for a total of $N_{2}=N^{\prime}_{2}+N^{\prime\prime}_{2}$ samples. The average numbers of samples used by the second stage are computed as follows. From (11), the conditional mean of $N_{i}$, $i=1,2$ given $s_{i}$ is
$\operatorname{E}[N_{1}\mid s_{1}]$ $=\frac{s_{1}}{p_{1}}+\frac{s_{1}-\alpha}{1-p_{1}}=\frac{s_{1}-\alpha p_{1}}{p_{1}(1-p_{1})},$ (120)
$\operatorname{E}[N_{2}\mid s_{2}]$ $=\frac{s_{2}-\alpha}{p_{2}}+\frac{s_{2}}{1-p_{2}}=\frac{s_{2}-\alpha(1-p_{2})}{p_{2}(1-p_{2})},$ (121)
where $\alpha=2$ for OR. Then, defining
$\bar{p}_{i}=p_{i}(1-p_{i}),$ (122)
and considering a generic $\alpha\geq 0$, it stems from (120) and (121) that $\operatorname{E}[N_{i}]$ is approximately inversely proportional to $\bar{p}_{i}$:
$\operatorname{E}[N_{1}]$ $=\operatorname{E}\left[\operatorname{E}[N_{1}\mid s_{1}]\right]=\frac{\operatorname{E}[s_{1}]-\alpha p_{1}}{\bar{p}_{1}}\leq\frac{\operatorname{E}[s_{1}]}{\bar{p}_{1}},$ (123)
$\operatorname{E}[N_{2}]$ $=\operatorname{E}\left[\operatorname{E}[N_{2}\mid s_{2}]\right]=\frac{\operatorname{E}[s_{2}]-\alpha(1-p_{2})}{\bar{p}_{2}}\leq\frac{\operatorname{E}[s_{2}]}{\bar{p}_{2}}.$ (124)
Using the same ideas as for the RR estimator (Section 2.1.1), it can be seen that to approximately achieve a given ratio of average sample sizes, considering only the samples used by the second stage for the moment, $\operatorname{E}[s_{1}]/\operatorname{E}[s_{2}]$ should be roughly proportional to
$\bar{\theta}=\frac{\bar{p}_{1}}{\bar{p}_{2}}.$ (125)
In view of this, the second difference from previous estimators is that the first stage in this case needs to use samples with parameters $\bar{p}_{i}$, $i=1,2$, rather than $p_{i}$, so as to acquire information about $\bar{\theta}$. Specifically, for $i=1,2$, the first stage applies IBS with $r_{i}$ successes to a sequence of samples with parameter $\bar{p}_{i}$. Denoting the number of samples used from this sequence by $\bar{M}_{i}$, it is clear from (11) that $\operatorname{E}[\bar{M}_{i}]=r_{i}/\bar{p}_{i}$. The samples with parameter $\bar{p}_{i}$ must be generated from samples with parameter $p_{i}$. A simple, efficient procedure for this will be given later, and it will be shown that with this procedure the total number of samples with parameter $p_{i}$ required by the first stage, $M_{i}$, has an average equal to $3r_{i}/(2\bar{p}_{i})$. Therefore, considering both sampling stages, the average numbers of samples from the two populations satisfy
$\frac{\operatorname{E}[M_{1}+N_{1}]}{\operatorname{E}[M_{2}+N_{2}]}=\frac{3r_{1}/2+\operatorname{E}[s_{1}]}{(3r_{2}/2+\operatorname{E}[s_{2}])\bar{\theta}}.$ (126)
This is analogous to (18), and in consequence $s_{1}$ and $s_{2}$ can be chosen using a similar approach as for RR, described next.
The first stage produces the variables $\bar{M}_{1}$ and $\bar{M}_{2}$, from which $X$ is defined as
$X=\frac{\bar{M}_{2}-\epsilon_{2}}{\bar{M}_{1}-\epsilon_{1}},$ (127)
where $\epsilon_{1}=\epsilon_{2}=1/2$ as in Section 2. For a target relative MSE given by $A$, the second-stage IBS parameters $s_{1}$ and $s_{2}$ are obtained from $X$ by solving the equation system formed by (20) and (33), where the values of the design parameters $\gamma$, $\delta_{1}$ and $\delta_{2}$ are yet to be specified; and the solutions will then have to be rounded to integer values, as usual. Once $s_{1}$, $s_{2}$ are known, the second stage is carried out, from which $N^{\prime}_{1}$, $N^{\prime\prime}_{1}$, $N^{\prime}_{2}$, $N^{\prime\prime}_{2}$ are obtained. According to Mendo [16, section 2], and taking into account that $\alpha=2$,
$\frac{(s_{1}-1)\,N^{\prime\prime}_{1}}{(s_{1}-2)\,(N^{\prime}_{1}-1)}$
is a conditionally unbiased estimator of $p_{1}/(1-p_{1})$ given $s_{1}$;
$\frac{(s_{2}-1)\,N^{\prime}_{2}}{(s_{2}-2)\,(N^{\prime\prime}_{2}-1)}$
is a conditionally unbiased estimator of $(1-p_{2})/p_{2}$ given $s_{2}$; and for $s_{1},s_{2}\geq 3$ the conditional variances of these estimators are respectively less than
$\frac{p_{1}^{2}}{(s_{1}-2)(1-p_{1})^{2}}\left(1-\frac{\bar{p}_{1}}{s_{1}-2+2p_{1}}\right)$
and
$\frac{(1-p_{2})^{2}}{(s_{2}-2)p_{2}^{2}}\left(1-\frac{\bar{p}_{2}}{s_{2}-2p_{2}}\right).$
Therefore, since the observations are independent,
$\hat{\psi}=\frac{(s_{1}-1)(s_{2}-1)\,N^{\prime\prime}_{1}N^{\prime}_{2}}{(s_{1}-2)(s_{2}-2)\,(N^{\prime}_{1}-1)(N^{\prime\prime}_{2}-1)}$ (128)
is an unbiased estimator of $\psi=p_{1}(1-p_{2})/(p_{2}(1-p_{1}))$, and for $s_{1},s_{2}\geq 3$
$\frac{\operatorname{E}\left[\hat{\psi}^{2}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]}{\psi^{2}}\leq\left(\frac{1}{s_{1}-2}\left(1-\frac{\bar{p}_{1}}{s_{1}-2+2p_{1}}\right)+1\right)\left(\frac{1}{s_{2}-2}\left(1-\frac{\bar{p}_{2}}{s_{2}-2p_{2}}\right)+1\right),$ (129)
which implies that, for any $p_{1},p_{2}\in(0,1)$,
$\frac{\operatorname{E}\left[\hat{\psi}^{2}\mid s_{1},s_{2}\right]}{\psi^{2}}<\frac{1}{s_{1}-2}+\frac{1}{s_{2}-2}+\frac{1}{(s_{1}-2)(s_{2}-2)}+1.$ (130)
Based on (130), the error function $e(s_{1},s_{2})$ for OR is defined as in (29) with
$\mu_{1}=2,\qquad\mu_{2}=2,\qquad\mu_{12}=1;$ (131)
and then, by the same reasoning as in (32), if the rounded values of $s_{1}$ and $s_{2}$ satisfy (31) this guarantees that $\operatorname{Var}[\hat{\psi}]/\psi^{2}<A$ for any $p_{1},p_{2}\in(0,1)$.
The procedure to generate samples with parameter $\bar{p}_{i}$ is as follows. Taking two samples with parameter $p_{i}$ as inputs is clearly sufficient to produce a sample with parameter $\bar{p}_{i}$. Namely, one possible criterion is (a) to output success if and only if the first input is a success and the second is a failure. However, if the first input happens to be a failure the second input need not be observed. This occurs with probability $1-p_{i}$. Alternatively, (b) the output can be defined to be success if the first input is a failure and the second is a success. In this case, the second input is not needed if the first is a success, which occurs with probability $p_{i}$. If criteria (a) or (b) are randomly chosen with equal probabilities, the average number of inputs required to produce an output is $1+(1-p_{i})/2+p_{i}/2=3/2$. This is an instance of a Bernoulli factory [14].
The $\bar{M}_{i}$ samples with parameter $\bar{p}_{i}$, $i=1,2$ required by the first stage are generated with the above method, requiring a total of $M_{i}$ samples of population $i$ as inputs. For the subsequent analysis it is necessary to characterize the relationship between $\operatorname{E}[M_{i}]$ and $\operatorname{E}[\bar{M}_{i}]$. It cannot be directly concluded from the preceding paragraph that $\operatorname{E}[M_{i}]/\operatorname{E}[\bar{M}_{i}]=3/2$, because the IBS stopping rule could conceivably introduce some deviation in this ratio. However, this equality turns out to be true. More generally, for an arbitrary Bernoulli factory, if the average number of inputs needed to produce an output is equal to some constant $\beta$, it can be seen that
$\operatorname{E}[M_{i}]=\beta\operatorname{E}[\bar{M}_{i}]=\frac{\beta r_{i}}{\bar{p}_{i}},$ (132)
where $\beta=3/2$ for the described method. To prove this, let $\beta_{\mathrm{s}}$ and $\beta_{\mathrm{f}}$ be the average number of inputs required to produce an output, conditioned on the output being success or failure respectively. Then
$\beta=\beta_{\mathrm{s}}\bar{p}_{i}+\beta_{\mathrm{f}}(1-\bar{p}_{i}).$ (133)
IBS with parameter $r_{i}$ is applied to the outputs of the Bernoulli factory, and consumes $\bar{M}_{i}$ of those outputs, of which $r_{i}$ are successes and $\bar{M}_{i}-r_{i}$ are failures. Therefore, using (11),
$\operatorname{E}[M_{i}]=\beta_{\mathrm{s}}r_{i}+\beta_{\mathrm{f}}\left(\operatorname{E}[\bar{M}_{i}]-r_{i}\right)=\beta_{\mathrm{s}}r_{i}+\frac{\beta_{\mathrm{f}}r_{i}(1-\bar{p}_{i})}{\bar{p}_{i}},$ (134)
which combined with (133) gives (132).
The characterization of $\operatorname{E}[M_{i}+N_{i}]$, $i=1,2$, as well as the ensuing selection of $\gamma$, $\delta_{1}$ and $\delta_{2}$, relies, as in the RR case, on using first-order approximations for $s_{1}$ and $s_{2}$ as functions of $X$ and $1/X$ respectively. Expressions (37)–(42) remain valid for OR, and (52) and (53) hold with $\theta$ replaced by $\bar{\theta}$:
$\operatorname{E}[X]$ $\approx\frac{r_{2}\bar{\theta}}{r_{1}-1},$ (135)
$\operatorname{E}\left[\frac{1}{X}\right]$ $\approx\frac{r_{1}}{(r_{2}-1)\bar{\theta}}.$ (136)
In addition, (54) and (55) have to be modified to take into account that the first stage uses samples with parameter $\bar{p}_{i}$, obtained by transforming samples with parameter $p_{i}$, $i=1,2$. Specifically, from (37), (38), (123), (124), (132), (135) and (136), and introducing the term $\xi$ to account for the effect of rounding $s_{1}$ and $s_{2}$,
$\operatorname{E}[M_{1}+N_{1}]$ $=\frac{3r_{1}/2+\operatorname{E}[s_{1}]-2p_{1}}{\bar{p}_{1}}\approx\frac{3r_{1}/2+\operatorname{E}[s_{1}]}{\bar{p}_{1}}\approx\frac{a_{1}+3r_{1}/2+\xi}{\bar{p}_{1}}+\frac{b_{1}r_{2}}{(r_{1}-1)\bar{p}_{2}},$ (137)
$\operatorname{E}[M_{2}+N_{2}]$ $=\frac{3r_{2}/2+\operatorname{E}[s_{2}]-2(1-2p_{2})}{\bar{p}_{2}}\approx\frac{a_{2}+3r_{2}/2+\xi}{\bar{p}_{2}}+\frac{b_{2}r_{1}}{(r_{2}-1)\bar{p}_{1}}.$ (138)
According to (39)–(42), (137) and (138), the condition (17) for the ratio of average sample sizes will be satisfied regardless of $p_{1},p_{2}$ if
$\frac{1}{A}+\frac{3r_{1}}{2}+\mu_{1}+\xi$ $=\frac{\lambda r_{1}}{\gamma(r_{2}-1)}\left(\frac{1}{A}+\delta_{1}+\mu_{1}\right),$ (139)
$\frac{\lambda}{\gamma}\left(\frac{1}{A}+\frac{3r_{2}}{2}+\mu_{2}+\xi\right)$ $=\frac{r_{2}}{r_{1}-1}\left(\frac{1}{A}+\delta_{2}+\mu_{2}\right).$ (140)
By analogy with (58) and (59), a simple solution to (139) and (140) is obtained if $3r_{1}/2+\mu_{1}=3r_{2}/2+\mu_{2}$ and $\delta_{1}+\mu_{1}=\delta_{2}+\mu_{2}$. The former condition is compatible with $r_{1},r_{2}\in\mathbb{N}$ only if $2(\mu_{1}-\mu_{2})/3\in\mathbb{Z}$. In particular, this holds if
$\mu_{1}=\mu_{2}.$ (141)
This is the case for OR estimation, and will also be true for LOR estimation (Section 5). Thus, in the sequel (141) will be assumed to hold. The two indicated conditions then reduce to
$r_{2}=r_{1},\qquad\delta_{2}=\delta_{1},$ (142)
which gives the solution to (139) and (140) as
$\gamma$ $=\lambda,$ (143)
$\delta_{1}$ $=\left(\frac{1}{A}+\frac{3r_{1}}{2}+\mu_{1}+\xi\right)\frac{r_{1}-1}{r_{1}}-\frac{1}{A}-\mu_{1}.$ (144)
An argument analogous to that used in Section 2.1.2 shows that (22) is satisfied.
The values of $s_{1}$, $s_{2}$ before rounding, obtained by solving (20) and (33) with $\delta_{1}$, $\delta_{2}$ and $\gamma$ as in (142)–(144), are again given by (64)–(66). The curvature function is defined as in (76) with the terms $1/A+r_{1}+\mu_{1}+\xi$ replaced by $1/A+3r_{1}/2+\mu_{1}+\xi$ and with $\mu_{1}=\mu_{2}$:
$\begin{split}c(A,r_{1},\xi)&=A\left(\left(\frac{1}{A}+\frac{3r_{1}}{2}+\mu_{1}+\xi\right)\frac{r_{1}-1}{r_{1}}-\frac{1}{A}\right)^{2}\\ &\quad+2\left(\left(\frac{1}{A}+\frac{3r_{1}}{2}+\mu_{1}+\xi\right)\frac{r_{1}-1}{r_{1}}-\frac{1}{A}\right)-\mu_{12},\end{split}$ (145)
from which the condition $c(A,r_{1},\xi)=0$ is
$\begin{split}&\left(\frac{3r_{1}}{2}+\mu_{1}+\xi\right)^{2}(r_{1}-1)^{2}A^{2}\\ &\quad+\left(2\left(\frac{3r_{1}}{2}+\mu_{1}+\xi\right)(r_{1}-1)^{2}-\mu_{12}r_{1}^{2}\right)A+1-2r_{1}=0.\end{split}$ (146)
Similarly to previous estimators, $\xi$ is set to 11 and $r_{i}$ is chosen as in (81), where $c(A,r_{1},\xi)$ is given by (145). Equivalently, the curve formed by the pairs $(A,r_{1})$ that solve (146) can be plotted, as shown in Figure 3; and then, for a given $A$, (81) corresponds to rounding up the value obtained from the curve, with a minimum of $3$. Defining
$\bar{\phi}=\sqrt{\bar{p}_{1}\bar{p}_{2}},$ (147)
this choice of $\xi$ and $r_{1}$ results, by analogy with Section 2.1.3, in the upper bounds
$\operatorname{E}[M_{1}+N_{1}]\bar{\phi}$ $<\left(\frac{1}{A}+\frac{3r_{1}}{2}+\mu_{1}+1\right)\left(\frac{1}{\sqrt{\lambda\bar{\theta}}}+\sqrt{\lambda\bar{\theta}}\right)\sqrt{\lambda},$ (148)
$\operatorname{E}[M_{2}+N_{2}]\bar{\phi}$ $<\left(\frac{1}{A}+\frac{3r_{1}}{2}+\mu_{1}+1\right)\left(\frac{1}{\sqrt{\lambda\bar{\theta}}}+\sqrt{\lambda\bar{\theta}}\right)\frac{1}{\sqrt{\lambda}}.$ (149)
The estimation procedure for OR is specified in Algorithm 2 (see Appendix A), where the properties of the estimator are also indicated. (The algorithm covers the LOR case as well, to be presented in Section 5.)
The relative MSE obtained from simulations is compared with the target $A$ in Figure 15. The simulation consists, as with previous estimators, of $10^{6}$ realizations for each combination of parameters. The difference between simulation results and target is seen to increase with $A$ and with $\phi$, as for RR. However, in OR the difference observed for large $\phi$ vanishes when $A$ is small, unlike in RR (compare the leftmost parts of Figures 6(a) and 15(a)). This is related to the fact that the factors $1-\bar{p}_{1}/(s_{1}-2+2p_{1})$ and $1-\bar{p}_{2}/(s_{2}-2p_{2})$ in (129), which are replaced by $1$ in the uniform bound (130), approach $1$ for large $s_{1}$, $s_{2}$. Thus when $A$ is small, which gives large values of $s_{1}$ and $s_{2}$, the uniform bound (130) is almost as good as (129). In RR, on the other hand, the corresponding factors are $(s_{1}-2)(1-p_{1})/(s_{1}-2+2p_{1})<1-p_{1}$ and $1-p_{2}$, as is seen comparing (27) and (28), and these do not tend to $1$ for large $s_{1}$, $s_{2}$. (In LRR the factors are more cumbersome, see Mendo [16, theorem 2]; but for large $s_{1}$, $s_{2}$ they tend to the same values $1-p_{1}$, $1-p_{2}$ as in RR.)
(a) $\lambda=1$, $\theta=16$, $\phi=0.1$
($p_{1}=0.4$, $p_{2}=0.025$)
(b) $\lambda=3$, $\theta=1/16$, $\phi=0.01$
($p_{1}=0.0025$, $p_{2}=0.04$)
Dashed line: target. Dots: simulation.
Figure 15: Relative MSE for OR with element sampling
Figure 16 shows simulation results for $\operatorname{E}[M_{i}+N_{i}]\bar{\phi}$, $i=1,2$. The values are very similar to those for $\operatorname{E}[M_{i}+N_{i}]\phi$ in RR (Figures 5(a) and 5(b)) except that for OR they are somewhat larger in the rightmost part of the graphs. This is due to the $3/2$ factor in OR, whose effect on the average sample sizes is only significant for large $A$.
(a) $\lambda=3$, $\theta=1/16$, $\phi=0.01$ ($p_{1}=0.0025$, $p_{2}=0.04$)
(b) $\lambda=1/3$, $\theta=1/16$, $\phi=0.01$ ($p_{1}=0.0025$, $p_{2}=0.04$)
Line: bound. Dots: simulation.
Figure 16: Normalized average sample sizes for OR with element sampling
The efficiency of the OR estimator with element sampling results from particularizing (87) to $\zeta=\psi$, with $\partial\psi/\partial p_{1}=\psi/\bar{p}_{1}$, $\partial\psi/\partial p_{2}=-\psi/\bar{p}_{2}$:
$\eta_{\mathrm{el}}=\frac{\displaystyle\frac{1}{\operatorname{E}[M_{1}+N_{1}]\bar{p}_{1}}+\frac{1}{\operatorname{E}[M_{2}+N_{2}]\bar{p}_{2}}}{\operatorname{Var}[\psi]/\psi^{2}}.$ (150)
Then, using (148) and (149) and considering that $\operatorname{Var}[\hat{\psi}]/\psi^{2}<A$, the efficiency can be bounded as
$\eta_{\mathrm{el}}>\frac{1}{1+A(3r_{1}/2+\mu_{1}+1)}.$ (151)
This is similar to the bound (89) for RR and LRR, the only differences being the $3/2$ factor in (151), arising from the Bernoulli factory, and the last factor in (89) (but note that the latter is approximately $1$ for small $\phi$). By the same arguments used for RR and LRR, $\eta_{\mathrm{el}}$ for OR approaches $1$ when $A$ tends to 0.
Figure 17 shows simulation results for $\eta_{\mathrm{el}}$, and compares them with the bound (151). In the simulation, $\eta_{\mathrm{el}}$ is computed as in previous sections, i.e. using sample averages in (150). The values are seen to be similar to those for RR and LRR, except that with large $A$ the efficiency is slightly lower for OR. Again, this is a consequence of the $3/2$ factor. As in RR and LRR, the difference between simulation and bound tends to be larger when $\phi$ is increased. However, in OR that difference vanishes for small $A$, unlike in the other cases (compare Figure 17(a) with Figure 9(b) or 13(a)). This agrees with the behavior of the MSE discussed earlier.
(a) $\lambda=1$, $\theta=16$, $\phi=0.1$
($p_{1}=0.4$, $p_{2}=0.025$)
(b) $\lambda=3$, $\theta=1/16$, $\phi=0.01$
($p_{1}=0.0025$, $p_{2}=0.04$)
Line: bound. Dots: simulation.
Figure 17: Efficiency for OR with element sampling
4.2 Group sampling
Group sampling for OR estimation consumes a number of groups $G$ given by (91), as with previous estimators, in order to provide the required amounts of samples $M_{i}+N_{i}$ of each population $i=1,2$; and $\operatorname{E}[G]$ can be approximately computed from (93). It is shown in Appendix B.2 that, for $\phi$ small, the term $\operatorname{E}[|\Delta|]$ in (93) can be expressed as
$\operatorname{E}\left[|\Delta|\right]\approx\operatorname{E}\left[\left|\frac{3r_{1}/2+s_{1}}{l_{1}\bar{p}_{1}}-\frac{3r_{2}/2+s_{2}}{l_{2}\bar{p}_{2}}\right|\right].$ (152)
Thus, proceeding as in Section 2.2.2 and taking into account (141),
$\begin{split}\operatorname{E}[G]\bar{\phi}&\approx\left(\frac{1}{A}+\frac{3r_{1}}{2}+\mu_{1}+\xi\right)\left(\frac{1}{l_{1}\sqrt{\bar{\theta}}}+\frac{\sqrt{\bar{\theta}}}{l_{2}}\right.\\ &\quad\left.+\frac{Y_{0}^{r_{1}-1}}{(Y_{0}+1)^{2r_{1}-1}\mathrm{B}(r_{1},r_{1})r_{1}}\left(\frac{1}{l_{1}\sqrt{\bar{\theta}}}+\frac{Y_{0}\sqrt{\bar{\theta}}}{l_{2}}\right)\right),\end{split}$ (153)
where
$Y_{0}=\frac{r_{1}}{2\lambda\bar{\theta}(r_{1}-1)}\left(\lambda\bar{\theta}-1+\sqrt{(\lambda\bar{\theta}-1)^{2}+\frac{4\lambda\bar{\theta}(r_{1}-1)^{2}}{r_{1}^{2}}}\right).$ (154)
The simulation results for $\operatorname{E}[G]\bar{\phi}$, not shown, are similar to those for $\operatorname{E}[G]\phi$ in RR and LRR, with the difference that for large $A$ the values are slightly greater in OR compared with those two cases, as it has been observed with element sampling.
The efficiency with group sampling is obtained particularizing (110) for $\zeta=\psi$ and approximating $\operatorname{Var}[\hat{\psi}]/\psi^{2}\approx A$:
$\eta_{\mathrm{gr}}\approx\frac{1/(l_{1}\sqrt{\bar{\theta}})+\sqrt{\bar{\theta}}/l_{2}}{A\operatorname{E}[G]\bar{\phi}},$ (155)
where $\operatorname{E}[G]\bar{\phi}$ is given by (153). Figure 18 represents the theoretical approximation (155), as well as results from simulation. As with element sampling, the efficiency for large $A$ is slightly lower than in RR and LRR (Figures 11 and 14), and for small $A$ there is almost no difference between simulation results and theoretical approximation even if $\phi$ is large.
(a) $l_{1}=1$, $l_{2}=1$, $\theta=16$, $\phi=0.1$
($p_{1}=0.4$, $p_{2}=0.025$)
(b) $l_{1}=3$, $l_{2}=1$, $\theta=1/16$, $\phi=0.01$
($p_{1}=0.0025$, $p_{2}=0.04$)
Line: approximation. Dots: simulation.
Figure 18: Efficiency for OR with group sampling
5 Estimation of log odds ratio
5.1 Element sampling
The estimation method for the LOR, $\Psi=\log(p_{1}(1-p_{2})/(p_{2}(1-p_{1})))$, has few differences compared to that presented for OR in Section 4. The second stage estimates $\log(p_{1}/(1-p_{1}))$ and $\log(p_{2}/(1-p_{2}))$ separately, using for each the method described in Mendo [16, section 3], which consists of two IBS procedures with the same parameter. Specifically, to estimate $\log(p_{i}/(1-p_{i}))$, $i=1,2$, IBS is applied to population $i$ until $s_{i}$ successes are obtained, which requires a random number $N_{i}^{\prime}$ of observations from this population, and then IBS is applied until $s_{i}$ failures are obtained, which requires $N_{i}^{\prime\prime}$ additional observations. These are the same steps as for OR, but with $\alpha=0$. Then, conditioned on $s_{1}$ and $s_{2}$, which are obtained in the first stage, $H_{N_{i}^{\prime}-1}+H_{N_{i}^{\prime\prime}-1}$ is a conditionally unbiased estimator of $\log(p_{i}/(1-p_{i}))$ with conditional variance less than $1/(s_{i}-5/4)$ for any $p_{i}\in(0,1)$ [16, theorem 3]. Therefore,
$\hat{\Psi}=-H_{N_{1}^{\prime}-1}+H_{N_{1}^{\prime\prime}-1}+H_{N_{2}^{\prime}-1}-H_{N_{2}^{\prime\prime}-1}$ (156)
is a conditionally unbiased estimator of $\Psi$ with
$\operatorname{Var}[\hat{\Psi}\mid s_{1},s_{2}]<\frac{1}{s_{1}-5/4}+\frac{1}{s_{2}-5/4}.$ (157)
It follows that $\hat{\Psi}$ is unconditionally unbiased, and, defining its error function as in (29) with
$\mu_{1}=5/4,\qquad\mu_{2}=5/4,\qquad\mu_{12}=0,$ (158)
inequality (31) guarantees that $\operatorname{Var}[\hat{\Psi}]<A$ for any $p_{1},p_{2}\in(0,1)$.
The number of samples from population $i=1,2$ used in the second stage, $N_{i}=N_{i}^{\prime}+N_{i}^{\prime\prime}$, has an average given by (123) and (124) with $\alpha=0$, that is, $\operatorname{E}[N_{i}]=\operatorname{E}[s_{i}]/\bar{p}_{i}$, where $\bar{p}_{i}$ is defined by (122). Thus, for the same reason as in OR estimation, the first stage for LOR estimation must be based on samples with parameters $\bar{p}_{i}$, $i=1,2$, generated from observations with parameters $p_{i}$. For each $i=1,2$, given $r_{i}$, a random number $\bar{M}_{i}$ of samples with parameter $\bar{p}_{i}$ are generated until $r_{i}$ successes are obtained, which in turn requires $M_{i}$ observations from population $i$. Then $X$ is computed from $\bar{M}_{1}$ and $\bar{M}_{2}$ as in (127). Since (141) holds, imposing the additional condition (142), $\gamma$ and $\delta_{1}$ are obtained from (143) and (144) as for OR, and (22) is satisfied. Following this, $s_{1}$ and $s_{2}$ are computed using (64)–(66), as for the other estimators, and then rounded with the restriction (31). The resulting average numbers of samples are bounded by (148) and (149).
The condition $c(A,r_{1},\xi)=0$, from which $r_{1}$ and $r_{2}$ are obtained, is expressed by (146). Because $\mu_{1}=\mu_{2}$ and $\mu_{12}=0$, this simplifies in the same way as for LRR, and has the positive solution
$A=\frac{1}{(3r_{1}/2+\mu_{1}+\xi)(r_{1}-1)}.$ (159)
The LOR estimator, like previous ones, uses $\xi=1$ and $r_{1}$ given by (81), which in this case, taking into account (159), is written as
$r_{1}=\displaystyle\max\left\{3,\left\lceil-1/4+\sqrt{(5/4)^{2}+2/(3A)}\right\rceil\right\}.$ (160)
As with previous estimators, this choice of $r_{1}$ yields average numbers of samples close to their minimum values with respect to this parameter (results not shown). The curve $c(A,r_{1},1)=0$, plotted in Figure 3, is almost indistinguishable from that for OR; and (160) corresponds to rounding up the ordinate values of this curve, with a minimum of $3$.
The estimation procedure for LOR is summarized in Algorithm 2 (see Appendix A), which also lists the properties of the estimator.
Particularizing (87) for $\zeta=\Psi$, with $\partial\Psi/\partial p_{1}=1/\bar{p}_{1}$, $\partial\Psi/\partial p_{2}=-1/\bar{p}_{2}$, and then using (148), (149) and the fact that $\operatorname{Var}[\hat{\Psi}]<A$, it follows that the efficiency with element sampling is bounded by (151) (with the value of $\mu_{1}$ corresponding to LOR and with $r_{1}$ computed accordingly). Figure 19 shows the results, which are very similar to those for OR (Figure 17).
(a) $\lambda=1$, $\theta=16$, $\phi=0.1$
($p_{1}=0.4$, $p_{2}=0.025$)
(b) $\lambda=3$, $\theta=1/16$, $\phi=0.01$
($p_{1}=0.0025$, $p_{2}=0.04$)
Line: bound. Dots: simulation.
Figure 19: Efficiency for LOR with element sampling
5.2 Group sampling
The group sampling procedure is analogous to that described for the other estimators. The average number of groups $\operatorname{E}[G]$ is computed from (93), where $\operatorname{E}[|\Delta|]$ satisfies (152), as shown in Appendix B.2. Thus $\operatorname{E}[G]$ is approximately given by (153) and (154), as for OR (but with the value of $\mu_{1}$ corresponding to LOR). Simulation results are omitted.
The efficiency with group sampling is expressed by (155), as for OR. The results, plotted in Figure 20, are not the same as in that case (Figure 18) because of the differences in $\mu_{1}$ and $r_{1}$, which affect $\operatorname{E}[G]$ and thus the efficiency; but they are seen to be very similar.
(a) $l_{1}=1$, $l_{2}=1$, $\theta=16$, $\phi=0.1$
($p_{1}=0.4$, $p_{2}=0.025$)
(b) $l_{1}=3$, $l_{2}=1$, $\theta=1/16$, $\phi=0.01$
($p_{1}=0.0025$, $p_{2}=0.04$)
Line: approximation. Dots: simulation.
Figure 20: Efficiency for LOR with group sampling
6 Conclusions
Two-stage sequential methods have been presented to estimate the RR $p_{1}/p_{2}$, the OR $p_{1}(1-p_{2})/(p_{2}(1-p_{1}))$ or their logarithmic versions using independent binary observations from two populations with parameters $p_{1}$ and $p_{2}$. The estimators are unbiased; guarantee that the relative mean square error, or the mean square error for the logarithmic versions, is less than a target value $A$ irrespective of $p_{1}$ and $p_{2}$; and approximately achieve a prescribed ratio of average sample sizes when samples are taken from each population individually (element sampling). The estimators can also be used with group sampling. In this case, samples are taken simultaneously from the two populations in fixed-size groups, and individual samples are extracted from those groups as needed, with a number of samples possibly discarded at the end of the process. The properties of unbiasedness and guaranteed accuracy are maintained with group sampling (and an exact sample size ratio is imposed by the sampling). Bounds and approximate expressions have been derived for the average sample sizes and the average number of sample groups, respectively. The estimation efficiency, defined in terms of the Cramér–Rao bound, has been characterized in the same way. The efficiency is generally good, both with element sampling and with group sampling; and is close to $1$ for small $A$. Algorithms 1 and 2 (Appendix A) specify the estimation procedure and summarize the properties of the estimators.
The described method can be extended to estimate other functions of $p_{1}$ and $p_{2}$, provided that an error function can be defined as in (29) and (63), and that a Bernoulli factory can be found, if needed, to generate samples with parameters equal to the probabilities to which $\operatorname{E}[N_{1}]$ and $\operatorname{E}[N_{2}]$ are approximately inversely proportional (as was done for OR and LOR), where $N_{1}$ and $N_{2}$ are the numbers of observations required in the second stage. In addition, the Bernoulli factory needs to have an average number of inputs per output equal to a constant $\beta$.
As an example, consider the estimation of $p_{1}p_{2}$. A straightforward approach would be to generate samples with parameter $p_{1}p_{2}$ from samples with parameters $p_{1}$ and $p_{2}$, using a procedure analogous to the Bernoulli factory described in Section 4, and then apply IBS to the generated samples. However, this does not give any control on the proportion of samples from the two populations, and the average number of samples required from each population scales with the inverse of $p_{1}p_{2}$. Instead, the method presented for RR and LRR can be used: two IBS processes with parameters $s_{1}$ and $s_{2}$ are respectively applied to the two populations (second stage), which requires $N_{1}$ and $N_{2}$ samples, where $s_{1}$ and $s_{2}$ are obtained from a previous pair of IBS processes with fixed parameters $r_{1}$ and $r_{2}$ (first stage). Then,
$\frac{(s_{1}-1)(s_{2}-1)}{(N_{1}-1)(N_{2}-1)}$
is an unbiased estimator of $p_{1}p_{2}$, and by means of (14) a target relative error can be guaranteed. The same expressions as in RR and LRR apply for the average sample sizes, average number of groups and efficiency. In particular, the average sample sizes with element sampling are approximately in the desired ratio, and, for $p_{1}/p_{2}$ fixed, they scale with the inverse of $\sqrt{p_{1}p_{2}}$.
Appendix A Estimation procedure and properties of the estimators
The estimation procedure for RR and LRR, and that for OR and LOR, are respectively described in Algorithms 1 and 2. The properties of the estimators are also summarized.
Algorithm 1 Estimator of $R$ or LRR
Input parameters:
  1. Target $A$, interpreted as relative MSE for RR, or MSE for LRR.
  2. Desired ratio of average sample sizes $\lambda$, or group sizes $l_{1}$, $l_{2}$.
Estimation procedure:
  1. 1.
    Define $\mu_{1},\mu_{2},\mu_{12}$ as in (30) for RR, or as in (116) for LRR.
    Define $\epsilon_{1}=\epsilon_{2}=1/2$ and $\xi=1$.
    Define $\lambda=l_{1}/l_{2}$ if group sampling is applied.
  2. 2.
    Compute $r_{1}$ from (81) with $c(A,r_{1},\xi)$ given by (76) (for LRR the explicit expression (119) can equivalently be used), or obtain it by rounding up the value from Figure 3, with a minimum of $3$. Compute $r_{2}$ as $r_{1}+\mu_{1}-\mu_{2}$.
  3. 3.
    Compute $\gamma$ and $\delta_{1}$ from (61) and (62).
  4. 4.
    (First sampling stage): For each $i=1,2$, repeatedly sample from population $i$ until $r_{i}$ successes are obtained. Let $M_{i}$ be the number of samples.
  5. 5.
    Compute $X$ from (21). Compute $s_{1}$, $s_{2}$ from (64)–(66), and then round one of them up and the other down, or both up, to fulfill (31).
  6. 6.
    (Second sampling stage): For each $i=1,2$, repeatedly sample from population $i$ until $s_{i}$ successes are obtained. Let $N_{i}$ be the number of samples.
  7. 7.
    Compute $\hat{\theta}$ from (24), or $\hat{\Theta}$ from (114).
Output:
  1. Estimation $\hat{\theta}$ or $\hat{\Theta}$.
Properties:
  1. Unbiased, with relative MSE (RR) or MSE (LRR) less than $A$.
  2. Average sample sizes: (79), (80); approximate ratio $\lambda$.
  3. Efficiency with element sampling: (89); approaches 11 for $A$, $\phi$ small.
  4. Average number of groups: (98), (109).
  5. Efficiency with group sampling: (112).
Algorithm 2 Estimator of OR or LOR
Input parameters:
  1. Target $A$, interpreted as relative MSE for OR, or MSE for LOR.
  2. Desired ratio of average sample sizes $\lambda$, or group sizes $l_{1}$, $l_{2}$.
Estimation procedure:
  1. 1.
    Define $\mu_{1},\mu_{2},\mu_{12}$ as in (131) for OR, or as in (158) for LOR.
    Define $\epsilon_{1}=\epsilon_{2}=1/2$ and $\xi=1$.
    Define $\alpha=2$ for OR, or $\alpha=0$ for LOR.
    Define $\lambda=l_{1}/l_{2}$ if group sampling is applied.
  2. 2.
    Compute $r_{1}$ from (81) with $c(A,r_{1},\xi)$ given by (145) (for LOR the explicit expression (160) can equivalently be used), or obtain it by rounding up the value from Figure 3, with a minimum of $3$. Set $r_{2}=r_{1}$.
  3. 3.
    Set $\gamma=\lambda$ and compute $\delta_{1}$ from (144).
  4. 4.
    (First sampling stage): For each $i=1,2$, using the method below, generate samples with parameter $\bar{p}_{i}$ until $r_{i}$ successes are obtained. Let $\bar{M}_{i}$ be the number of generated samples.
    To generate a sample with parameter $\bar{p}_{i}$, choose one of these options equally likely:
    • (a)
      Take a sample from population $i$. If it is a failure, output failure. Else, take another sample from population $i$ and output the opposite of its value.
    • (b)
      Take a sample from population $i$. If it is a success, output failure. Else, take another sample from population $i$ and output its value.
    Let $M_{i}$ be the total number of samples used from population $i$.
  5. 5.
    Compute $X$ from (127). Compute $s_{1}$, $s_{2}$ from (64)–(66), and then round one of them up and the other down, or both up, to fulfill (31).
  6. 6.
    (Second sampling stage): From population $1$, take as many samples as needed, $N^{\prime}_{1}$, to obtain $s_{1}$ successes; then as many as needed, $N^{\prime\prime}_{1}$, to obtain $s_{1}-\alpha$} failures. From population $2$, take as many samples as needed, $N^{\prime}_{2}$}, to obtain $s_{2}-\alpha$} successes; then as many as needed, $N^{\prime\prime}_{2}$, to obtain $s_{2}$ failures. Let $N_{1}=N_{1}^{\prime}+N_{1}^{\prime\prime}$ and $N_{2}=N_{2}^{\prime}+N_{2}^{\prime\prime}$.
  7. 7.
    Compute $\hat{\psi}$ from (128), or $\hat{\Psi}$ from (156).
Output:
  1. Estimation $\hat{\psi}$ or $\hat{\Psi}$.
Properties:
  1. Unbiased, with relative MSE (OR) or MSE (LOR) less than $A$.
  2. Average sample sizes: (148), (149); approximate ratio $\lambda$.
  3. Efficiency with element sampling: (151); approaches 11 for $A$ small.
  4. Average number of groups: (153), (154).
  5. Efficiency with group sampling: (155).
Appendix B Approximation of $\operatorname{E}[|\Delta|]$
B.1 For relative risk and log relative risk
The random variable $M_{1}$ in RR and LRR has a negative binomial distribution with parameters $r_{1}$ and $p_{1}=\phi\sqrt{\theta}$. For $\phi\rightarrow 0$, the distribution of $M_{1}\phi$ tends to a gamma distribution with location parameter $r_{1}$ and scale parameter $1/\sqrt{\theta}$, because
$\begin{split}\lim_{\phi\rightarrow 0}\Pr[M_{1}\phi\leq x]&=\lim_{\phi\rightarrow 0}\sum_{k=r_{1}}^{\lfloor x/\phi\rfloor}\binom{k-1}{r_{1}-1}\left(\phi\sqrt{\theta}\right)^{r_{1}}\left(1-\phi\sqrt{\theta}\right)^{k-r_{1}}\\ &=\frac{\theta^{r_{1}/2}}{(r_{1}-1)!}\int_{0}^{x}t^{r_{1}-1}\exp(-t\sqrt{\theta})\,\mathrm{d}\mkern 0.8mut.\end{split}$ (161)
Likewise, the distribution of $M_{2}\phi$ tends to that of a gamma random variable with location parameter $r_{2}$ and scale parameter $\sqrt{\theta}$. By the continuous mapping theorem [21, theorem 2.3], the variable $X$ defined in (21) converges in distribution to the ratio of these two gamma random variables, and therefore $Y=X/\theta$ and $1/Y$ converge in distribution to beta prime random variables with parameters $r_{2}$, $r_{1}$ and $r_{1}$, $r_{2}$ respectively [8, section 4.4]. It is easy to see that $Y^{2}$ and $1/Y^{2}$ are uniformly integrable as $\phi\rightarrow 0$, which implies [6, theorem 25.12] that the variances of $Y$ and $1/Y$ converge to those of the referred beta prime distributions. Thus, approximating $\operatorname{Var}[X]$ and $\operatorname{Var}[1/X]$ by their values for $\phi\rightarrow 0$, the following expressions are obtained:
$\operatorname{Var}[X]$ $\approx\frac{r_{2}(r_{1}+r_{2}-1)\theta^{2}}{(r_{1}-2)(r_{1}-1)^{2}},$ (162)
$\operatorname{Var}\left[\frac{1}{X}\right]$ $\approx\frac{r_{1}(r_{1}+r_{2}-1)}{(r_{2}-2)(r_{2}-1)^{2}\theta^{2}}.$ (163)
The value of $\operatorname{E}[|\Delta|]$ obviously depends both on the mean of $\Delta$ and on the variability of $\Delta$ with respect to its mean. In view of (92), this variability can be understood as arising from three sources: (i) the variability of $M_{1}$ and $M_{2}$ as they directly appear in that expression; (ii) the variability of $N_{1}$ and $N_{2}$ conditioned on $s_{1}$ and $s_{2}$; and (iii) the variability of $N_{1}$ and $N_{2}$ caused by variations of $s_{1}$ and $s_{2}$ (since the values of $s_{1}$ and $s_{2}$ are obtained from $X$ this is another, indirect effect of $M_{1}$ and $M_{2}$). As will be seen, the contributions of the first two sources of variability are small compared to that of the third, and can be neglected with little error. This statement can be made more precise by expressing it in terms of variance. Applying the law of total variance [5, theorem 12.2.6],
$\operatorname{Var}\left[\Delta\right]=\operatorname{E}\left[\operatorname{Var}\left[\Delta\mid M_{1},M_{2}\right]\right]+\operatorname{Var}\left[\operatorname{E}\left[\Delta\mid M_{1},M_{2}\right]\right].$ (164)
Substituting (92) into (164), and noting that $N_{1}$, $N_{2}$ depend on $M_{1}$, $M_{2}$ only through $s_{1}$ and $s_{2}$,
$\begin{split}\displaystyle\operatorname{Var}\left[\Delta\right]&=\operatorname{Var}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\right]+\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]+\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\\ &\quad+2\operatorname{Cov}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}},\,\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right].\end{split}$ (165)
The first three summands on the right-hand side of (165) correspond to (i), (ii) and (iii) respectively, and the fourth represents the statistical relationship between (i) and (iii) (both of which stem from the variability of $M_{1}$ and $M_{2}$).
The first summand in (165) can be written, taking into account that $M_{1}$ and $M_{2}$ are independent, as $\operatorname{Var}[M_{1}/l_{1}]+\operatorname{Var}[M_{2}/l_{2}]$, and then applying (13) gives
$\operatorname{Var}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\right]\phi^{2}l_{1}l_{2}=\frac{r_{1}(1-p_{1})}{\lambda\theta}+r_{2}(1-p_{2})\lambda\theta\approx\frac{r_{1}}{\lambda\theta}+(r_{1}+\mu_{1}-\mu_{2})\lambda\theta.$ (166)
The second summand is computed as follows. Noting that $N_{1}$ and $N_{2}$ are conditionally independent given $s_{1}$ and $s_{2}$, the conditional variance $\operatorname{Var}[N_{1}/l_{1}-N_{2}/l_{2}\mid s_{1},s_{2}]$ can be expressed as $\operatorname{Var}[N_{1}/l_{1}\mid s_{1}]+\operatorname{Var}[N_{2}/l_{2}\mid s_{2}]$, and using (13) again yields
$\begin{split}\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\phi^{2}l_{1}l_{2}&=\operatorname{E}\left[\frac{s_{1}(1-p_{1})}{\lambda\theta}\right]+\operatorname{E}\left[s_{2}(1-p_{2})\lambda\theta\right]\\ &\approx\frac{\operatorname{E}[s_{1}]}{\lambda\theta}+\operatorname{E}[s_{2}]\lambda\theta.\end{split}$ (167)
Then, from (37)–(42), (52), (53) and (60)–(62), and including the rounding term $\xi$ when computing $\operatorname{E}[s_{1}]$ and $\operatorname{E}[s_{2}]$,
$\begin{split}\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\phi^{2}l_{1}l_{2}&\approx\frac{1}{\lambda\theta}\left(\frac{1}{A}+\mu_{1}+\xi\right)+\lambda\theta\left(\frac{1}{A}+\mu_{2}+\xi\right)\\ &\quad+2\left(\frac{1}{A}+r_{1}+\mu_{1}+\xi\right).\end{split}$ (168)
As for the third summand, from (11), (37) and (38) it can be written as
$\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\phi^{2}l_{1}l_{2}=\operatorname{Var}\left[\frac{s_{1}}{l_{1}p_{1}}-\frac{s_{2}}{l_{2}p_{2}}\right]\phi^{2}l_{1}l_{2}\\ \approx\frac{b_{1}^{2}\operatorname{Var}[X]}{\lambda\theta}+\lambda\theta b_{2}^{2}\operatorname{Var}\left[\frac{1}{X}\right]+2b_{1}b_{2}\operatorname{Cov}\left[X,-\frac{1}{X}\right].$ (169)
The term $\operatorname{Cov}[X,-1/X]$ is easily obtained using (52), (53) and (60):
$\operatorname{Cov}\left[X,-\frac{1}{X}\right]=-1+\operatorname{E}\left[X\right]\operatorname{E}\left[\frac{1}{X}\right]\approx\frac{r_{1}+r_{2}-1}{(r_{1}-1)(r_{2}-1)}.$ (170)
Substituting (40), (42), (162), (163) and (170) into (169) and using (60)–(62),
$\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\phi^{2}l_{1}l_{2}\approx\left(\frac{1}{A}+r_{1}+\mu_{1}+\xi\right)^{\!2}(2r_{1}+\mu_{1}-\mu_{2}-1)\\ \cdot\left(\frac{\lambda\theta}{(r_{1}-2)(r_{1}+\mu_{2}-\mu_{1})}+\frac{1}{\lambda\theta r_{1}(r_{1}+\mu_{2}-\mu_{1}-2)}+\frac{2}{r_{1}(r_{1}+\mu_{2}-\mu_{1})}\right).$ (171)
Lastly, the fourth summand in (165) can be bounded using the Cauchy-Schwarz inequality [5, proposition 6.2.8]:
$\left|2\operatorname{Cov}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}},\,\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\right|\\ \leq 2\sqrt{\operatorname{Var}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\right]\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]}.$ (172)
It will be established next that (166), (168) and the right-hand side of (172) are much smaller than (171). To this end, the following observations will be useful. From (60) and (63) it stems that $r_{1}\approx r_{2}=r_{1}+\mu_{1}-\mu_{2}$ with little approximation error. In practice, the target $A$ will typically be smaller, or much smaller, than $1$. For example, a relative RMSE of $10\%$ corresponds to $A=0.01$. On the other hand, $\xi=1$ is small compared to $1/A$, and so are the values of $r_{1}$ obtained from (81); namely $r_{1}\approx\sqrt{1/A}$, as can be seen from (82) or in Figure 3. In the example, $A=0.01$ gives $r_{1}=9$ for RR and $10$ for LRR.
To show that (166) is much smaller than (171), it is convenient to study the cases $\lambda\theta\approx 1$, $\lambda\theta\gg 1$ and $\lambda\theta\ll 1$ separately. For $\lambda\theta\approx 1$, the approximations in the above paragraph imply that (166) and (171) reduce to $2r_{1}$ and $8(1/A+r_{1})^{2}/r_{1}$ respectively, and their ratio,
$\frac{r_{1}^{2}}{4(1/A+r_{1})^{2}},$
is much smaller than $1$. For $\lambda\theta\gg 1$, (166) and (171) are approximated as $r_{1}\lambda\theta$ and $2\lambda\theta(1/A+r_{1})^{2}/r_{1}$ respectively, and their ratio is
$\frac{r_{1}^{2}}{2(1/A+r_{1})^{2}},$
which is again small compared with 11. The case $\lambda\theta\ll 1$ gives the same result. Thus, for any value of $\lambda\theta$, the first summand in (165) can be approximately neglected in comparison with the third.
Proceeding analogously to compare (168) with (171), for $\lambda\theta\approx 1$ their approximate values are $4/A+2r_{1}$ and $8(1/A+r_{1})^{2}/r_{1}$ respectively, with a ratio
$\frac{(4/A+2r_{1})r_{1}}{8(1/A+r_{1})^{2}}=\frac{(1/A+r_{1}/2)r_{1}}{2(1/A+r_{1})^{2}}<\frac{r_{1}}{2(1/A+r_{1})},$
which is significantly smaller than $1$. For $\lambda\theta\gg 1$, (168) and (171) are approximated as $\lambda\theta/A$ and $2\lambda\theta(1/A+r_{1})^{2}/r_{1}$ respectively, and this gives a ratio
$\frac{(1/A)r_{1}}{2(1/A+r_{1})^{2}}<\frac{r_{1}}{2(1/A+r_{1})},$
which is again small compared with $1$. For $\lambda\theta\ll 1$ the result is the same. Thus the second summand in (165) can also, to a good approximation, be neglected in comparison with the third.
As for (172), dividing its right-hand side by the left-hand side of (171) gives twice the square root of the ratio between (166) and (171). Therefore the fourth summand in (165) is also significantly smaller than the third.
The conclusion of the preceding analysis is that (165) can be approximated by keeping only the third summand in the right-hand side, as it is significantly larger than the others. This means that the variability in $\Delta$ is mostly due to the variability of $N_{1}$ and $N_{2}$ caused by variations of $s_{1}$ and $s_{2}$, i.e. (iii) as defined above. The variability of $M_{1}$ and $M_{2}$, (i), and that of $N_{1}$ and $N_{2}$ conditioned on $s_{1}$ and $s_{2}$, (ii), are comparatively smaller. Therefore, to compute $\operatorname{E}[|\Delta|]$, the variables $M_{1}$, $M_{2}$ in (92) can be replaced by their means, and $N_{1}$, $N_{2}$ can be replaced by their conditional means given $s_{1}$, $s_{2}$, which yields (94).
B.2 For odds ratio and log odds ratio
The law of total variance [5, theorem 12.2.6] conditioning on $\bar{M}_{1}$ and $\bar{M}_{2}$ gives, for OR and LOR,
$\operatorname{Var}\left[\Delta\right]=\operatorname{E}\left[\operatorname{Var}\left[\Delta\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right]+\operatorname{Var}\left[\operatorname{E}\left[\Delta\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right].$ (173)
The variables $N_{1}^{\prime}$, $N_{1}^{\prime\prime}$, $N_{2}^{\prime}$, $N_{2}^{\prime\prime}$, and therefore $N_{1}$, $N_{2}$, depend on $M_{1}$, $M_{2}$ only through $\bar{M}_{1}$, $\bar{M}_{2}$. This implies that $M_{1}$, $M_{2}$, $N_{1}$, $N_{2}$, are conditionally independent given $\bar{M}_{1}$, $\bar{M}_{2}$. Thus, using (92), the first term in the right-hand side of (173) is written as
$\operatorname{E}\left[\operatorname{Var}\left[\Delta\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right]=\operatorname{E}\left[\operatorname{Var}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right]+\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right],$ (174)
whereas the second term is
$\begin{split}\displaystyle\operatorname{Var}\left[\operatorname{E}\left[\Delta\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right]&=\operatorname{Var}\left[\operatorname{E}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right]+\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right]\\ &\quad+2\operatorname{Cov}\left[\operatorname{E}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\!\!,\,\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right].\end{split}$ (175)
On the other hand, making use of the law of total variance again,
$\operatorname{E}\left[\operatorname{Var}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right]+\operatorname{Var}\left[\operatorname{E}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right]=\operatorname{Var}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\right].$ (176)
Combining (174)–(176) with (173), and noting that conditioning $N_{1}$ or $N_{2}$ on $\bar{M}_{1}$, $\bar{M}_{2}$ is equivalent to conditioning on $s_{1}$, $s_{2}$, the following expression is obtained for $\operatorname{Var}[\Delta]$:
$\begin{split}\displaystyle\operatorname{Var}\left[\Delta\right]&=\operatorname{Var}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\right]+\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]+\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\\ &\quad+2\operatorname{Cov}\left[\operatorname{E}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\!\!,\,\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right].\end{split}$ (177)
This is similar to the decomposition (165) in RR and LRR, where only the covariance term is different. Although the approximate expressions of the summands for small $\phi$, to be computed next, are different from that case, it will be shown that for OR and LOR the third summand again dominates the other three.
The first summand in (177) is written, thanks to the independence of $M_{1}$ and $M_{2}$, as
$\operatorname{Var}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\right]=\frac{\operatorname{Var}[M_{1}]}{l_{1}^{2}}+\frac{\operatorname{Var}[M_{2}]}{l_{2}^{2}}.$ (178)
From the law of total variance, $\operatorname{Var}[M_{i}]$, $i=1,2$ is expressed as
$\operatorname{Var}\left[M_{i}\right]=\operatorname{E}\left[\operatorname{Var}\left[M_{i}\mid\bar{M}_{i}\right]\right]+\operatorname{Var}\left[\operatorname{E}\left[M_{i}\mid\bar{M}_{i}\right]\right].$ (179)
The first stage applies IBS to samples with parameter $\bar{p}_{i}$, using a number $\bar{M}_{i}$ of those samples, of which $r_{i}$ are successes. Each sample with parameter $\bar{p}_{i}$ is generated by the Bernoulli factory described in Section 4.1, taking samples with parameter $p_{i}$ as inputs. With this factory, the average number of inputs needed to produce an output is $3/2$; a success output always uses $2$ inputs, and a failure output uses either $1$ or $2$ inputs. Let $\pi_{i}$ denote the probability that $2$ inputs are used, conditioned on the output being a failure. Then $3/2=2\bar{p}_{i}+(1+\pi_{i})(1-\bar{p}_{i})$, from which
$\pi_{i}=1-\frac{1}{2(1-\bar{p}_{i})}=\frac{1-2\bar{p}_{i}}{2(1-\bar{p}_{i})}.$ (180)
Thus, conditioned on $\bar{M}_{i}$, the average number of required inputs is
$\operatorname{E}[M_{i}\mid\bar{M}_{i}]=2r_{i}+(1+\pi_{i})(\bar{M}_{i}-r_{i})=\frac{(3-4\bar{p}_{i})\bar{M}_{i}+r_{i}}{2(1-\bar{p}_{i})}.$ (181)
The term $\operatorname{Var}[M_{i}\mid\bar{M}_{i}]$ in (179) equals the variance of a binomial random variable with parameters $\bar{M}_{i}-r_{i}$ and $\pi_{i}$. Therefore, computing $\operatorname{E}[\bar{M}_{i}]$ from (11) and substituting (180),
$\operatorname{E}\left[\operatorname{Var}\left[M_{i}\mathrel{}\middle|\mathrel{}\bar{M}_{i}\right]\right]=\pi_{i}(1-\pi_{i})\operatorname{E}[\bar{M}_{i}-r_{i}]=\frac{(1-2\bar{p}_{i})r_{i}}{4\bar{p}_{i}(1-\bar{p}_{i})}.$ (182)
On the other hand, using (181) and computing $\operatorname{Var}[\bar{M}_{i}]$ from (13),
$\operatorname{Var}\left[\operatorname{E}\left[M_{i}\mathrel{}\middle|\mathrel{}\bar{M}_{i}\right]\right]=\frac{(3-4\bar{p}_{i})^{2}}{4(1-\bar{p}_{i})^{2}}\operatorname{Var}\left[\bar{M}_{i}\right]=\frac{(3-4\bar{p}_{i})^{2}r_{i}}{4\bar{p}_{i}^{2}(1-\bar{p}_{i})}.$ (183)
From (179), (182) and (183),
$\operatorname{Var}\left[M_{i}\right]=\frac{(14\bar{p}_{i}^{2}-23\bar{p}_{i}+9)r_{i}}{4\bar{p}_{i}^{2}(1-\bar{p}_{i})}=\frac{(9-14\bar{p}_{i})r_{i}}{4\bar{p}_{i}^{2}}.$ (184)
Thus, for $\phi$ small, which implies $\bar{\phi}$ small, substituting (184) into (178) and using (142),
$\operatorname{Var}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\right]\bar{\phi}^{2}l_{1}l_{2}=\frac{9r_{1}}{4}\left(\frac{1-14\bar{p}_{1}/9}{\lambda\bar{\theta}}+(1-14\bar{p}_{2}/9)\lambda\bar{\theta}\right)\approx\frac{9r_{1}}{4}\left(\frac{1}{\lambda\bar{\theta}}+\lambda\bar{\theta}\right).$ (185)
The second summand in (177) is computed as follows. Using the fact that $N_{1}^{\prime}$, $N_{1}^{\prime\prime}$, $N_{2}^{\prime}$, $N_{2}^{\prime\prime}$ are conditionally independent given $s_{1}$, $s_{2}$, taking into account (13) and approximating for $\phi$ small,
$\begin{split}&\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\bar{\phi}^{2}l_{1}l_{2}\\ &\ =\frac{\operatorname{E}[s_{1}](1-p_{1})\bar{\phi}^{2}}{\lambda p_{1}^{2}}+\frac{\operatorname{E}[s_{1}-\alpha]p_{1}\bar{\phi}^{2}}{\lambda(1-p_{1})^{2}}+\frac{\lambda\operatorname{E}[s_{2}-\alpha](1-p_{2})\bar{\phi}^{2}}{p_{2}^{2}}+\frac{\lambda\operatorname{E}[s_{2}]p_{2}\bar{\phi}^{2}}{(1-p_{2})^{2}}\\ &\ =\left(\frac{\operatorname{E}[s_{1}]((1-p_{1})^{3}+p_{1}^{3})-\alpha p_{1}^{3}}{\lambda\bar{p}_{1}^{2}}+\frac{\lambda\left(\operatorname{E}[s_{2}]((1-p_{2})^{3}+p_{2}^{3})-\alpha(1-p_{2})^{3}\right)}{\bar{p}_{2}^{2}}\right)\bar{\phi}^{2}\\ &\ =\frac{\operatorname{E}[s_{1}](1-3\bar{p}_{1})-\alpha p_{1}^{3}}{\lambda\bar{\theta}}+\lambda\bar{\theta}\left(\operatorname{E}[s_{2}](1-3\bar{p}_{2})-\alpha(1-p_{2})^{3}\right)\\ &\ \approx\frac{\operatorname{E}[s_{1}]}{\lambda\bar{\theta}}+\lambda\bar{\theta}(\operatorname{E}[s_{2}]-\alpha),\end{split}$ (186)
with $\alpha$ equal to $2$ for OR and 0 for LOR. Using (37) and (38), including the rounding term $\xi$, and then substituting (39)–(42), (135), (136) and (141)–(144), (186) becomes
$\begin{split}\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\bar{\phi}^{2}l_{1}l_{2}&\approx\frac{1}{\lambda\bar{\theta}}\left(\frac{1}{A}+\mu_{1}+\xi\right)+\lambda\bar{\theta}\left(\frac{1}{A}+\mu_{1}+\xi-\alpha\right)\\ &\quad+2\left(\frac{1}{A}+\frac{3r_{1}}{2}+\mu_{1}+\xi\right).\end{split}$ (187)
As for the third summand, using (120)–(122) and then taking into account (37) and (38),
$\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\bar{\phi}^{2}l_{1}l_{2}\\ \approx\frac{b_{1}^{2}\operatorname{Var}[X]}{\lambda\bar{\theta}}+\lambda\bar{\theta}b_{2}^{2}\operatorname{Var}\left[\frac{1}{X}\right]+2b_{1}b_{2}\operatorname{Cov}\left[X,-\frac{1}{X}\right].$ (188)
For $\phi$ small, $\operatorname{Var}[X]$, $\operatorname{Var}[1/X]$ and $\operatorname{Cov}[X,-1/X]$ are approximately given by the same expressions (162), (163) and (170) as in RR and LRR except with $\theta$ replaced by $\bar{\theta}$ and with $r_{1}=r_{2}$. According to this, and making use of (141)–(144), (40) and (42), (188) yields
$\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\bar{\phi}^{2}l_{1}l_{2}\\ \approx\left(\frac{1}{A}+\frac{3r_{1}}{2}+\mu_{1}+\xi\right)^{\!2}(2r_{1}-1)\left(\frac{1}{r_{1}(r_{1}-2)}\left(\lambda\bar{\theta}+\frac{1}{\lambda\bar{\theta}}\right)+\frac{2}{r_{1}^{2}}\right).$ (189)
The fourth summand in (177) is bounded by means of the Cauchy-Schwarz inequality as in Appendix B.1; and then, using (181) and the fact that $(3-4\bar{p}_{i})/(2(1-\bar{p}_{i}))$ is less than $3/2$ (and approaches that value for $\phi$ small),
$\begin{split}&\left|2\operatorname{Cov}\left[\operatorname{E}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\!\!,\,\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\right|\\ &\quad\leq 2\sqrt{\operatorname{Var}\left[\operatorname{E}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right]\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]}\\ &\quad<3\sqrt{\operatorname{Var}\left[\frac{\bar{M}_{1}}{l_{1}}-\frac{\bar{M}_{2}}{l_{2}}\right]\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]}.\end{split}$ (190)
The variables $\bar{M}_{1}$ and $\bar{M}_{2}$ are independent. Combined with (13) and (142), this yields
$\operatorname{Var}\left[\frac{\bar{M}_{1}}{l_{1}}-\frac{\bar{M}_{2}}{l_{2}}\right]\bar{\phi}^{2}l_{1}l_{2}=\frac{r_{1}(1-\bar{p}_{1})}{\lambda\bar{\theta}}+r_{2}(1-\bar{p}_{2})\lambda\bar{\theta}\approx r_{1}\left(\frac{1}{\lambda\bar{\theta}}+\lambda\bar{\theta}\right).$ (191)
From (185), (LABEL:eq:_Cov_vasaf_vasas_OR) and (191) it follows that the fourth summand is asymptotically bounded for $\phi$ small as
$\left|2\operatorname{Cov}\left[\operatorname{E}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\!\!,\,\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\right|\\ \quad<2\sqrt{\operatorname{Var}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\right]\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]}.$ (192)
Using (185), (187), (189) and (192), a similar assessment as in Appendix B.1 can be made: for $\phi$ small; typical values of $A$, for which $1/A$ is large and $r_{1}$ is considerably smaller, namely $r_{1}\approx\sqrt{2/(3A)}$; both for OR ($\mu_{1}=\mu_{2}=2$, $\alpha=2$) and LOR ($\mu_{1}=\mu_{2}=5/4$, $\alpha=0$); and for any $\lambda$ and $\bar{\theta}$, the third summand in the right-hand side of (177) dominates the result for $\operatorname{Var}[\Delta]$. Specifically, the first summand divided by the third takes the following approximate values, where the first corresponds to the case $\lambda\bar{\theta}\approx 1$ and the second to $\lambda\bar{\theta}\gg 1$ and to $\lambda\bar{\theta}\ll 1$:
$\frac{(3r_{1}/2)^{2}}{4(1/A+3r_{1}/2)^{2}},\quad\frac{(3r_{1}/2)^{2}}{2(1/A+3r_{1}/2)^{2}}.$
The ratio of the second and third summands has, for the three cases, the upper bound
$\frac{r_{1}}{2(1/A+3r_{1}/2)}.$
Lastly, the absolute value of the fourth summand in (177) divided by the third is, in view of (192), less than twice the square root of the ratio of first to third summands. Thus, all these ratios are seen to be significantly less than $1$.
It is concluded from the above analysis that the variability of $\Delta$ in OR and LOR, as in RR and LRR, is mostly caused by that of $N_{1}$ and $N_{2}$ induced by the variations of $s_{1}$ and $s_{2}$. This implies that to compute $\operatorname{E}[|\Delta|]$ each variable $M_{i}$, $i=1,2$ in (92) can be replaced by its mean, given by (132), and $N_{i}$ can be replaced by its conditional mean given $s_{i}$, which is approximately $s_{i}/\bar{p}_{i}$ according to (120) and (121). This establishes (152).
Acknowledgments
The author would like to thank Dr. Ben O’Neill for advice on terminology.